### v1

# Spurious reconstruction from brain activity

Abstract
The rapid advances in brain decoding, particularly visual image reconstruction, have sparked discussions about the potential societal implications and ethical considerations surrounding neurotechnology. As these methods aim to recover perceived images from brain activity and achieve prediction over diverse images beyond training samples (zero-shot prediction), it is crucial to critically assess their capabilities and limitations to prevent misguided public expectations and inform future regulations. Our case study of recent text-guided reconstruction methods, which leverage a large-scale dataset (the Natural Scene Dataset, NSD) and text-to-image diffusion models, reveals significant limitations in their generalizability. We found a notable decrease in performance when applying these methods to a different dataset, which was designed to prevent category overlaps between training and test sets. UMAP visualization of the text features with NSD images showed a limited diversity of distinct semantic and visual clusters, with substantial overlap between training and test sets. Formal analysis and simulations demonstrated that clustered training samples can lead to “output dimension collapse,” restricting the output feature dimensions predictable from brain activity. Diversifying the training set to ensure a broader feature distribution improved generalizability beyond the trained clusters. However, text features alone are insufficient for a complete mapping to the visual space, even if perfectly predicted from brain activity. We argue that recent photo-like reconstructions may primarily be a blend of classification into trained categories and the generation of convincing yet inauthentic images through text-to-image diffusion (hallucination). To achieve genuine zero-shot prediction, diverse datasets and compositional representations spanning the image space are essential. As neurotechnology advances, engaging in interdisciplinary discussions involving neuroscientists, ethicists, policymakers, and the public is crucial to ensure responsible development and application of these techniques. These discussions should be grounded in a clear understanding of the current capabilities and limitations of the technology, as well as a careful consideration of the potential ethical and societal impacts.
Keywords Visual image reconstruction  $\cdot$ brain decoding  $\cdot$ zero-shot prediction  $\cdot$ methodology  $\cdot$ generative AIs
1 Introduction
Brain decoding has been widely used in the neuroscience field, revealing specific contents of the mind (Haxby et al., 2001; Kamitani and Tong, 2005; Soon et al., 2008; Horikawa et al., 2013). As brain decoding is sometimes referred to as “mind-reading” in popular media (Somers, 2021; Whang, 2023; Raasch, 2023), it has attracted significant attention beyond the scientific community due to its potential for real-world applications in medicine and industry. Such neurotechnology has also started to affect future ethical discussions and legal regulations (UNESCO, 2023). To prevent misguiding public expectations and policies, scientists need to carefully assess the current status of brain decoding techniques and clarify the possibilities and limitations.
One of the major challenges in brain decoding is the limited amount of brain data we can collect. The current brain measurement devices are costly, yielding far less brain data than the amounts typically used in image or text processing within the field of computer science and AI (Deng et al., 2009; Schuhmann et al., 2022). Although we have gradually increased the amount of brain data per subject (Van Essen et al., 2012; Allen et al., 2022; Naselaris et al., 2021; Hebart et al., 2023; Xu et al., 2024), it remains impractical to collect brain data covering the full range of cognitive states and perceptual experiences. Consequently, classification-based decoding approaches, primarily developed in the early stages of this field, are insufficient to uncover the neural code in general or natural conditions since decodable information is confined to the same stimuli or predefined categories used in a training phase.
To overcome this limitation, several decoding methods have been proposed to enable the prediction of novel contents from brain activities that are not encountered during the training phase. Kay et al. (2008) proposed a general visual decoding approach via a statistical encoding model that predicted fMRI voxel values from image features. It successfully identified novel test images from a set of 1000 candidates. Mitchell et al. (2008) utilized co-occurrence rates of specific verb sets for nouns to predict fMRI brain activity during the perception of line drawing images, demonstrating the ability to predict unseen nouns. Brouwer and Heeger (2009) constructed a color-tuning model and predicted brain activity while the subjects were presented with color stimuli. As their training stimuli cover most of the color space, their methods can successfully identify novel colors not included in the training dataset. Horikawa and Kamitani (2017) utilized deep neural network (DNN) features to decode brain activity measured while subjects perceived natural images. They showed successful prediction of novel object categories not encountered during the training phase.
In the field of machine learning, “zero-shot” prediction refers to the ability of a model to accurately predict or classify novel contents that were not encountered during the training phase (Larochelle et al., 2008; Palatucci et al., 2009). This technique has found widespread application across various domains, including image classification (Radford et al., 2021), image generation (Ramesh et al., 2021), and natural language processing (Brown et al., 2020). The concept of zero-shot prediction can be considered analogous to brain decoding techniques that aim to interpret brain activity patterns associated with previously unseen stimuli or experiences. Both approaches seek to generalize knowledge gained from a limited set of training data to novel situations, enabling the interpretation of new information without explicit prior exposure. To achieve effective zero-shot prediction, the model often utilizes a compositional representation of the output (Lake et al., 2017; Higgins et al., 2018). Compositional representation refers to the ability to understand and generate novel combinations of previously learned concepts or features. By learning the underlying structure and relationships between different elements, the model can generalize its knowledge to new, unseen instances.
Visual image reconstruction is another prominent example of zero-shot prediction in brain decoding. This task aims to recover perceived novel images that were not encountered during the training phase, effectively reconstructing visual experiences from brain activity patterns (Stanley et al., 1999; Miyawaki et al., 2008). As our perceptual visual experiences cannot be fully covered by limited brain data, reconstruction methods require strong generalizability. Miyawaki et al. (2008) conducted a study that demonstrated the reconstruction of perceived arbitrary 10x10 binary-contrast images from brain activity. They built multiple modular decoders to predict the local contrasts of each location and combined their predictions. This approach leverages the compositional representation of the visual field, which is organized retinotopically in the early visual cortex. Incorporating cortical organization into the model’s architecture can significantly improve its ability to perform zero-shot prediction and reconstruct novel visual experiences from brain activity. Although the training stimuli were only 400 random images, it was possible to reconstruct an arbitrary image from a set of possible $2^{100}$ instances, including geometric shapes such as crosses and alphabets. Similarly, Shen et al. (2019b) replaced local decoders with DNN feature decoders. Although their training stimuli were 1200 natural images, they demonstrated reconstructing novel images, including artificial images, which were not part of the training set. These successes suggest that the proposed reconstruction models capture rich and comprehensive information about the general aspects of the neural code, beyond merely the information defined by the training data (Kriegeskorte and Douglas, 2019). Developing reliable reconstruction methods also enables further analysis of subjective visual experiences, such as visual imagery (Shen et al., 2019b), attention (Horikawa and Kamitani, 2022), and illusion (Cheng et al., 2023). Enabling the decoding of novel brain states that were never encountered during the training phase can be a promising approach to neural mind-reading (Kamitani and Tong, 2005).
Visual image reconstruction pipelines can be divided into three main components: translator, latent features, and generator (Figure 1). The translator serves to convert a brain activity pattern into a latent feature space via linear regression (Shen et al., 2019b; Seeliger et al., 2018; Mozafari et al., 2020; Ozcelik et al., 2022; Kavasidis et al., 2017), or nonlinear transformation (Qiao et al., 2020). Latent features serve as surrogate representations of perceived visual images. While the local contrasts in Miyawaki et al. (2008) can be seen as a primitive form of latent features, recent studies often use DNN features, such as the intermediate output of recognition models (Shen et al., 2019b; Cheng et al., 2023). The generator visualizes translated latent features into images. Some studies have used the pretraining image generative models for the generator module (Mozafari et al., 2020; Qiao et al., 2020; Ozcelik et al., 2022; Kavasidis et al., 2017). Image optimization can also be regarded as a generator (Shen et al., 2019b). Direct mapping from brain activity to images using DNNs can also be considered to implicitly contain these components (Shen et al., 2019a; Beliy et al., 2019; Ren et al., 2021; Gaziv et al., 2022; Lin et al., 2022; Chen et al., 2023a).
Figure 1: An overview of the visual image reconstruction pipeline. Brain activity recorded while perceiving a visual image is fed into a translator module. The translator maps the brain activity to latent features. The generator module converts the latent features into an image corresponding to the original visual stimulus.
Recent advances in generative AI, particularly in text-to-image generation, have naturally given rise to expectations that these techniques could provide a shortcut for visual image reconstruction by leveraging semantic representations. In addition, there has been a growing trend towards collecting neural datasets using a wide range of diverse visual and semantic content. This shift aims to capture a more comprehensive and ecologically valid representation of the human experience (Naselaris et al., 2021). Researchers have started to collect large-scale datasets, such as the Natural Scene Dataset (NSD) (Allen et al., 2022) and the THINGS-fMRI dataset (Hebart et al., 2023), which incorporate a broader range of fMRI data induced by diverse visual stimuli with text or category annotations. Notably, recent studies leveraging the latest generative AI techniques with semantic representation and large-scale datasets have reported photo-like reconstructions from brain activity (Takagi and Nishimoto, 2023b; Ozcelik and VanRullen, 2023; Scotti et al., 2023; Bai et al., 2023; Chen et al., 2023b; Lan et al., 2023; Benchetrit et al., 2024). They utilized contrastive language-image pretraining (CLIP) features (Radford et al., 2021) as latent features and text-to-image diffusion models (Ramesh et al., 2021; Xu et al., 2022) as generators.
While these recent approaches show promising results, it remains uncertain whether these methods truly achieve zero-shot reconstruction due to several factors. The complex model architectures employed in these studies, along with the use of a large-scale dataset, make it challenging to interpret and understand the underlying mechanisms driving the reconstruction process. To fully assess the zero-shot prediction capabilities of these approaches, it is essential to rigorously test their generalizability across different datasets and to provide detailed analyses of the individual model components. This includes evaluating the performance of the translators, latent features, and generators used in these methods. Furthermore, the characterization of the diversity of stimuli in the datasets and the latent representations has not been thoroughly explored. It is unclear whether the recently proposed datasets, such as the NSD, are optimally designed to capture the full range of human visual experiences and to support the development of truly generalizable prediction models.
In the following, we begin with a case study that critically tests the methods proposed by Takagi and Nishimoto (2023b) and Ozcelik and VanRullen (2023), which originally used the NSD. Our findings reveal several issues, including the failure of replication with another dataset specifically designed to avoid overlap between training and test sets (Shen et al., 2019b), and the questionable post-hoc selection procedures of Takagi and Nishimoto (2023b) that can produce convincing reconstructions even with random brain data. We identify the lack of diversity represented by the limited number of clusters in the NSD dataset as a potential factor contributing to these issues. We also demonstrate the failure of zero-shot prediction in the feature space and the inability to recover a stimulus from its latent features, suggesting a misspecification of the model. Based on these findings, we conclude that the apparent photo-like reconstructions are essentially a result of classification to clusters shared between training and test sets, combined with hallucinations by the generative model.
In the formal analysis and simulation section, we aim to uncover the general factors underlying the issues and observations highlighted in the case study. We first describe the phenomenon of “output dimension collapse” in the translation from brain activity to the latent feature space. We show that the regression model trained on the clustered outputs becomes specialized to the training examples and its output collapses into a subspace of the training dataset. Our simulations with clustered data demonstrate that out-of-sample prediction can be achieved with an increasing number of training clusters, suggesting that zero-shot prediction is possible if stimulus diversity and a compositional representation are available. We also explore the preservation of image information at hierarchical layers of DNNs and discuss the caveats associated with evaluating reconstructions using identification metrics alone. Finally, we provide speculation on how we are fooled by the seemingly realistic reconstructions generated by AI models.
We conclude with recommendations for critically testing reconstruction models. Our findings highlight the importance of rigorous evaluation, the use of diverse datasets, and the need for careful specification and characterization of model components to advance the field of visual image reconstruction and develop truly generalizable prediction capabilities.
2 Results
2.1 Case study
We investigated two types of recent generative AI-based reconstruction methods, StableDiffusionReconstruction (Takagi and Nishimoto, 2023b) and Brain-diffuser (Ozcelik and VanRullen, 2023), as well as their validation dataset, the NSD (Allen et al., 2022). Both reconstruction methods utilize CLIP features (Radford et al., 2021) to effectively apply recent text-to-image diffusion models in visual image reconstruction analysis. CLIP text features are obtained from the average of five text annotations corresponding to the stimulus image. These text annotation information are only used during training to map the brain activity into CLIP text features. In the test phase, they directly predict CLIP text features from brain activity during image perception. Henceforth, these two reconstruction methods will be referred to together as text-guided reconstruction methods.
The StableDiffusionReconstruction method (Takagi and Nishimoto, 2023b) uses the components of the Stable Diffusion model (Ramesh et al., 2021), the VAE features (Kingma and Welling, 2014) and CLIP text features as latent features. Similarly, the Brain-diffuser method (Ozcelik and VanRullen, 2023) utilizes the component of another type of diffusion model (Xu et al., 2022), CLIP text features, and CLIP vision features as well as VDVAE features (Child, 2021) as latent features. Both methods first generate initial images from translated VAE/VDVAE features. These initial images are then passed through the image-to-image pipeline of the diffusion model conditioned on the translated CLIP features, producing the final reconstructed images. They validated the reconstruction performance using the NSD dataset, preparing training and test data based on the data split provided by the NSD study. Thanks to the authors of these studies who made efforts to make their datasets and scripts publicly available, we were able to conduct our replication analysis effectively. We compared the reconstructed results of these two text-guided reconstruction methods with those from a previous image reconstruction method, iCNN (Shen et al., 2019b). For more information on datasets and reconstruction methods, see Material and Methods (“Datasets” or “Reconstruction methods”) or the original studies (Allen et al., 2022; Shen et al., 2019b; Takagi and Nishimoto, 2023b; Ozcelik and VanRullen, 2023).
2.1.1 Observations: Failed replication and convincing reconstruction from random data
We first confirmed the reproduction of the findings of the text-guided reconstruction methods, particularly those of the Brain-diffuser study (Ozcelik and VanRullen, 2023) when applied to the NSD dataset (Figure 2a). The reconstructed images generated by the Brain-diffuser method effectively captured most of the layout and semantics of the test images. Although the reconstructed images produced by the StableDiffusionReconstruction method (Takagi and Nishimoto, 2023b) showed slightly degraded performance compared to the original paper, they still successfully captured the semantics of the test images. Notably, the iCNN method (Shen et al., 2019b) also performed well on the NSD dataset, with reconstructed images capturing the dominant structures of the objects within the images, consistent with the findings reported in the original study.
To further investigate the generalizability of the text-guided reconstruction methods, we attempted to replicate their performance using a different dataset, Deeprecon, which was originally collected for the study by Shen et al. (2019b). The Deeprecon dataset was specifically designed to avoid overlap between training and test sets, making it an ideal benchmark for evaluating the zero-shot prediction capabilities of reconstruction methods. However, the original Deeprecon dataset lacked the text annotations required by the text-guided reconstruction methods. To facilitate a fair comparison, we collected five text annotations for each training stimulus in the Deeprecon dataset through crowd-sourcing and used them to generate CLIP text features.
Despite the inclusion of these text annotations, the text-guided reconstruction methods failed to achieve the same level of performance on the Deeprecon dataset as they did on the NSD dataset. The reconstructed images produced by the text-guided methods exhibited photo-like appearances but suffered from largely degraded quality compared to their performance on the NSD dataset (Figure 2b). Notably, the text-guided reconstruction methods generated photo-like reconstructions even for simple geometric shapes present in the Deeprecon dataset, which deviated significantly from the original stimuli. In contrast, the iCNN method (Shen et al., 2019a) consistently provided faithful reconstructions for both the NSD and Deeprecon datasets, despite its simplicity compared to the text-guided methods.
Figure 2: Image reconstruction results. (a) Reconstruction results for the NSD. The first row shows the original test images, followed by reconstructions using StableDiffusionReconstruction, Brain-Diffuser, and iCNN methods. (b) Reconstruction results for the Deeprecon dataset. The format is the same as in (a).
Upon further investigation, we noted a questionable post-hoc image selection procedure. In the StableDiffusionReconstruction study (Takagi and Nishimoto, 2023b), they presented the reconstruction results by the following procedure: “We generated five images for each test image and selected the generated images with highest PSM” (PSM means perceptual similarity metrics), as illustrated in Figure 3a. This selection might lead readers or peer reviewers, particularly those not specialized in the brain decoding field, to overestimate the effectiveness of the methods and potentially lead to a distorted understanding of the actual reconstruction performance. Note that Brain-diffuser study (Ozcelik and VanRullen, 2023) did not executed such procedures and in their following report (Takagi and Nishimoto, 2023a), they updated the image presentation procedure more fairly as: “we generated five images with different stochastic noise and selected three images randomly.”
Figure 3: Questionable post-hoc selection. (a) Procedure in the StableDiffusionReconstruction method. Five images were generated, and the one that most closely resembled the test images was selected. (b) Reconstruction using shuffled brain activity. The left panels show the images generated from the VAE features given shuffled brain activity and CLIP text features given original brain activity. The right panels show the images generated from the VAE and CLIP text features given shuffled brain activity.
To examine the impact of this post-hoc selection procedure, we conducted an experiment using random brain data (Figure 3b). Instead of feeding the test brain data into pretrained translators, we shuffled the activities within each voxel of the NSD test set independently to create random brain data. Surprisingly, when the random brain data were input into the VAE feature translator, which contributes to producing initial images, plausible images were obtained by generating five images and selecting the best one. Even more strikingly, when the random brain data were input into both VAE and CLIP text feature translators, we still obtained convincing results by simply generating images five times and conducting the aforementioned selection. These observations are inexplicable since the artificially created brain data should completely lack any information related to the original visual stimuli.
These observations raise perplexing questions about the performance and generalizability of recent text-guided reconstruction methods. The significant deterioration in performance when simply changing the dataset from NSD to Deeprecon highlights the need for a deeper understanding of the factors contributing to the success of these methods on the NSD dataset. Moreover, the ability to obtain plausible reconstructions from random brain data by merely generating multiple images and selecting the best ones suggests that there may be fundamental issues with both the evaluation dataset and the components of the reconstruction methods themselves. In the following sections, we will conduct a thorough investigation into the potential problems associated with the NSD dataset and each component of the text-guided reconstruction pipeline.
2.1.2 Lack of diversity in the stimulus set
First, we examined the characteristics and limitations of the NSD dataset itself. To characterize the diversity of stimuli in the datasets and their latent representations, we focused on the CLIP features, which are used as latent features in recent reconstruction methods. We employed uniform manifold approximation and projection (UMAP) (McInnes et al., 2018) to visualize the CLIP text features of the NSD stimuli (see Materials and Methods “UMAP visualization”). The visualization revealed approximately 40 distinct clusters, with considerable overlap between the training and test sets (Figure 4a). Interestingly, we were able to describe the stimulus images in each cluster using a single semantic word, such as airplane, giraffe, or tennis. Despite the NSD containing around 30,000 brain samples per subject, the diversity of the presented stimuli was quite limited to just around 40 semantic concepts. In contrast, the Deeprecon dataset, which was specifically designed to differentiate object categories between training and test data, exhibited less overlap between the two sets (Supplementary Figure S1).
To further investigate the similarity between the training and test stimuli, we conducted an analysis using a state-of-the-art perceptual similarity metric called DreamSim (Fu et al., 2023) (Figure 4b). For each test image in the NSD, we used DreamSim to identify the most similar training images based on their perceptual similarity scores. We observed that the training images extracted using DreamSim were highly similar to the test images in the NSD, not only in terms of semantic concepts but also in terms of overall layout and visual composition. This finding suggests that the NSD test images are heavily biased towards the distribution of the training data, with a significant overlap in the visual and semantic features present in both sets. Such a strong bias raises concerns about the true reconstruction performance of methods evaluated on this dataset, as the impressive results may be largely attributable to the methods’ ability to memorize and reproduce the characteristics of the training data rather than generalizing to novel stimuli.
In contrast, when we applied the same DreamSim-based analysis to the Deeprecon dataset, we found that the extracted training images were substantially different from the test images. This observation indicates a clearer separation between the training and test distributions in the Deeprecon dataset, with the test images containing novel visual and semantic features that are not well-represented in the training set. The distinct differences between the NSD and Deeprecon datasets in terms of stimulus similarity highlight the importance of carefully designing evaluation benchmarks that can effectively assess the generalization capabilities of visual image reconstruction methods.
Figure 4: Dataset analysis. (a) UMAP visualization of CLIP text features in NSD. The center figure shows the scatter plot of the UMAP embedding of CLIP text features in NSD. Green points represent training samples, while orange points represent test samples. The surrounding images show the images in cluster. The presented images in each cluster were randomly selected. (b) Similarity between training and test images using the DreamSim metric. The left side shows results for the NSD, and the right side shows results for the Deeprecon dataset. The top 5 training images with highest similarity to each test image were selected.
2.1.3 Failed zero-shot prediction in the feature space
Given the significant semantic and structural similarities between the NSD test images and the training images, it is crucial to assess whether the predictions made from brain activity truly demonstrate generalization beyond the training set (i.e., zero-shot prediction). To evaluate the zero-shot prediction capability of the translators used in the reconstruction methods, we conducted an $N+1$ way identification analysis, where $N$ represents all of the training samples and $1$ refers to the test sample (Figure 5a). We calculated the percentage of translated features that successfully identified the corresponding test sample in the NSD dataset.
The performance of the CLIP features was nearly $0\%$ (Figure 5b). This extremely low performance suggests that the CLIP feature translator fails to achieve zero-shot prediction, even within the distribution covered by the training set. In other words, the CLIP features translated from brain activity are unable to accurately identify novel test samples, indicating a lack of generalizability beyond the training data. In contrast, the VGG19 features (Simonyan and Zisserman, 2015) employed in the iCNN method demonstrated moderate identification performance at the latent features on the intermediate layers. This finding suggests that the VGG19 features possess a certain degree of generalizability, enabling them to identify test samples that were not encountered during training.
Figure 5: Zero-shot sample identification. (a) Schematic diagram of novel sample identification: calculating the similarity between predicted features and test features or each of the training features, and counting how many predicted features were more similar to the true features than other candidate training features. (b) Results for NSD. The left side shows the identification accuracies evaluated on the CLIP features, which were used in the Brain-diffuser methods. The right side shows the accuracies evaluated on several VGG19 features used in the iCNN method.
Furthermore, we investigated whether the CLIP feature translator enables the prediction of novel semantic concepts not in the training. We redesigned the dataset split to ensure no semantic concepts were shared between them as in the previous zero-shot prediction study (Mitchell et al., 2008; Brouwer and Heeger, 2009). We first applied $k$-means clustering to the UMAP embedding space of the NSD’s CLIP text features (Supplementary Figure S2a). We set the number of clusters as $40$ based on visual inspection of the UMAP results. Based on these clustering results, we performed a hold-out analysis: when predicting samples within a cluster (e.g., the ski cluster), we excluded samples from that cluster in the training set (Figure 6a; hold-out split conditon). As a control, we also prepared a naive data-split condition where the training sample size is the same as in the hold-out split condition but allows overlapping semantic concepts. When we visualized the properties of predicted features in hold-out analysis by transforming the predicted features into the previous UMAP embedding space (Figure 4a), we observed that the predicted features tend to diverge largely from their original clusters and move into other clusters (Figure 6b).
Figure 6: Zero-shot cluster identification. (a) Conceptual diagram of hold-out analysis. For predictions of test samples in a cluster, training samples belonging to the cluster were excluded. The naive split condition serves as a control, using the same number of training samples as in the hold-out split condition but allowing semantic concepts to overlap between training and test sets. (b) Representative example of feature translation under the hold-out split condition. The green and orange points represent training and test samples, respectively. The purple diamonds represent the translated (predicted) latent features. The gray lines represent the transition from the test sample to the corresponding predicted feature. (c) Two types of evaluation of predicted latent features. The left side shows the cluster identification accuracy, which compares predicted features to the average features of training samples within each cluster. The right side shows the pairwise sample identification accuracy, comparing predicted features with the test sample feature or with features of other test samples. (d) Feature prediction results. The four surrounding panels show the semantic clusters observed in the UMAP. In each panel, the upper graphs show the cluster identification performances for CLIP text and CLIP vision features. The expected chance level is $1/40$. The lower graphs show the pairwise sample identification performances for CLIP text and CLIP vision features. The expected chance level is $1/2$.
To quantitatively assess the performance of the feature translator, we employed two metrics: cluster identification accuracy and pairwise sample identification accuracy. Cluster identification accuracy focuses on evaluating the translator’s ability to generate features that correctly identify the semantic cluster to which a test sample belongs (Figure 6c; left). In this analysis, we compare the predicted features of each test sample to the average features of the training samples within each semantic cluster. The accuracy is then calculated as the percentage of predicted features that successfully identify the original semantic cluster of their corresponding test samples in the feature space. Pairwise sample identification accuracy is a commonly used metric in the evaluation of feature prediction and reconstruction performance (Beliy et al., 2019; Shen et al., 2019b, a; Mozafari et al., 2020; Qiao et al., 2020; Ren et al., 2021; Gaziv et al., 2022; Takagi and Nishimoto, 2023b; Ozcelik and VanRullen, 2023; Ferrante et al., 2023; Scotti et al., 2023; Denk et al., 2023; Koide-Majima et al., 2024). This analysis assesses the translator’s ability to generate features that are more similar to the actual test features than to the features of other samples in the test set (Figure 6c; right). For each test sample, we compare its predicted features with two sets of features: the actual test features and the features of another randomly selected sample from the test set. The accuracy is then calculated as the average winning rate of the predicted features against all other test samples, indicating the proportion of cases where the predicted features are more similar to the actual test features than to the features of other samples.
Figure 6d presents a comparison of the identification performance of the feature translator in each semantic cluster between the hold-out and naive split conditions, as measured by two evaluation metrics: cluster identification accuracy and pairwise sample identification accuracy. The cluster identification accuracy in the hold-out split condition exhibited a substantial drop compared to the naive split condition across all semantic clusters. Notably, the cluster identification accuracy was frequently found to be $0\%$ in the hold-out split condition, exposing a severe limitation of the translator in novel semantic concepts that were absent from the training set (see also Supplementary Figure S2b). This finding strongly suggests that the CLIP feature translator functions primarily as a “classifier,” heavily relying on the semantic features learned during the training phase.
Intriguingly, despite the feature translator’s complete failure to identify semantic concepts in the hold-out split condition, the pairwise sample identification accuracy often surpassed the chance level across the semantic clusters (Figure 6d). The insensitivity of the pairwise sample identification accuracy to the translator’s inability to capture semantic concepts can be attributed to its emphasis on comparing the relative similarity between the predicted features and the actual test features, without taking into account the broader categorical context. This can lead to an inflated pairwise sample identification accuracy, potentially misrepresenting the true reconstruction performance and the translator’s ability to generalize to novel clusters. This observation raises concerns about the adequacy of using identification accuracy as the sole metric for evaluating reconstruction performance, a practice that has been widely adopted in many studies (Beliy et al., 2019; Shen et al., 2019a, b; Mozafari et al., 2020; Qiao et al., 2020; Ren et al., 2021; Gaziv et al., 2022; Takagi and Nishimoto, 2023b; Ozcelik and VanRullen, 2023; Ferrante et al., 2023; Scotti et al., 2023; Denk et al., 2023; Koide-Majima et al., 2024).
2.1.4 Failed recovery of a stimulus from its latent features
Finally, we conducted a rigorous evaluation of the generator component, which typically consists of diffusion models in generative AI-based reconstruction methods. To ensure that a visual image reconstruction method has the potential to faithfully reproduce an individual’s perceived visual experiences, it is crucial that the method can recover the original images with a high degree of perceptual similarity when the neural translation from brain activity to latent features is perfect. However, it has been unclear whether generative AI-based reconstruction methods meet this fundamental requirement. To address this question, we performed a recovery check analysis by reconstructing images using the true latent features of target images. Instead of using latent features transformed from brain activity, we directly input the latent features derived from the target images into the generator. While generative AI-based reconstruction methods generated images that were semantically similar to the target, the resulting images were not perceptually similar to the original (Figure 7a). In contrast, iCNN method yielded results that closely resembled the actual target images, demonstrating their superior ability to capture and replicate the original visual content.
Figure 7: Recovery check. (a) Reconstruction from the true latent features. The left column shows the true images, followed by each method’s recovery results. (b) Reconstruction using each component of the Brain-diffuser method. The left column shows the true images, followed by the recovery results of each component of the Brain-diffuser method.
To further investigate the recovery performance, we conducted a recovery check on each latent feature of the Brain-diffuser method (Figure 7b). Interestingly, reconstructions from VDVAE features, which are used for generating initial images in the Brain-diffuser, exhibited a high degree of similarity to the target images. However, the images generated by CLIP features through the diffusion models showed significant deviations from the original targets. These findings suggest that text-guided reconstruction methods may not be well-suited for visual image reconstruction tasks, as they fail to faithfully recover the original visual images. Instead, they tend to create images based on their latent features, such as CLIP features, which can lead to a phenomenon known as “hallucination” in the field of generative AIs. Hallucination refers to an output that appears plausible but is actually incorrect or misleading, raising concerns about the reliability and accuracy of the model (Rawte et al., 2023). These methods seem to prioritize generating semantically similar images rather than faithfully reconstructing the visual content perceived by the individual.
These findings may provide an explanation for why the text-guided reconstruction methods performed well only on the NSD dataset. The results of the case study demonstrated that the text-guided reconstruction methods struggle to reconstruct (Figure 2) or identify (Figure 6) test samples that lie beyond the distribution of the training set. Such limitations suggest that these methods lack true generalization capabilities and are unable to accurately reconstruct novel visual experiences that differ significantly from the examples they were trained on. Moreover, even when the test samples belonged to the same distribution as the training set, the translators of the text-guided reconstruction methods had difficulty correctly identifying those test samples (Figure 5). This observation indicates that the translators may not have learned a sufficiently robust and generalizable mapping between brain activity patterns and the corresponding latent features, further limiting their ability to faithfully reconstruct the perceived visual experiences.
The case study revealed that the NSD test stimuli are highly similar to the training set (Figure 4), with a significant overlap in their visual and semantic features. Given this similarity, the impressive reconstruction results achieved by the recent text-guided reconstruction methods on the NSD dataset should not be interpreted as evidence of zero-shot reconstruction capabilities. Instead, a more plausible interpretation is that these methods primarily function as a combination of “classification” and “hallucination.”
In this context, the translator component of the text-guided reconstruction methods functions primarily as a classifier, predicting categorical information present in the training data rather than capturing the fine-grained details of the visual experience. This limitation may explain why convincing reconstructions can be obtained even from random brain data through post-hoc selection. Due to the limited variety in the outputs generated by the reconstruction models, repeated trials can eventually produce images that are semantically and visually similar to the target stimulus. The apparent plausibility and semantic consistency of these generated images can be attributed to the capabilities of the diffusion model, which learns to generate realistic-looking images based on the semantic information. While these images may seem convincing at first glance, they do not accurately reflect the specific visual experience of the individual. This phenomenon of hallucination raises serious concerns about the reliability and validity of the text-guided reconstruction methods when evaluated on the NSD dataset.
2.2 Formal analysis and simulation
In the above case study, we have identified several issues with these text-guided reconstruction methods and the dataset, especially the cluster structure of CLIP latent features, the lack of diversity in NSD, and the misspecification of the latent representation for image reconstruction, which results in the inability of diffusion models to faithfully recover the original images from their latent features. However, it is crucial to recognize that the findings of the case study are not merely specific to CLIP, NSD, or diffusion models. Instead, these issues likely reflect more fundamental problems that can arise in the development and evaluation of brain decoding and visual image reconstruction methods. Thus in this section, we extend the problems identified in the case study into formal analyses and simulations in generalized settings, aiming to provide a more comprehensive understanding of the factors that contribute to the limitations of current reconstruction methods and explore strategies for mitigating these issues.
2.2.1 Output dimension collapse
First, we revisit the properties of the translator and highlight the issues that arise when the prediction targets have insufficient diversity. Let us consider the general situation of predicting a feature $\mathbf{y}\in\mathbb{R}^{D}$ from a brain activity pattern $\mathbf{x}\in\mathbb{R}^{D}$ using a linear regression model. For the training set, we consider brain activity matrix $X_{\mathrm{tr}}\in\mathbb{R}^{N\times D}$ and feature value matrix $Y_{\mathrm{tr}}\in\mathbb{R}^{N\times D}$, where $X_{\mathrm{tr}}$ consists of $N$ samples of $D$-dimensional brain activity vectors $\mathbf{x}$ and $Y_{\mathrm{tr}}$ consists of $N$ feature value vectors $\mathbf{y}$. We then train the linear (ridge) regression model using this training data. Given a regularization parameter $\lambda$, the weight of ridge regression model is analytically derived as $W=(X_{\mathrm{tr}}^{\top}X_{\mathrm{tr}}+\lambda I)^{-1}X_{\mathrm{tr}}^{\top}Y_{\mathrm{tr}}$, where $I$ is the $D\times D$ identity matrix. The predicted feature value $\hat{\mathbf{y}}_{\mathrm{te}}$ for the test brain activity data $\mathbf{x}_{\mathrm{te}}$ can be represented as:
$\hat{\mathbf{y}}_{\mathrm{te}}$ $=W^{\top}\mathbf{x}_{\mathrm{te}}$ (1)
$=Y_{\mathrm{tr}}^{\top}X_{\mathrm{tr}}(X_{\mathrm{tr}}^{\top}X_{\mathrm{tr}}+\lambda I)^{-1}\mathbf{x}_{\mathrm{te}}$ (2)
$=Y_{\mathrm{tr}}^{\top}\mathbf{m}=\sum_{i}^{N}m_{i}\mathbf{y}_{\mathrm{tr}}^{(i)},$ (3)
where ${\bf m}=X_{\rm tr}(X_{\rm tr}^{\top}X_{\mathrm{tr}}+\lambda I)^{-1}\mathbf{x}_{\mathrm{te}}\in\mathbb{R}^{N\times 1}$ and $\mathbf{y}_{\mathrm{tr}}^{(i)}$ is the $i$th training feature vector. This transformation indicates that the predicted value is always represented as a linear combination of the features in the training set. This property is not limited to ridge regression but is generally applicable to ordinary ridgeless linear regression and related linear models.
Next, we consider a scenario where the diversity of the target features is small. This situation can arise when the feature space exhibits a clustered structure and the training data lacks sufficient diversity, as observed in the case study with the CLIP text features and the NSD dataset. When the training feature values have limited diversity, the predicted values from brain activity, which are represented as linear combinations of these features, also become constrained. Consequently, the prediction from brain data to features effectively becomes a projection onto a low-dimensional subspace formed by the training data.
To illustrate this phenomenon, we conducted a simple simulation to examine the distribution of predicted values from a linear regression model trained on clustered features (Figure 8). We generated clustered features by sampling from a Gaussian mixture distributions in a high-dimensional space. The corresponding brain activity data was generated from a multivariate Gaussian distribution. We then trained a linear regression model to predict the clustered feature values from the randomly generated brain data.
Figure 8: Output dimension collapse. The left panel shows the data distribution of source brain activity in the first two dimensions of high-dimensional space. The right panel shows the distribution of target latent features in the first two principal component (PC) dimensions. If the target features form a cluster structure in the high-dimensional space (left panel green points), the predictions (gray colors) fall into a subspace created by the training clusters, regardless of the input data (right panel).
The simulation results clearly demonstrate the impact of clustered features on the predicted values. The trained model projects arbitrary brain data onto the subspace defined by the training features, resulting in predicted values that are confined to the vicinity of the training clusters. This observation highlights the limitation of training linear regression models with clustered features: the trained model’s predictions are inherently constrained by the diversity and structure of the training data.
This phenomenon, which we refer to as “output dimension collapse,” has important implications for the generalization capability of linear regression models in the context of brain decoding and visual image reconstruction. When the training data lacks diversity and forms distinct clusters in the feature space, the translator overly adapts to the subspace formed by the training data, regardless of the potential of the latent feature space. Consequently, the translator’s outputs become similar to patterns in the training set, irrespective of the inputs, severely limiting the model’s ability to predict novel or out-of-distribution samples. Note that output dimension collapse is not inherently caused by the limitations of linear regression models but rather stems from a lack of diversity in the outputs of the training data. This phenomenon can occur regardless of the type of regression model employed, including nonlinear models. In fact, nonlinear models may be even more susceptible to output dimension collapse, as the increased flexibility and complexity of these models can potentially exacerbate the issue.
Output dimension collapse may explain why plausible reconstructions were obtained even from random brain data by merely generating images several times in the case study. The NSD’s lack of semantic diversity causes the translator to adapt only to the feature patterns of the training set, restricting its outputs to the subspace formed by the training data. As a result, convincing images could be found even from random data through questionable post-hoc selection.
2.2.2 Simulation with clustered features: What makes prediction compositional
The case study revealed that the NSD exhibits limited diversity (Figure 4) and poses difficulties for zero-shot prediction (Figure 5 and 6). These observations suggest that the translator of CLIP features suffers from output dimension collapse due to the lack of semantic diversity in NSD. To explore potential strategies for mitigating output dimension collapse and achieving flexible predictions, we conducted simulation analyses using clustered data to assess generalization performance beyond the training set.
Our simulation involved teacher-student learning with Gaussian mixture models (see Materials and Methods: “Simulation with clustered data”). To imitate the feature translation situation from brain activity, we first generated feature data and assumed a scenario where observation noise affects the brain data. To simulate a situation where the dataset has cluster structures and to control diversity effectively, the training feature $\mathbf{y}\in\mathbb{R}^{D}$ was generated from a $D$-dimensional mixture of Gaussians (Figure 9a). The brain activity data $\mathbf{x}\in\mathbb{R}^{D}$ was generated from $\mathbf{y}$ multiplied by teacher weight and adding observation noise.
Figure 9: Simulation analysis of predicting features with cluster structure. (a) Illustration of target latent features setting. The latent features were generated from Gaussian mixture distributions. $\sigma_{r}^{2}$ is a parameter of the sample variance of each cluster. $\sigma_{s}^{2}$ controls the distance across clusters. In-distribution test samples are generated from the same clusters as in the training set. Out-of-distribution (OOD) test samples are generated from a novel cluster not included in the training set. (b) Cluster identification accuracy. The $x$ axis represents the number of clusters used in the training and the $y$ axis represents the cluster identification performance. Green line shows the results of In-distribution test samples and orange line shows those for OOD samples. Dashed curve indicates the chance level. (c) Illustration of a possible scenario where sufficient data diversity is needed to achieve generalizability. The upper illustration shows that the training data should cover the whole latent feature space, requiring an exponential order relative to the feature dimension. The lower one shows that the training data is covering only the effective axes of the latent feature space, leading to a linear order relative to the feature dimension. (d) Relation between latent feature dimension and sufficient number of clusters for generalization. The $x$ axis represents the dimension of target latent features. The $y$ axis represents the number of clusters that achieve above $0.5$ cluster identification accuracy. Three lines indicate the results from different $\sigma_{s}^{2}/\sigma_{r}^{2}$. (e) Representative example of generalizability beyond the training set. Light green points indicate the in-distribution test samples while light orange points indicate OOD test samples. Green and orange points represent the predictions.
We trained a ridge regression model on large training data samples and obtained the student weight. To simulate the situation where the trained model encounters clusters that are not available at the training phase, we generated two types of test samples, in-distribution and out-of-distribution. In-distribution test samples were generated using the same probability distribution as the training data, while out-of-distribution (OOD) samples were generated from a different distribution. For these two types of predicted features, we calculated the cluster identification accuracy (Figure 6c), using $C+1$ cluster centers, which included $C$ centers from the training set and one additional center for the OOD test set.
We first examined zero-shot prediction performance as the number of clusters in the training data was increased while fixing the feature dimension $D$ and the cluster variance ratio between each cluster constant (Figure 9b). While the cluster identification performance of in-distribution samples were perfect, that of OOD samples showed different patterns depending on the number of clusters in the training data. When the number of training clusters was low, cluster identification accuracy was $0\%$. This indicates that the behavior of the translator became more similar to a classifier, difficult to generalize beyond the training set, as observed in NSD cases (see also Figure 6d). On the other hand, as the number of clusters in the training data increased, it became possible to identify the novel clusters, achieving the same performance as in-distribution test samples. This observation indicates the importance of the diversity of the training dataset. We also emphasize that large numbers of training data do not necessarily address the problem. All of these results were obtained with a sufficiently large number of training data, and the number of training clusters was varied while keeping the number of training data fixed. Also, we observed qualitatively similar results in increasing the data diversity by controlling the cluster variance ratio (Supplementary Figure S3a).
Next, we investigate how diverse the training data needs to be to ensure sufficient generalization. There are two possible scenarios for diversifying training samples: either by densely sampling the entire space so that there are no remaining gaps or by uniformly sampling to the extent that it covers the entire dimension of the feature space (Figure 9c). The former scenario requires an exponentially larger number of samples relative to the dimension while the latter scenario only requires up to a linear order. We sought to reveal which of these two scenarios was more likely to be true by varying the dimensions of the feature space and identifying the number of clusters required for generalization (see also Supplementary Figure S3b). Here, we defined the number of clusters required for generalization as the point at which the prediction accuracy of OOD samples exceeds $50\%$. The relationship between the dimension of the feature space and the number of clusters necessary for generalization appears to be linear (Figure 9d). This finding suggests that achieving generalization does not necessarily require an exponentially large diversity that fills the entire feature space. Instead, it suffices to have a number of clusters that cover the effective dimensions within the feature space. Although obtaining a large amount of brain data is hard work, it is important for a dataset to contain sufficiently diverse stimuli covering the effective dimensions of the target feature space to achieve zero-shot prediction.
We also confirmed this phenomenon by a simple and transparent example ($D=2$, Figure 9e). The training data covers sufficient axes in the feature space, enabling the prediction of locations not present in the training set. Based on this low dimensional intuition, we argue that successful zero-shot prediction requires that training data leads to representations that can serve as a basis spanning the feature space. Leveraging such bases effectively enables the model to predict novel samples by predicting each basis and combining them. This compositional representation, spanning the target feature space is crucial for zero-shot prediction (Schug et al., 2024) and reconstructing arbitrary visual images from limited brain data.
2.2.3 Preserved image information across hierarchical DNN layers
To enable the reconstruction for arbitrary visual images, compositional latent features that can be appropriately mapped into the image space are essential. Although it is often assumed that the hierarchical feature representations in DNNs, especially convolutional neural network (CNN), discard pixel-level information through their hierarchical processing with progressively expanding receptive fields, this is not necessarily the case.
For example, the latent features of auto-encoder models (Hinton and Salakhutdinov, 2006; Kingma and Welling, 2014; van den Oord et al., 2017) can represent images in a low-dimensional space while preserving their reversibility, which is reasonable considering that the output is trained to match the input. Furthermore, Mahendran and Vedaldi (2015) showed that input images can be reconstructed with reasonable accuracy even from relatively high-level layers of a DNN designed for an object recognition task. It has also been argued that large receptive field sizes do not necessarily impair neural coding capacity as long as the number and density of units remain constant (Zhang and Sejnowski, 1999; Majima et al., 2017). These results challenge the notion that higher-level layers in DNNs discard all pixel-level information and highlight the potential for utilizing intermediate representations in visual image reconstruction. To further illustrate this point, we performed a recovery check on each intermediate layer of the VGG19 network used in the iCNN methods (Figure 10; see also Figure 7). Given the DNN features of a target image, we optimized input pixel values to make the image’s latent features similar to the targets (See Materials and Methods “Recovery check of a single layer by iCNN”). We observed that input images can be recovered with reasonable accuracy from relatively high-level layers (around the 11th layers, the 19th layers as a total). Furthermore, by introducing image generator networks to add constraints on image statistics, reasonable recovery can be achieved from even higher layers. By utilizing a weak image prior (Ulyanov et al., 2017), which contains only information about the structure of images without any prior information on natural images, input images can be recovered from the 13th layer. When using an image generator that has learned natural image information (Shen et al., 2019b; Dosovitskiy and Brox, 2016), input images can be recovered even from the 15th layer.
Figure 10: Recovery check: single layer used in iCNN method. For three images, the different optimization algorithm is applied to recover the visual input. The left one is pixel optimization, which directly optimizes the pixel values so that the loss between that generated image feature and target image features is minimized. The middle one is optimization with week image prior, which optimizes the parameters of image prior models and their latent features simultaneously (see Material and methods). The right one is optimization with pretrained image prior, which optimizes the latent features of the fixed image generator model so that the loss between the output image features and target image features is minimized.
These findings underscore the significance of intermediate DNN representations that can effectively recover the input image in visual image reconstruction studies. Conversely, utilizing high-level image features or features from other modalities, such as text annotations, for which the corresponding images are challenging to recover, is not a rational choice for visual image reconstruction tasks. Although recent text-to-image models and decoded text features can easily generate images, the outputs should not be interpreted as reconstruction results. Instead, it is more reasonable to interpret them as visualizations of decoded semantic information.
2.2.4 Caveat with evaluation by identification
Figure 11: Pairwise sample identification with structured data. The $x$ axis represents the number of categories in the test set. The $y$ axis represents the pairwise sample identification accuracy. The lines show the different accuracy $q$ in the situations where the candidate sample is from a different category than those in the test samples. The representative similarity matrices corresponding to these conditions are shown in the figure above.
Pairwise identification has been a standard metric for evaluating latent feature decoding (Horikawa and Kamitani, 2017) or reconstruction performance (Beliy et al., 2019; Shen et al., 2019a, b; Mozafari et al., 2020; Qiao et al., 2020; Ren et al., 2021; Gaziv et al., 2022; Takagi and Nishimoto, 2023b; Ozcelik and VanRullen, 2023; Ferrante et al., 2023; Scotti et al., 2023; Denk et al., 2023; Koide-Majima et al., 2024). However, our analysis revealed that even with difficulties in accurately identifying semantic clusters the test samples belong to, pairwise identification performance still surpassed chance levels (Figure 6d). Here we critically examine this metric and demonstrate that significant results can be easily obtained when the target or predicted features exhibit certain structures.
Pairwise identification is calculated as the accuracy with which the predicted features (reconstructed image or its features) can correctly identify the true ones, in pairs consisting of a true sample and one of the remaining samples in the test set. If the candidate feature belongs to the same category as the true feature, we can expect the identification to be difficult. Conversely, if the candidate feature belongs to a different category from the true feature, identification would become easier.
Here we assume the test set comprises $k$ categories and all test samples are equally distributed across each category for simplicity. We model the situation mentioned above as the expected identification accuracy is set to chance level (i.e., 0.5) when a candidate belongs to the same category as the true feature. In contrast, the expected identification accuracy when a candidate belongs to a different category from the true feature is set to $q$, ranging between 0.5 to 1 to mimic the ease of identification across categories. Assuming the number of test samples is large enough, the pairwise identification becomes
$\mathrm{Acc}=\frac{1}{k}\cdot 0.5+\left(1-\frac{1}{k}\right)\cdot q$ (4)
(See Materials and Methods “Expected identification accuracy in imprecise reconstructions” for the derivation). Figure 11 shows the line plots of expected values of pairwise identification accuracy as a function of the number of categories in the test set, while presenting the possible similarity structure of the test set. Even with failure to identify within categories completely and success only between two categories, pairwise identification accuracy can still reach a high value up to $75\%$. This finding highlights a potential limitation of the pairwise identification metric, as judging reconstruction performance solely based on it can be misleading.
2.2.5 How are we fooled by hallucinations of generative AIs
Figure 12: The dangerous interpretation of photo-like appearances as accurately reflecting the perceived visual image. The left Venn diagram represents the interpretation of photo-like reconstructions. $T$ denotes the event where the generator’s output accurately reflects the visual image and $R$ denotes the event where the generator’s output has a photo-like appearance. Introducing generative AIs (diffusion models) might increase the overlap between $T$ and $R$; however, it may also significantly increase the areas unrelated to $T$. The right equation represents the probability that the generator’s output accurately reflects the visual image given that the generator’s output has a photo-like appearance as inferred from the left Venn diagram.
Generative AIs have recently made remarkable progress, with models now capable of producing high-resolution and hyper-realistic images from text input (Ramesh et al., 2021) or generating text of a quality indistinguishable from human-written content (Brown et al., 2020). However, due to the complex internal structure of these models and the vast amounts of data they are trained on, we are often fooled by the outputs of generative AIs. For instance, when searching for an unfamiliar topic using a large language model (LLM) in our daily life, we may not realize that the model is creating false concepts. Someone may have mistaken generative AI show real-time response in the promotion, despite the fact that these promotions were edited (Coldewey, 2023). We might also believe that these models are unbiased or can represent all possible viewpoints, even though they inherently contain biases from their training data and developers (Messeri and Crockett, 2024). As we have observed that the generative AI-based reconstruction methods exhibit photo-like appearances but poor generalizability (Figure 2), could similar issues be occurring in visual image reconstruction studies as well?
The goal of visual image reconstruction is to generate images from brain activity that precisely mirror visual experience. However, there appears to be a prevalent focus among the general public, reviewers, and even researchers on achieving outputs that are as photorealistic as possible, rather than emphasizing the accuracy of these reconstructions. This shift in focus raises questions about the extent to which these photorealistic reconstructions truly represent the actual visual experiences.
Traditionally, we have held two beliefs: 1) generating photo-like images from brain activity is challenging, and 2) if the reconstruction pipeline effectively captures the brain representation under natural image perception, the model’s output should also appear naturalistic or photo-like. Based on these beliefs, we tend to consider photo-like reconstructions as an indication of accurately reflecting the actual visual experience.
This heuristic can be formalized as follows:
Assumption 1: $\Pr[R]\ll\Pr[\bar{R}]\approx 1,$ (5)
Assumption 2: $\Pr[R\mid T]\approx 1,$ (6)
Conclusion: $\Pr[T\mid R]\text{ is high},$ (7)
where $T$ represents the event of the model’s output truthfully reflects the visual image and $R$ represents the event of the model’s output has a realistic appearance. $\bar{R}$ represents a complementary event of $R$. $\Pr[T\mid R]$ is the probability that how likely we should conclude that the model’s output actually reflects the subject’s visual experience based solely on the fact that the output is photo-like. In fact, this heuristic is reasonable when the above two assumptions are true and $\Pr[T]$ is not extremely low.
However, recent developments in generative AIs, such as diffusion models, have made it easily to produce convincing photo-like outputs, subverting the first assumption i.e, $\Pr[R]\gg\Pr[\bar{R}]$. Consequently, it becomes invalid to infer that the visual images are accurately reflected in the generator outputs solely because they appear photo-like. Rather, as shown in Figure 12, the probability $\Pr[T\mid R]$ may become smaller as the generative AIs produce more convincing outputs (see also Figure 7). This perspective emphasizes the need for careful evaluation of reconstruction performance, considering the possibility of hallucinations by generators. While pursuing photo-like reconstructions to improve reconstruction fidelity is undoubtedly important, it would be counterproductive to obsess over naturalistic appearance to the point of neglecting the original goal of reconstructing perceived visual images.
3 Discussion
In this study, we critically examined recent generative AI-based visual image reconstruction methods to assess their true capabilities and limitations. Our primary goals were to (1) investigate the performance of these methods on different datasets, (2) identify potential issues and pitfalls in their methodology and evaluation, and (3) provide insights and recommendations for future research in this field. We conducted a case study focusing on text-guided reconstruction methods and their validation on the Natural Scene Dataset (NSD). Our findings revealed several concerns, including the failure to replicate the reconstruction performance on a different dataset, the use of problematic post-hoc image selection procedures, the lack of diversity and limited number of clusters in the NSD stimulus set, the failure of zero-shot prediction by the translator component, and the inability to accurately recover original stimuli by the generator component. Formal analysis and simulations further demonstrated the phenomenon of output dimension collapse, the importance of compositional representations for achieving zero-shot prediction, and the potential pitfalls of relying solely on identification metrics to evaluate reconstruction performance. Moreover, we highlighted that photo-like appearance does not necessarily imply accurate reflection of the perceived visual images. Based on these findings, we argue that the reconstructions from the recent text-guided reconstruction methods are, in large part, the result of a combination of classification and hallucination. Our study emphasizes the need for more rigorous evaluation and careful interpretation of results in visual image reconstruction research, particularly when using generative AI-based methods.
While our study critically examined the limitations of recent text-guided, diffusion-based methods for visual image reconstruction, it is important to acknowledge that these approaches provide new and potentially promising directions for brain decoding research. Although we emphasized the importance of achieving zero-shot prediction, it is crucial to recognize that most brain decoding studies focus on classification tasks, which, while not zero-shot, have provided insights into neural representations (Kamitani and Tong, 2005; Haxby et al., 2001). Moreover, the visualization of semantic contents (e.g., the supplementary movies in Horikawa et al. (2013)), can have significant utility in visual communication, even if it does not constitute zero-shot prediction. It is also worth noting that the individual components of these text-guided, diffusion-based methods are already being utilized in various brain decoding applications. For instance, text latent features derived from deep neural networks and large language models (LLMs) have shown promise in analyzing semantic information from brain activity (Tang et al., 2023; Caucheteux et al., 2023; Zhou et al., 2024). Furthermore, diffusion models are not limited to text-to-image generation; they can be trained to generate images from visual latent features, as demonstrated by Cheng et al. (2023), enabling the successful reconstruction of subjective experiences. By leveraging these components and exploring their potential synergies, researchers can continue to push the boundaries of brain decoding and visual image reconstruction, while being mindful of the challenges and limitations highlighted in our study.
The recent trend of collecting and sharing large-scale visual neural datasets, such as those by Hebart et al. (2023) and Xu et al. (2024), is a welcome development in the field of neuroscience. These datasets provide valuable resources for researchers to investigate brain function and advance our understanding of visual processing. The NSD is a particularly notable example, as it was created with the goal of extensively sampling brain responses to a wide range of natural visual stimuli (Allen et al., 2022; Naselaris et al., 2021). The NSD has been widely utilized in various studies (Prince et al., 2022; Conwell et al., 2023; Gifford et al., 2023), demonstrating its value to the research community. While our results suggest that the semantic and visual diversity of the NSD stimuli may not be as high as initially thought, and there is substantial overlap between the training and test sets provided by the NSD authors, this does not diminish the overall importance and usefulness of the dataset. However, to fully leverage the NSD and other publicly available large-scale datasets for developing generalizable and zero-shot prediction models, it is crucial to carefully consider the data split between training and test sets. When aiming for generalizable predictions, such as in visual image reconstruction, researchers should verify whether there are significantly similar stimuli included in both the training and test sets. If necessary, redesigning the training and test split can ensure that the models are tested on their ability to predict novel stimuli. Moreover, recent advancements in functional alignment and inter-site neural code conversion methods (Haxby et al., 2011; Yamada et al., 2015; Wang et al., 2024) hold promise for combining datasets from different sources, enabling truly larger-scale data analysis in neuroscience. These techniques allow researchers to align brain activity patterns across individuals and measurement sites even when stimuli are not shared across datasets. By leveraging these methods, researchers can pool data from various sources, increasing the sample size and diversity of the combined dataset, mitigating the limitations of individual datasets, and enhancing the development of generalizable and zero-shot prediction models.
Investigating neural responses to natural stimuli is a highly valuable approach to understand brain function and representation (Nastase et al., 2020; Hasson et al., 2020). As our brains have developed while being exposed to natural scenes, it is crucial to use natural stimuli especially in model training. However, we should not forget that we are also capable of perceiving non-natural stimuli like artificial images. We would like to emphasize that there are the potential pitfalls when relying too heavily on evaluations based solely on natural stimuli. With the increasing scale of neural data and the growing complexity of analysis pipelines, there is a risk that the learned mappings may produce unexpected shortcuts, just as we have demonstrated that the generative AI-based reconstruction methods exploited the semantic and visual overlap between training and test sets. Drawing inspiration from comparative and developmental psychology, where researchers often employ simple stimuli and tasks to measure cognitive abilities in infants for better experimental control and precise inferences (Kominsky et al., 2022; Messeri and Crockett, 2024) , the evaluation of visual image reconstruction should not be limited to complex natural stimuli alone. While natural stimuli are essential for ensuring ecological validity and understanding how the brain processes real-world information, it is equally important to assess the performance in controllable and transparent manners.
We discussed output dimension collapse and related issues with clustered data found in natural image features. It is important to note that these issues can arise in various situations. Several studies have used brain data where the stimuli shared the category information between the training and test sets. Kavasidis et al. (2017) collected a dataset of EEG signals recorded during natural image perception. Their visual stimuli consisted of 2,000 images selected from 40 object categories in ImageNet (50 images per category) and the test set contains the same categories that are included in the training set. Denk et al. (2023) attempted to develop music reconstruction methods from fMRI activity patterns. Their music stimuli consisted of 540 music pieces selected from 10 music genres and the test set contains the same genres as the training set (Nakai et al., 2021). Orima et al. (2024) attempted to reconstruct perceived texture images from EEG signals. Their texture stimuli consisted of 191 images of 21 natural textures and they performed a reconstruction analysis with a leave-one-out manner. It should be carefully examined whether these studies may suffer from output dimension collaplse, merely decoding category information from brain activity and generating stimuli based on this classified category information, similar to our interpretation of the text-guided reconstruction methods. Note that for other issues in the data from Kavasidis et al. (2017), see Li et al. (2018); Xu et al. (2024). Our inspection of the data from Nishimoto et al. (2011) revealed that many frames in the test movie stimuli were almost identical to those in the training set (Supplementary Figure S4). This is presumably because temporally adjacent video frames were split between the training and test stimuli. While our preliminary analysis of their motion energy features did not exhibit peculiar clustering (note also that their model was an encoding model where the output is brain activity), caution should be exercised if the dataset is used for models that extract other features of the videos.
One of the remaining challenges in visual image reconstruction is the development of metrics for evaluating the quality and accuracy of the reconstructed images. The first and most critical step in assessing reconstruction results is to confirm a qualitative similarity between the reconstructed images and the perceived images through visual inspection across a diverse range of test sets. Following this, quantitative metrics should be employed for a more objective, high-throughput evaluation. However, as our analysis has suggested, it can be misleading to evaluate reconstruction by heavily relying on identification performance based on the relative similarity among alternatives (Koide-Majima et al., 2024). Even in cases where the reconstructed images only capture superficial information, such as categories or overall brightness, identification metrics can still be high. While identification performance can provide a useful benchmark, it should not be the sole metric for evaluating reconstruction quality. It is crucial to develop more appropriate similarity metrics that can accurately measure the perceptual similarity between the reconstructed and original images. One promising approach is to leverage image quality assessment (IQA) techniques from the computer vision field (Fu et al., 2023; Ding et al., 2020). These techniques are designed to quantify the perceptual quality of images and can be adapted to the specific requirements of visual image reconstruction.
Visual image reconstruction methods have gained attention not only from neuroscientists but also from the general public and policymakers, sparking discussions about their potential applications and risks (UNESCO, 2023). These stakeholders often contemplate the possibilities of seamless information communication through the brain, such as in brain-machine interfaces (BMIs), or the dangers of unauthorized access to private information from brain activity. This interest may stem from the perception that brain activity data can be obtained easily and reliably in real-time. However, current technology and analysis methods fall short of these expectations. For instance, most reconstruction methods analyze previously acquired brain data offline. Additionally, the brain data used for reconstructing images are often averaged over multiple presentations of the test image, with only a few studies demonstrating single-trial reconstruction results (Miyawaki et al., 2008; Cheng et al., 2023). Further, It has been argued that subject cooperation is essential for reliably training and testing decoding models (Tang et al., 2023). Public expectations are thus set too high, and it is challenging to meet these demands quickly. It is essential to explicitly state these limitations to avoid disappointment and prevent governments or companies from making misguided decisions. Key limitations include the lack of real-time analysis, reliance on averaged brain data, the necessity of subject cooperation, and the need for further model fine-tuning using the test brain dataset (Beliy et al., 2019; Chen et al., 2023a). It is crucial to avoid making overly optimistic claims about the ability to reconstruct arbitrary images, as the applicability can be highly dependent on specific training data and conditions.
4 Materials and Methods
4.1 Datasets
We utilized two datasets: the natural scene dataset (NSD) (Allen et al., 2022) and Deeprecon dataset (Shen et al., 2019b). Both datasets comprise visual stimuli and corresponding fMRI activity collected when subjects perceived the stimuli. In the NSD dataset, eight subjects were presented with MSCOCO images (Lin et al., 2014), yielding 30,000 brain activity samples per subject, which is three times the amount provided by the Deeprecon dataset. The Deeprecon dataset includes fMRI activity data from subjects presented with both ImageNet images (Deng et al., 2009) and artificial images. It contains roughly 8,000 brain samples per subject. Since this dataset is designed for evaluating reconstruction performance, the test data is carefully selected. The test natural images were selected from ImageNet, which were in categories different from those used in the training. The artificial images were only used as test data to check the generalizability performance of the proposed reconstruction methods. In both datasets, we adopted the train/test split used in previous studies and utilized data from the first subject (S1 in NSD, Subject 1 in Deeprecon). Text-guided reconstruction methods require text annotations of images.. For the NSD, text annotations accompanying the MSCOCO database were used. For the Deeprecon dataset, we collected captions for each experimental stimulus via crowd workers on Amazon Mechanical Turk, yielding five captions per image. The captions of training stimuli are publicly available at the GitHub repository.
4.2 Reconstruction methods
We utilized three image reconstruction methods: StableDiffusionReconstruction (Takagi and Nishimoto, 2023b), Brain-diffuser (Ozcelik and VanRullen, 2023) and iCNN (Shen et al., 2019b). Each method employs two common steps: first, translating brain activity patterns into latent features of the stimuli, and second, generating images from these latent features using an image generator (see also Figure 1). In the StableDiffusionReconstruction method, the latent features are the VAE (Kingma and Welling, 2014) features calculated from stimulus images and CLIP text features (Radford et al., 2021) from the image annotations. The generator is StableDiffusion (Rombach et al., 2022). They first generate low-resolution images from the translated VAE features, and those images are further fed into the StableDiffusion model with translated text features to generate images. The generated images are regarded as reconstructed images from brain activity. In the Brain-diffuser method, the latent features are the VDVAE (Child, 2021) and CLIP vision features from stimulus images and CLIP text features from the image annotations. The generator is Versatile diffusion (Xu et al., 2022). Similar to StableDiffusionReconstruction, low-resolution images are first generated from the translated VDVAE features, and these images are further used for the input of the versatile diffusion model with the translated vision and text features. The generated images are regarded as reconstructed images from brain activity. In the iCNN method, the latent features are the intermediate output of VGG19 (Simonyan and Zisserman, 2015) layer from stimulus images. As a generator, they used the pretrained image generator (Dosovitskiy and Brox, 2016) and they solved the optimization problem to minimize the discrepancy between the VGG19 features calculated from the generated images and the transformed VGG19 features. Well-optimized images are regarded as reconstructed images.
4.3 UMAP visualization
To investigate dataset diversity, we employed Uniform Manifold Approximation and Projection (UMAP), a non-linear dimensionality reduction technique (McInnes et al., 2018) to learn a projection from a latent features space to a lower dimension (UMAP embedding space). We used both the training and test CLIP text features for learning the UMAP projection. These features were combined and standardized beforehand. The UMAP hyperparameters follow the official UMAP guide for clustering usage with cosine distance as a distance metric. The learned UMAP was also used to project the features predicted from brain activity (Figure 4). After standardizing the predicted features using the same mean and standard deviation parameters used in UMAP projection learning, we projected the decoded features into the UMAP embedding space.
4.4 Simulation with clustered data
We conducted a simulation analysis to examine the generalization performance beyond the training data (Figure 6). This analysis involves a teacher-student learning task using Gaussian mixture distributions. We generated both input and target pairs from teachers, then examined the prediction performance of regression models of students. To imitate the feature translation situation from brain activity, we first generated feature data and assumed a scenario where observation noise affects the brain data.
The training sample of latent feature data $\mathbf{y}\in\mathbb{R}^{D}$ was generated from a $D$-dimensional space using Gaussian mixture distributions, formulated as:
$p_{\mathrm{tr}}(\mathbf{y})=\frac{1}{C}\sum_{c=1}^{C}\mathcal{N}(\boldsymbol{\mu}_{c}^{\mathrm{tr}},\sigma_{r}^{2}I),\quad\boldsymbol{\mu}_{c}^{\mathrm{tr}}\sim\mathcal{N}(\mathbf{0},\sigma_{s}^{2}),$ (8)
where $C$ is the number of clusters in the training set. $\sigma_{r}^{2}$ is the scalar value representing the variance of the Gaussian distribution corresponding to each cluster. $\sigma_{s}^{2}$ is the scalar value representing the variance of the distribution of cluster centers ($\boldsymbol{\mu}_{c}^{\mathrm{tr}}$). Brain activity data, $\mathbf{x}\in\mathbb{R}^{D}$, were generated using teacher weights $\bar{A}\in\mathbb{R}^{D\times D}$ and incorporating observation noise $\boldsymbol{\xi}$ with $\mathbf{x}=\bar{A}^{\top}\mathbf{y}+\boldsymbol{\xi}$, where $\boldsymbol{\xi}\sim\mathcal{N}(\mathbf{0},\sigma_{n}^{2}I)$. $\sigma^{2}_{n}$ is the scalar value representing the variance of observation noise. Using $N$ training samples, the ridge regression model was trained. With the hyperparameter $\lambda$, the weight of ridge regression model $W$ can be calculated analytically:
$W=(X_{\mathrm{tr}}^{\top}X_{\mathrm{tr}}+\lambda I)^{-1}X_{\mathrm{tr}}^{\top}Y_{\mathrm{tr}}.$ (9)
For testing the model performance, we prepared two types of test samples, in-distribution test samples and out-of-distribution test samples. In-distribution test samples are generated using the same parameter as $p_{\mathrm{tr}}(\mathbf{y})$. Out-of-distribution (OOD) test samples were generated in two steps. First, $C_{\mathrm{ood}}$ different cluster centers $\boldsymbol{\mu}_{c}^{\mathrm{ood}}$ were obtained by sampling from a Gaussian distribution: $\boldsymbol{\mu}_{c}^{\mathrm{ood}}\sim\mathcal{N}(\mathbf{0},\sigma_{s}^{2})$. Then OOD test samples were generated from these new cluster centers:
$p_{\mathrm{ood}}(\mathbf{y})=\frac{1}{C_{\mathrm{ood}}}\sum_{c=1}^{C_{\mathrm{ood}}}\mathcal{N}(\boldsymbol{\mu}_{c}^{\mathrm{ood}},\sigma_{r}^{2}I).$ (10)
To evaluate the model’s zero-shot performance, we evaluate $C+1$ way cluster identification analyis on the $C$ training clusters and the one cluster to which each OOD test sample belongs. Similar to Figure 6c, we caluculated the proportion of predicted features from test brain data that were correctly classified as belonging to the original cluster centers compared with the other candidate clusters. The similarity between predicted features and each cluster center was calculated using the correlation coefficient, with the highest similarity determining the cluster assignment, where the chance level is $1/(C+1)$. We evaluated the cluster identification performance for each of the OOD clusters and reported the median performance. Note that, for in-distribution test samples, we also conducted a C + 1 way identification analysis on the C training clusters and one selected OOD cluster for fair comparison.
In the simulation analysis, we mainly focused on the cluster identification performance for the dimension $D$, for the number of clusters in the training set $C$ or the distance among cluster centers for large sample size $N=500000$. Other parameters fixed in this simulation analysis as follows: $\sigma_{r}^{2}=10/D$, $\sigma_{s}^{2}=100/D$, $\bar{A}_{ij}\sim\mathcal{N}(0,D^{-1/2})$, $\sigma_{n}^{2}=0.1$, $\lambda=1.0$, and $C_{\mathrm{ood}}=32$. We parameterized the scale of the variances ($\sigma_{r}^{2}$ and $\sigma_{s}^{2}$) and the teacher weights $\bar{A}$ using the dimension $D$ so that the order of magnitude of the output invariant from the dimension.
4.5 Recovery check of a single layer by iCNN
We performed a recovery check analysis using a single layer from the iCNN method at Figure 10. Briefly, the iCNN method generates an image by optimizing pixel values to make the image’s latent features similar to the target latent features (Shen et al., 2019b). In the pixel optimization condition (the left columns of each recovery image in Figure 10), we directly optimized the pixel values of images to minimize the mean squared loss between the latent feature of the image as well as the total-variance (TV) loss of pixel values (Mahendran and Vedaldi, 2015).
Additionally, the iCNN method can incorporate image generator networks (middle and right columns of each recovery image in Figure 10) to add constraints on image statistics. Instead of optimizing the pixel values, we optimized the parameters related to the generator networks to minimize the mean squared loss between the latent features obtained through the generator networks and the target latent features. As a weak image prior, we used Deep image prior (DIP) (Ulyanov et al., 2017). DIP utilizes a hierarchical U-Net architecture as an inherent prior for image tasks, capturing the statistical regularities of images without relying on a specific dataset. This model works effectively by optimizing a randomly initialized neural network that can be used as an image prior in various inverse problems such as denoising, super-resolution, and inpainting tasks. In our analysis, DIP started with a U-Net initialized with random noise. Subsequently, the latent features and parameters of DIP were optimized to minimize the difference between the network’s output and target DNN features. For the pretrained image prior, we used the same generator model as in Shen et al. (2019b) (Dosovitskiy and Brox, 2016), optimizing the latent features of the pretrained networks.
4.6 Expected identification accuracy in imprecise reconstructions
A pairwise identification accuracy is a metric defined on three types of samples: the test sample, the predicted sample, and the candidate sample selected from a test set as $\mathbf{y},\mathbf{\hat{y}},\mathbf{y}_{-}\in\mathcal{Y}$, respectively. We define a function $S\colon\mathcal{Y}\times\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$ that takes the aforementioned triplet as the input and output whether the predicted sample was much close to the test sample than the candidate sample as
$S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})=\begin{cases}1&\left(\operatorname{sim}(\hat{\mathbf{y}},\mathbf{y})\geq\operatorname{sim}(\hat{\mathbf{y}},\mathbf{y}_{-})\right)\\ 0&\left(\text{otherwise}\right)\end{cases}$ (11)
where $\operatorname{sim}(\cdot,\cdot)$ is an arbitrary function that evaluates a similarity between two samples. The pairwise identification accuracy $\mathrm{Acc}$ over $n$ test samples is defined as
$\mathrm{Acc}=\frac{1}{n(n-1)}\sum_{i}^{n}\sum_{j\neq i}^{n}S(\hat{\mathbf{y}}^{(i)},\mathbf{y}^{(i)},\mathbf{y}_{-}^{(i)}).$ (12)
Now we consider a scenario where the translator only decodes semantic information (e.g., category) and cannot decode information about its precise visual appearance. Suppose the test set contains a categorical structure like NSD stimuli, we model such a scenario as
$\mathbb{E}_{p(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-}\mid(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in Z)}[S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})]$ $=0.5,$ (13)
$\mathbb{E}_{p(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-}\mid(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in\bar{Z})}[S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})]$ $=q\text{ where }q\in[0.5,1].$ (14)
$Z$ is a set of triplet of which the test sample and the candidate sample belong to the same category. $\bar{Z}$ is a complementary set of $Z$. $\mathbb{E}_{p(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-}\mid\cdot)}[S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})]$ represents the pairwise identification accuracy in the conditional expectation form. If the candidate sample belongs to the same category as the test sample, pairwise identification is challenging because of the poor prediction of the translator. On the other hand, if the candidate samples belongs to a different category than the test sample, the test sample is easily identified only from the semantic information.
Here, we assume that the test set contains $k$ categories in total and all samples are equally distributed across each category for simplicity. If we have sufficiently large number of test samples, the above identification accuracy can be approximated as
$\mathrm{Acc}$ $\underset{\phantom{n\to\infty}}{=}\frac{1}{n(n-1)}\left(\sum_{(\mathbf{y},\mathbf{\hat{y}},\mathbf{y}_{-})\in Z}S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})\right)+\frac{1}{n(n-1)}\left(\sum_{(\mathbf{y},\mathbf{\hat{y}},\mathbf{y}_{-})\in\bar{Z}}S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})\right)$ (15)
$\underset{\phantom{n\to\infty}}{=}\frac{|Z|}{n(n-1)}\left(\frac{1}{|Z|}\sum_{(\mathbf{y},\mathbf{\hat{y}},\mathbf{y}_{-})\in Z}S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})\right)+\frac{|\bar{Z}|}{n(n-1)}\left(\frac{1}{|\bar{Z}|}\sum_{(\mathbf{y},\mathbf{\hat{y}},\mathbf{y}_{-})\in\bar{Z}}S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})\right)$ (16)
=$\underset{n\to\infty}{=}\frac{1}{k}\cdot\mathbb{E}_{p(\mathbf{\hat{y}},\mathbf{y},\mathbf{y}_{-}\mid(\mathbf{y},\mathbf{\hat{y}},\mathbf{y}_{-})\in Z)}[S]+\left(1-\frac{1}{k}\right)\cdot\mathbb{E}_{p(\mathbf{\hat{y}},\mathbf{y},\mathbf{y}_{-}\mid(\mathbf{y},\mathbf{\hat{y}},\mathbf{y}_{-})\in\bar{Z})}[S]$ (17)
$\underset{\phantom{n\to\infty}}{=}\frac{1}{k}\cdot 0.5+\left(1-\frac{1}{k}\right)\cdot q,$ (18)
where $|Z|=n\left(n/k-1\right)$, and $|\bar{Z}|=n\left(n-n/k\right)$. We used the assumption on the large sample size at the third equality.
5 Acknowledgements
We thank our laboratory team, especially Eizaburo Doi, Hideki Izumi, Matthias Mildenberger and for their invaluable comments and suggestions on the manuscript. This work was supported by the Japan Society for the Promotion of Science (JSPS: KAKENHI grants JP20H05954, JP20H05705 to Y.K, JP21K17821 to Y.N, and 22KJ1801 to K.S.), Japan Science and Technology Agency (JST: CREST grants JPMJCR18A5, and JPMJCR22P3 to Y.K.), and New Energy and Industrial Technology Development Oprganization (NEDO: commissioned project, JPNP20006 to Y.K.).
Appendix A Supplementary figures
Figure S1: UMAP visualization of CLIP text features in Deeprecon. Blue points represent training samples, while red points represent test samples.
Figure S2: Supplementary information of hold out analysis. (a) The clustering results, applying the $k$-Means clustering to the UMAP embedding of CLIP text features. Based on these clustering results, the assignment of training and test samples under the hold-out split condition was determined. (b) Confusion probability matrix of cluster identification results. The top row shows the cluster identification results under the hold-out split conditions. The left side is for CLIP text features and the right side is for CLIP vision features. The bottom row shows the cluster identification results under the naive split conditions. In each matrix, the $(i,j)$ cell indicates the ratio of samples assigned by cluster $i$ are predicted to cluster $j$. The empty cell means the ratio is 00.
Figure S3: Additional results about simulation with clustered data. (a) Cluster identification accuracy as a function of the cluster variance ratio. The $x$ axis indicates the cluster variance ratio $\sigma_{r}^{2}/\sigma_{s}^{2}$ and the $y$ axis represents the cluster identification performance. Green lines show the results of in-distribution test samples and orange lines show those for OOD samples. The number of clusters used in training is increasing as the color lightens. Similar to Figure 9b in the main figure, the OOD identification performance become closer to the in distribution samples as the cluster variance ratio is increasing. (b). Heatmap plot of cluster identification accuracy varying the number of clusters in training and dimension. The left matrix shows the cluster identification results for the in-distribution test samples. The right matrix shows the cluster identification results for the OOD test samples. In each matrix, the $(i.j)$ cell indicates the cluster identification accuracy when the dimension is $i$ and the number of clusters used in the training is $j$.
Figure S4: Frame simirality in Nishimoto et al. (2011). We first detected scenes from the training and test movie stimuli. Then, for the first and last frames of each test scene, we searched for the frame from the training movies with the closest Euclidean distance. Our analysis revealed that 37 out of the 48 scenes in the test set contained frames that were nearly identical to those in the training set.

### v3

@@
 Abstract
-The rapid advances in brain decoding, particularly visual image reconstruction, have sparked discussions about the potential societal implications and ethical considerations surrounding neurotechnology.
-As these methods aim to recover perceived images from brain activity and achieve prediction over diverse images beyond training samples (zero-shot prediction), it is crucial to critically assess their capabilities and limitations to prevent misguided public expectations and inform future regulations.
-Our case study of recent text-guided reconstruction methods, which leverage a large-scale dataset (the Natural Scene Dataset, NSD) and text-to-image diffusion models, reveals significant limitations in their generalizability.
-We found a notable decrease in performance when applying these methods to a different dataset, which was designed to prevent category overlaps between training and test sets.
-UMAP visualization of the text features with NSD images showed a limited diversity of distinct semantic and visual clusters, with substantial overlap between training and test sets.
-Formal analysis and simulations demonstrated that clustered training samples can lead to “output dimension collapse,” restricting the output feature dimensions predictable from brain activity.
-Diversifying the training set to ensure a broader feature distribution improved generalizability beyond the trained clusters.
-However, text features alone are insufficient for a complete mapping to the visual space, even if perfectly predicted from brain activity.
-We argue that recent photo-like reconstructions may primarily be a blend of classification into trained categories and the generation of convincing yet inauthentic images through text-to-image diffusion (hallucination).
-To achieve genuine zero-shot prediction, diverse datasets and compositional representations spanning the image space are essential.
-As neurotechnology advances, engaging in interdisciplinary discussions involving neuroscientists, ethicists, policymakers, and the public is crucial to ensure responsible development and application of these techniques.
-These discussions should be grounded in a clear understanding of the current capabilities and limitations of the technology, as well as a careful consideration of the potential ethical and societal impacts.
-Keywords Visual image reconstruction  $\cdot$ brain decoding  $\cdot$ zero-shot prediction  $\cdot$ methodology  $\cdot$ generative AIs
-1 Introduction
-Brain decoding has been widely used in the neuroscience field, revealing specific contents of the mind (Haxby et al., 2001; Kamitani and Tong, 2005; Soon et al., 2008; Horikawa et al., 2013).
-As brain decoding is sometimes referred to as “mind-reading” in popular media (Somers, 2021; Whang, 2023; Raasch, 2023), it has attracted significant attention beyond the scientific community due to its potential for real-world applications in medicine and industry.
-Such neurotechnology has also started to affect future ethical discussions and legal regulations (UNESCO, 2023).
-To prevent misguiding public expectations and policies, scientists need to carefully assess the current status of brain decoding techniques and clarify the possibilities and limitations.
+Advances in brain decoding, particularly visual image reconstruction, have sparked discussions about the societal implications and ethical considerations of neurotechnology.
+As these methods aim to recover visual experiences from brain activity and achieve prediction beyond training samples (zero-shot prediction), it is crucial to assess their capabilities and limitations to inform public expectations and regulations.
+Our case study of recent text-guided reconstruction methods, which leverage a large-scale dataset (Natural Scene Dataset, NSD) and text-to-image diffusion models, reveals limitations in their generalizability.
+We found poor performance when applying these methods to a different dataset designed to prevent category overlaps between training and test sets.
+UMAP visualization of the text features with NSD images showed a limited diversity of semantic and visual clusters, with overlap between training and test sets.
+Formal analysis and simulations demonstrated that clustered training samples can lead to “output dimension collapse,” restricting predictable output feature dimensions.
+Simulations further showed that diversifying the training set improved generalizability.
+However, text features alone are insufficient for mapping to the visual space.
+We argue that recent realistic reconstructions may primarily be a blend of classification into trained categories and generation of inauthentic images through text-to-image diffusion (hallucination).
+Diverse datasets and compositional representations spanning the image space are essential for genuine zero-shot prediction.
+Interdisciplinary discussions grounded in understanding the current capabilities and limitations, as well as ethical considerations, of the technology are crucial for its responsible development.
+Author summary
+Brain decoding has gained attention not only in neuroscience but also in fields like computer science, and even policy-making.
+However, claims about its predictive performance often tend to be overly optimistic.
+Here, we critically examine recent brain decoding methods, particularly visual image reconstruction methods that use generative AIs to generate seemingly realistic images from brain activity.
+Our analyses reveal that the apparent success of these methods largely stems from exploiting similarities between training and test stimuli in the experiment and from the hallucination of recent generative AIs, failing to achieve the genuine zero-shot reconstruction of novel images.
+We introduce the concept of “output dimension collapse,” a fundamental limitation in predicting within a target space, and discuss conditions necessary for true zero-shot prediction in terms of compositional representations and image recovery from latent features.
+We also highlight the limitations of current evaluation metrics and provide a framework for interpreting seemingly realistic reconstructions.
+By examining current limitations and proposing improvement strategies, we aim to encourage rigorous development and evaluation of brain decoding techniques.
+Introduction
+Brain decoding has been widely used in the neuroscience field, revealing specific contents of the mind [1, 2, 3, 4].
+As brain decoding is sometimes referred to as “mind-reading” in popular media [5, 6, 7], it has attracted significant attention beyond the scientific community due to its potential for real-world applications in medicine and industry.
+Such neurotechnology has also started to affect future ethical discussions and legal regulations [8].
+To prevent misleading public expectations and policies, scientists need to carefully assess the current status of brain decoding techniques and clarify the possibilities and limitations.
 One of the major challenges in brain decoding is the limited amount of brain data we can collect.
-The current brain measurement devices are costly, yielding far less brain data than the amounts typically used in image or text processing within the field of computer science and AI (Deng et al., 2009; Schuhmann et al., 2022).
-Although we have gradually increased the amount of brain data per subject (Van Essen et al., 2012; Allen et al., 2022; Naselaris et al., 2021; Hebart et al., 2023; Xu et al., 2024), it remains impractical to collect brain data covering the full range of cognitive states and perceptual experiences.
+The current brain measurement devices are costly, yielding far less brain data than the amounts typically used in image or text processing within the field of computer science and AI [9, 10].
+Although we have gradually increased the amount of brain data per subject [11, 12, 13, 14, 15], it remains impractical to collect brain data covering the full range of cognitive states and perceptual experiences.
 Consequently, classification-based decoding approaches, primarily developed in the early stages of this field, are insufficient to uncover the neural code in general or natural conditions since decodable information is confined to the same stimuli or predefined categories used in a training phase.
-To overcome this limitation, several decoding methods have been proposed to enable the prediction of novel contents from brain activities that are not encountered during the training phase.
+To overcome this limitation, several decoding methods have been developed to enable the prediction of novel contents from brain activities that are not encountered during the training phase.
 Kay et al.
-(2008) proposed a general visual decoding approach via a statistical encoding model that predicted fMRI voxel values from image features.
-It successfully identified novel test images from a set of 1000 candidates.
+[16] proposed a general visual decoding approach via a statistical encoding model that predicted fMRI voxel values from image features.
+It successfully identified novel test images from a set of 1,000 candidates.
 Mitchell et al.
-(2008) utilized co-occurrence rates of specific verb sets for nouns to predict fMRI brain activity during the perception of line drawing images, demonstrating the ability to predict unseen nouns.
-Brouwer and Heeger (2009) constructed a color-tuning model and predicted brain activity while the subjects were presented with color stimuli.
-As their training stimuli cover most of the color space, their methods can successfully identify novel colors not included in the training dataset.
-Horikawa and Kamitani (2017) utilized deep neural network (DNN) features to decode brain activity measured while subjects perceived natural images.
+[17] utilized co-occurrence rates of specific verb sets for nouns and built a computational model to predict fMRI voxel values while thinking about nouns in presented line drawing images.
+Their model demonstrated the ability to predict voxel values for novel nouns not seen during the training phase.
+Brouwer and Heeger [18] constructed a color-tuning model and predicted brain activity while the subjects were presented with color stimuli.
+As their training stimuli covered most of the color space, their methods successfully identified novel colors not included in the training dataset.
+Horikawa and Kamitani [19] utilized deep neural network (DNN) features to decode brain activity measured while subjects perceived natural images.
 They showed successful prediction of novel object categories not encountered during the training phase.
-In the field of machine learning, “zero-shot” prediction refers to the ability of a model to accurately predict or classify novel contents that were not encountered during the training phase (Larochelle et al., 2008; Palatucci et al., 2009).
-This technique has found widespread application across various domains, including image classification (Radford et al., 2021), image generation (Ramesh et al., 2021), and natural language processing (Brown et al., 2020).
+In the field of machine learning, “zero-shot” prediction refers to the ability of a model to accurately predict or classify novel contents not encountered during the training phase [20, 21].
+This capability has emerged in various applications across different domains, including image classification [22], image generation [23], and natural language processing [24].
 The concept of zero-shot prediction can be considered analogous to brain decoding techniques that aim to interpret brain activity patterns associated with previously unseen stimuli or experiences.
 Both approaches seek to generalize knowledge gained from a limited set of training data to novel situations, enabling the interpretation of new information without explicit prior exposure.
-To achieve effective zero-shot prediction, the model often utilizes a compositional representation of the output (Lake et al., 2017; Higgins et al., 2018).
+To achieve effective zero-shot prediction, the model often utilizes a compositional representation of the output [25, 26].
 Compositional representation refers to the ability to understand and generate novel combinations of previously learned concepts or features.
@@
 Visual image reconstruction is another prominent example of zero-shot prediction in brain decoding.
-This task aims to recover perceived novel images that were not encountered during the training phase, effectively reconstructing visual experiences from brain activity patterns (Stanley et al., 1999; Miyawaki et al., 2008).
+This task aims to recover perceived novel images that were not encountered during the training phase, effectively reconstructing visual experiences from brain activity patterns [27, 28].
 As our perceptual visual experiences cannot be fully covered by limited brain data, reconstruction methods require strong generalizability.
 Miyawaki et al.
-(2008) conducted a study that demonstrated the reconstruction of perceived arbitrary 10x10 binary-contrast images from brain activity.
+[28] conducted a study demonstrating the reconstruction of perceived arbitrary $10\times 10$ binary-contrast images from brain activity.
 They built multiple modular decoders to predict the local contrasts of each location and combined their predictions.
@@
 Similarly, Shen et al.
-(2019b) replaced local decoders with DNN feature decoders.
-Although their training stimuli were 1200 natural images, they demonstrated reconstructing novel images, including artificial images, which were not part of the training set.
-These successes suggest that the proposed reconstruction models capture rich and comprehensive information about the general aspects of the neural code, beyond merely the information defined by the training data (Kriegeskorte and Douglas, 2019).
-Developing reliable reconstruction methods also enables further analysis of subjective visual experiences, such as visual imagery (Shen et al., 2019b), attention (Horikawa and Kamitani, 2022), and illusion (Cheng et al., 2023).
-Enabling the decoding of novel brain states that were never encountered during the training phase can be a promising approach to neural mind-reading (Kamitani and Tong, 2005).
-Visual image reconstruction pipelines can be divided into three main components: translator, latent features, and generator (Figure 1).
-The translator serves to convert a brain activity pattern into a latent feature space via linear regression (Shen et al., 2019b; Seeliger et al., 2018; Mozafari et al., 2020; Ozcelik et al., 2022; Kavasidis et al., 2017), or nonlinear transformation (Qiao et al., 2020).
-Latent features serve as surrogate representations of perceived visual images.
-While the local contrasts in Miyawaki et al.
-(2008) can be seen as a primitive form of latent features, recent studies often use DNN features, such as the intermediate output of recognition models (Shen et al., 2019b; Cheng et al., 2023).
+[29] replaced local decoders with DNN feature decoders.
+Although their training stimuli were $1,200$ natural images, they demonstrated reconstructing novel images, including artificial images, which were not part of the training set.
+These successes suggest that the proposed reconstruction models capture rich and comprehensive information about the general aspects of the neural code, beyond merely the information defined by the training data [30].
+Developing reliable reconstruction methods also enables further analysis of subjective visual experiences, such as visual imagery [29], attention [31], and illusion [32].
+Enabling the decoding of novel brain states that were never encountered during the training phase can be a promising approach to neural mind-reading [2].
+Visual image reconstruction pipelines typically comprise three main components: translator, latent features, and generator (Fig 1).
+The translator converts brain activity patterns into a latent feature space, employing either linear regression [29, 33, 34, 35] or nonlinear transformation [36].
+Latent features serve as surrogate representations of perceived visual images, evolving from primitive forms like local contrasts [28] to more sophisticated DNN features, such as intermediate outputs of recognition models [19, 29].
+In our current work, we reframe this process as “translation” rather than “feature decoding,” a term we used in previous studies.
+This terminology acknowledges two important points: first, brain activity itself can be considered a latent representation of an image or perception of it, and second, it helps avoid potential ambiguity between image encoding/decoding and brain encoding/decoding processes.
+This new perspective conceptualizes the process as a translation between two latent spaces: from neural representation to machine representation.
+Fig 1: Visual image reconstruction pipeline.
+The first step involves translating brain activity patterns into machine/AI latent representations.
+While this process typically involves brain decoding of latent features using machine learning, here we term it “translation.” This terminology acknowledges that brain activity itself can be viewed as a latent representation of an image, framing the process as a translation from neural to machine latent representations.
+In the second step, a generator module takes these translated latent features as input and converts them into a visual image that corresponds to the content represented by the original brain activity.
 The generator visualizes translated latent features into images.
-Some studies have used the pretraining image generative models for the generator module (Mozafari et al., 2020; Qiao et al., 2020; Ozcelik et al., 2022; Kavasidis et al., 2017).
-Image optimization can also be regarded as a generator (Shen et al., 2019b).
-Direct mapping from brain activity to images using DNNs can also be considered to implicitly contain these components (Shen et al., 2019a; Beliy et al., 2019; Ren et al., 2021; Gaziv et al., 2022; Lin et al., 2022; Chen et al., 2023a).
-Figure 1: An overview of the visual image reconstruction pipeline.
-Brain activity recorded while perceiving a visual image is fed into a translator module.
-The translator maps the brain activity to latent features.
-The generator module converts the latent features into an image corresponding to the original visual stimulus.
-Recent advances in generative AI, particularly in text-to-image generation, have naturally given rise to expectations that these techniques could provide a shortcut for visual image reconstruction by leveraging semantic representations.
+Some studies have used pretraining image generative models for the generator module [34, 36, 35].
+Image optimization can also be regarded as a generator [29].
+End-to-end mapping from brain activity to images using DNNs can also be considered to contain these components implicitly or as a generator-only method [37, 38, 39, 40, 41, 42, 43].
+Recent advances in generative AI, particularly in text-to-image generation, have naturally given rise to expectations that these techniques could provide a valuable tool for visual image reconstruction by leveraging semantic representations.
 In addition, there has been a growing trend towards collecting neural datasets using a wide range of diverse visual and semantic content.
-This shift aims to capture a more comprehensive and ecologically valid representation of the human experience (Naselaris et al., 2021).
-Researchers have started to collect large-scale datasets, such as the Natural Scene Dataset (NSD) (Allen et al., 2022) and the THINGS-fMRI dataset (Hebart et al., 2023), which incorporate a broader range of fMRI data induced by diverse visual stimuli with text or category annotations.
-Notably, recent studies leveraging the latest generative AI techniques with semantic representation and large-scale datasets have reported photo-like reconstructions from brain activity (Takagi and Nishimoto, 2023b; Ozcelik and VanRullen, 2023; Scotti et al., 2023; Bai et al., 2023; Chen et al., 2023b; Lan et al., 2023; Benchetrit et al., 2024).
-They utilized contrastive language-image pretraining (CLIP) features (Radford et al., 2021) as latent features and text-to-image diffusion models (Ramesh et al., 2021; Xu et al., 2022) as generators.
+This shift aims to capture a more comprehensive and ecologically valid representation of the human experience [13].
+Researchers have started to collect large-scale fMRI datasets, such as the Natural Scene Dataset (NSD) [12] and the THINGS-fMRI dataset [14], which incorporate a broader range of brain data induced by diverse visual stimuli with text or category annotations.
+Notably, recent studies leveraging the latest generative AI techniques with semantic representation and large-scale datasets have reported realistic reconstructions from brain activity [44, 45, 46, 47, 48].
+They utilized contrastive language-image pretraining (CLIP) features [22] as latent features and text-to-image diffusion models [23, 49] as generators.
 While these recent approaches show promising results, it remains uncertain whether these methods truly achieve zero-shot reconstruction due to several factors.
@@
 To fully assess the zero-shot prediction capabilities of these approaches, it is essential to rigorously test their generalizability across different datasets and to provide detailed analyses of the individual model components.
-This includes evaluating the performance of the translators, latent features, and generators used in these methods.
+This test includes evaluating the performance of the translators, latent features, and generators used in these methods.
 Furthermore, the characterization of the diversity of stimuli in the datasets and the latent representations has not been thoroughly explored.
 It is unclear whether the recently proposed datasets, such as the NSD, are optimally designed to capture the full range of human visual experiences and to support the development of truly generalizable prediction models.
-In the following, we begin with a case study that critically tests the methods proposed by Takagi and Nishimoto (2023b) and Ozcelik and VanRullen (2023), which originally used the NSD.
-Our findings reveal several issues, including the failure of replication with another dataset specifically designed to avoid overlap between training and test sets (Shen et al., 2019b), and the questionable post-hoc selection procedures of Takagi and Nishimoto (2023b) that can produce convincing reconstructions even with random brain data.
-We identify the lack of diversity represented by the limited number of clusters in the NSD dataset as a potential factor contributing to these issues.
-We also demonstrate the failure of zero-shot prediction in the feature space and the inability to recover a stimulus from its latent features, suggesting a misspecification of the model.
-Based on these findings, we conclude that the apparent photo-like reconstructions are essentially a result of classification to clusters shared between training and test sets, combined with hallucinations by the generative model.
+In the following, we begin with a case study that critically tests the methods proposed by Takagi and Nishimoto [44] and Ozcelik and VanRullen [45], which originally used the NSD.
+Our findings reveal several issues, including the failure to replicate the same level of reconstruction with another dataset specifically designed to avoid overlap between training and test sets [29], and the questionable post-hoc selection procedures of Takagi and Nishimoto that can produce convincing reconstructions even with random brain data.
+We identify the lack of diversity represented by the limited number of semantic clusters in the NSD dataset as a potential factor contributing to these issues.
+We also demonstrate the failure of zero-shot prediction in the latent feature space and the inability to recover a stimulus from its latent features, suggesting a misspecification of the model.
+Based on these findings, we conclude that the apparent realistic reconstructions are essentially a result of classification to clusters shared between training and test sets, combined with hallucinations by the generative model.
 In the formal analysis and simulation section, we aim to uncover the general factors underlying the issues and observations highlighted in the case study.
 We first describe the phenomenon of “output dimension collapse” in the translation from brain activity to the latent feature space.
-We show that the regression model trained on the clustered outputs becomes specialized to the training examples and its output collapses into a subspace of the training dataset.
+We show that the regression model trained on the clustered targets becomes specialized to the training examples, and its outputs collapse into a subspace of the training set.
 Our simulations with clustered data demonstrate that out-of-sample prediction can be achieved with an increasing number of training clusters, suggesting that zero-shot prediction is possible if stimulus diversity and a compositional representation are available.
-We also explore the preservation of image information at hierarchical layers of DNNs and discuss the caveats associated with evaluating reconstructions using identification metrics alone.
+We also discuss the caveats associated with evaluating reconstructions using identification metrics alone and explore the preservation of image information at hierarchical layers of DNNs.
 Finally, we provide speculation on how we are fooled by the seemingly realistic reconstructions generated by AI models.
@@
 Our findings highlight the importance of rigorous evaluation, the use of diverse datasets, and the need for careful specification and characterization of model components to advance the field of visual image reconstruction and develop truly generalizable prediction capabilities.
-2 Results
-2.1 Case study
-We investigated two types of recent generative AI-based reconstruction methods, StableDiffusionReconstruction (Takagi and Nishimoto, 2023b) and Brain-diffuser (Ozcelik and VanRullen, 2023), as well as their validation dataset, the NSD (Allen et al., 2022).
-Both reconstruction methods utilize CLIP features (Radford et al., 2021) to effectively apply recent text-to-image diffusion models in visual image reconstruction analysis.
+Results
+Case study
+We investigated two types of recent generative AI-based reconstruction methods, StableDiffusionReconstruction [44] and Brain-Diffuser [45], as well as their validation dataset, the Natural Scene Dataset (NSD) [12].
+Both reconstruction methods utilize CLIP features [22] to effectively apply recent text-to-image diffusion models in visual image reconstruction analysis.
 CLIP text features are obtained from the average of five text annotations corresponding to the stimulus image.
-These text annotation information are only used during training to map the brain activity into CLIP text features.
-In the test phase, they directly predict CLIP text features from brain activity during image perception.
-Henceforth, these two reconstruction methods will be referred to together as text-guided reconstruction methods.
-The StableDiffusionReconstruction method (Takagi and Nishimoto, 2023b) uses the components of the Stable Diffusion model (Ramesh et al., 2021), the VAE features (Kingma and Welling, 2014) and CLIP text features as latent features.
-Similarly, the Brain-diffuser method (Ozcelik and VanRullen, 2023) utilizes the component of another type of diffusion model (Xu et al., 2022), CLIP text features, and CLIP vision features as well as VDVAE features (Child, 2021) as latent features.
+This text annotation information is only used during training to map the brain activity into the CLIP text features.
+In the test phase, they directly translate text features from brain activity during image perception.
+Hence, these two reconstruction methods will be referred to together as text-guided reconstruction methods.
+The StableDiffusionReconstruction method [44] uses the components of the Stable Diffusion model [50], the VAE features [51], and CLIP text features as latent features.
+Similarly, the Brain-Diffuser method [45] utilizes the component of another type of diffusion model [49], CLIP text features, CLIP vision features, and VDVAE features [52] as latent features.
 Both methods first generate initial images from translated VAE/VDVAE features.
@@
 Thanks to the authors of these studies who made efforts to make their datasets and scripts publicly available, we were able to conduct our replication analysis effectively.
-We compared the reconstructed results of these two text-guided reconstruction methods with those from a previous image reconstruction method, iCNN (Shen et al., 2019b).
-For more information on datasets and reconstruction methods, see Material and Methods (“Datasets” or “Reconstruction methods”) or the original studies (Allen et al., 2022; Shen et al., 2019b; Takagi and Nishimoto, 2023b; Ozcelik and VanRullen, 2023).
-2.1.1 Observations: Failed replication and convincing reconstruction from random data
-We first confirmed the reproduction of the findings of the text-guided reconstruction methods, particularly those of the Brain-diffuser study (Ozcelik and VanRullen, 2023) when applied to the NSD dataset (Figure 2a).
+We compared the reconstructed results of these two text-guided reconstruction methods with those from a previous image reconstruction method, iCNN [29].
+For more information on datasets and reconstruction methods, see Material and Methods (“Datasets” or “Reconstruction methods”) or the original studies [12, 29, 44, 45].
+Observations: Failed replication and convincing reconstruction from random data
+We first confirmed the reproduction of the findings of the text-guided reconstruction methods, particularly those of the Brain-Diffuser study [45] when applied to the NSD dataset (Fig 2A).
 The reconstructed images generated by the Brain-Diffuser method effectively captured most of the layout and semantics of the test images.
-Although the reconstructed images produced by the StableDiffusionReconstruction method (Takagi and Nishimoto, 2023b) showed slightly degraded performance compared to the original paper, they still successfully captured the semantics of the test images.
-Notably, the iCNN method (Shen et al., 2019b) also performed well on the NSD dataset, with reconstructed images capturing the dominant structures of the objects within the images, consistent with the findings reported in the original study.
+Although the reconstructed images produced by the StableDiffusionReconstruction method [44] showed slightly degraded performance compared to the original paper, they still successfully captured the semantics of the test images.
+Notably, the iCNN method [29] also performed well on the NSD dataset, with reconstructed images capturing the dominant structures of the objects within the images, consistent with the findings reported in the original study.
+Fig 2: Comparison of image reconstruction results across datasets and methods.
+(A) Reconstruction results for the Natural Scene Dataset (NSD).
+The first row shows the original test images, followed by reconstructions using StableDiffusionReconstruction, Brain-Diffuser, and iCNN methods.
+(B) Reconstruction results for the Deeprecon dataset, presented in the same format as (A).
 To further investigate the generalizability of the text-guided reconstruction methods, we attempted to replicate their performance using a different dataset, Deeprecon, which was originally collected for the study by Shen et al.
-(2019b).
-The Deeprecon dataset was specifically designed to avoid overlap between training and test sets, making it an ideal benchmark for evaluating the zero-shot prediction capabilities of reconstruction methods.
+[29].
+The Deeprecon dataset was explicitly designed to avoid overlap between training and test sets, making it an ideal benchmark for evaluating the zero-shot prediction capabilities of reconstruction methods.
 However, the original Deeprecon dataset lacked the text annotations required by the text-guided reconstruction methods.
 To facilitate a fair comparison, we collected five text annotations for each training stimulus in the Deeprecon dataset through crowd-sourcing and used them to generate CLIP text features.
-Despite the inclusion of these text annotations, the text-guided reconstruction methods failed to achieve the same level of performance on the Deeprecon dataset as they did on the NSD dataset.
-The reconstructed images produced by the text-guided methods exhibited photo-like appearances but suffered from largely degraded quality compared to their performance on the NSD dataset (Figure 2b).
-Notably, the text-guided reconstruction methods generated photo-like reconstructions even for simple geometric shapes present in the Deeprecon dataset, which deviated significantly from the original stimuli.
-In contrast, the iCNN method (Shen et al., 2019a) consistently provided faithful reconstructions for both the NSD and Deeprecon datasets, despite its simplicity compared to the text-guided methods.
-Figure 2: Image reconstruction results.
-(a) Reconstruction results for the NSD.
-The first row shows the original test images, followed by reconstructions using StableDiffusionReconstruction, Brain-Diffuser, and iCNN methods.
-(b) Reconstruction results for the Deeprecon dataset.
-The format is the same as in (a).
+Despite including these text annotations, the text-guided reconstruction methods failed to achieve the same level of performance on the Deeprecon dataset as they did on the NSD dataset.
+The reconstructed images produced by the text-guided methods exhibited realistic appearances but suffered from largely degraded quality compared to their performance on the NSD dataset (Fig 2B).
+Notably, the text-guided reconstruction methods generated realistic reconstructions even for simple geometric shapes in the Deeprecon dataset, which deviated significantly from the original stimuli.
+In contrast, the iCNN method consistently provided faithful reconstructions for both the NSD and Deeprecon datasets, despite its simplicity compared to the text-guided methods.
+These results suggest that the text-guided methods may not generalize well across different datasets.
+The ability to generate realistic but inaccurate reconstructions, especially for simple shapes, indicates that these methods might rely more on training stimuli than on actual brain activity information.
 Upon further investigation, we noted a questionable post-hoc image selection procedure.
-In the StableDiffusionReconstruction study (Takagi and Nishimoto, 2023b), they presented the reconstruction results by the following procedure: “We generated five images for each test image and selected the generated images with highest PSM” (PSM means perceptual similarity metrics), as illustrated in Figure 3a.
+In the StableDiffusionReconstruction study [44], they presented the reconstruction results by the following procedure: “We generated five images for each test image and selected the generated images with highest PSM” (PSM means perceptual similarity metrics, and it was calculated from early/middle/late layers of several image recognition deep neural networks [DNNs]), as illustrated in Fig 3A.
 This selection might lead readers or peer reviewers, particularly those not specialized in the brain decoding field, to overestimate the effectiveness of the methods and potentially lead to a distorted understanding of the actual reconstruction performance.
-Note that Brain-diffuser study (Ozcelik and VanRullen, 2023) did not executed such procedures and in their following report (Takagi and Nishimoto, 2023a), they updated the image presentation procedure more fairly as: “we generated five images with different stochastic noise and selected three images randomly.”
-Figure 3: Questionable post-hoc selection.
-(a) Procedure in the StableDiffusionReconstruction method.
-Five images were generated, and the one that most closely resembled the test images was selected.
-(b) Reconstruction using shuffled brain activity.
-The left panels show the images generated from the VAE features given shuffled brain activity and CLIP text features given original brain activity.
-The right panels show the images generated from the VAE and CLIP text features given shuffled brain activity.
-To examine the impact of this post-hoc selection procedure, we conducted an experiment using random brain data (Figure 3b).
-Instead of feeding the test brain data into pretrained translators, we shuffled the activities within each voxel of the NSD test set independently to create random brain data.
-Surprisingly, when the random brain data were input into the VAE feature translator, which contributes to producing initial images, plausible images were obtained by generating five images and selecting the best one.
-Even more strikingly, when the random brain data were input into both VAE and CLIP text feature translators, we still obtained convincing results by simply generating images five times and conducting the aforementioned selection.
-These observations are inexplicable since the artificially created brain data should completely lack any information related to the original visual stimuli.
+Note that the Brain-Diffuser study [45] did not execute such procedures, and in their subsequent report [53], they updated the image presentation procedure more fairly as: “we generated five images with different stochastic noise and selected three images randomly.”
+Fig 3: Analysis of post-hoc selection in the StableDiffusionReconstruction method [44].
+(A) Selection procedure.
+Five images were generated, and the one most closely resembling the test image was selected based on perceptual similarity metrics (PSM).
+(B) Images were generated from the CLIP text features translated from original brain activity and the VAE (vision) features from shuffled brain activity.
+Examples selected from five generations are shown with the test images.
+(C) Images were generated from the CLIP text and the VAE (vision) features, both translated from shuffled brain activity.
+Examples of selected images are shown with the text images.
+Examples are shown as in (B).
+To examine the impact of this post-hoc selection procedure, we conducted an experiment using random brain data.
+Instead of feeding the test brain data into trained translators, we shuffled the activities within each voxel of the NSD test set independently to create random brain data.
+Specifically, the brain data of the NSD test set is in a matrix shape, with rows representing stimulus samples and columns representing voxels.
+To generate the random brain data, we selected one column (voxel) of the matrix and randomly shuffled its values.
+This process was repeated for all voxels independently.
+Surprisingly, when the random brain data were input into the VAE feature translator, which contributes to producing initial images, plausible images were obtained by generating five images and selecting the best one (Fig 3B).
+Even more strikingly, when the random brain data were input into both VAE and CLIP text feature translators, we still obtained convincing results by simply generating images five times and conducting the selection mentioned above (Fig 3C).
+These observations are inexplicable because the artificially created brain data should completely lack any information related to the original visual stimuli.
 These observations raise perplexing questions about the performance and generalizability of recent text-guided reconstruction methods.
-The significant deterioration in performance when simply changing the dataset from NSD to Deeprecon highlights the need for a deeper understanding of the factors contributing to the success of these methods on the NSD dataset.
+The performance significantly deteriorated when we changed from the NSD to Deeprecon dataset.
+This highlights the need to better understand why these methods succeed with the NSD.
 Moreover, the ability to obtain plausible reconstructions from random brain data by merely generating multiple images and selecting the best ones suggests that there may be fundamental issues with both the evaluation dataset and the components of the reconstruction methods themselves.
-In the following sections, we will conduct a thorough investigation into the potential problems associated with the NSD dataset and each component of the text-guided reconstruction pipeline.
-2.1.2 Lack of diversity in the stimulus set
+In the following sections, we will thoroughly investigate the potential problems associated with the NSD dataset and each component of the text-guided reconstruction pipeline.
+Lack of diversity in the stimulus set
 First, we examined the characteristics and limitations of the NSD dataset itself.
-To characterize the diversity of stimuli in the datasets and their latent representations, we focused on the CLIP features, which are used as latent features in recent reconstruction methods.
-We employed uniform manifold approximation and projection (UMAP) (McInnes et al., 2018) to visualize the CLIP text features of the NSD stimuli (see Materials and Methods “UMAP visualization”).
-The visualization revealed approximately 40 distinct clusters, with considerable overlap between the training and test sets (Figure 4a).
+To characterize the diversity of stimuli in the datasets and their latent representations, we focused on the CLIP text features, which are used as latent features in text-guided reconstruction methods.
+We employed uniform manifold approximation and projection (UMAP) [54] to visualize the CLIP text features of the NSD stimuli (see Materials and Methods “UMAP visualization”).
+The visualization revealed approximately $40$ distinct clusters, with considerable overlap between the training and test sets (Fig 4A).
 Interestingly, we were able to describe the stimulus images in each cluster using a single semantic word, such as airplane, giraffe, or tennis.
 Despite the NSD containing around $30,000$ brain samples per subject, the diversity of the presented stimuli was quite limited to just around $40$ semantic concepts.
-In contrast, the Deeprecon dataset, which was specifically designed to differentiate object categories between training and test data, exhibited less overlap between the two sets (Supplementary Figure S1).
-To further investigate the similarity between the training and test stimuli, we conducted an analysis using a state-of-the-art perceptual similarity metric called DreamSim (Fu et al., 2023) (Figure 4b).
+Note that we performed UMAP visualization using the parameters recommended in the official guide for clustering, but a similar cluster structure was also observed with the default parameters (S1 Fig).
+In contrast, the Deeprecon dataset, which was specifically designed to differentiate object categories between training and test data, exhibited less overlap between the two sets (S2 Fig).
+Fig 4: Dataset diversity and similarity between training and test stimuli.
+(A) UMAP visualization of CLIP text features in the NSD.
+The center figure shows the scatter plot of the UMAP embedding of CLIP text features.
+The gray points represent training samples, while the orange points represent test samples.
+The surrounding images were randomly selected from each cluster.
+(B, C) Similarity between training and test images.
+For each example test image (the one on the left of each row), the five training images with the highest similarity were selected using the DreamSim metric and displayed.
+This analysis was performed for the NSD (B) and Deeprecon (C) datasets, using the same procedure.
+To further investigate the similarity between the training and test stimuli, we analyzed using a state-of-the-art perceptual similarity metric called DreamSim [55].
 For each test image in the NSD, we used DreamSim to identify the most similar training images based on their perceptual similarity scores.
-We observed that the training images extracted using DreamSim were highly similar to the test images in the NSD, not only in terms of semantic concepts but also in terms of overall layout and visual composition.
-This finding suggests that the NSD test images are heavily biased towards the distribution of the training data, with a significant overlap in the visual and semantic features present in both sets.
-Such a strong bias raises concerns about the true reconstruction performance of methods evaluated on this dataset, as the impressive results may be largely attributable to the methods’ ability to memorize and reproduce the characteristics of the training data rather than generalizing to novel stimuli.
-In contrast, when we applied the same DreamSim-based analysis to the Deeprecon dataset, we found that the extracted training images were substantially different from the test images.
-This observation indicates a clearer separation between the training and test distributions in the Deeprecon dataset, with the test images containing novel visual and semantic features that are not well-represented in the training set.
+We found that the training images extracted using DreamSim were highly similar to the test images in the NSD, not only in terms of semantic concepts but also in terms of overall layout and visual composition (Fig 4B).
+These findings suggest that the distribution of NSD test images is heavily biased towards that of the training images, with a significant overlap in the visual and semantic features present in both sets.
+Such a strong bias raises concerns about the actual reconstruction performance of methods evaluated on this dataset.
+The impressive results may be largely attributable to the methods’ ability to memorize and reproduce the characteristics of the training stimuli rather than generalizing them to novel stimuli.
+In contrast, when we applied the same DreamSim-based analysis to the Deeprecon dataset, we found that the extracted training images were substantially different from the test images (Fig 4C).
+This observation supports the careful design of the train-test split in the Deeprecon dataset, ensuring that the test images contain novel visual and semantic features that are not well-represented in the training set.
 The distinct differences between the NSD and Deeprecon datasets in terms of stimulus similarity highlight the importance of carefully designing evaluation benchmarks that can effectively assess the generalization capabilities of visual image reconstruction methods.
-Figure 4: Dataset analysis.
-(a) UMAP visualization of CLIP text features in NSD.
-The center figure shows the scatter plot of the UMAP embedding of CLIP text features in NSD.
-Green points represent training samples, while orange points represent test samples.
-The surrounding images show the images in cluster.
-The presented images in each cluster were randomly selected.
-(b) Similarity between training and test images using the DreamSim metric.
-The left side shows results for the NSD, and the right side shows results for the Deeprecon dataset.
-The top 5 training images with highest similarity to each test image were selected.
-2.1.3 Failed zero-shot prediction in the feature space
-Given the significant semantic and structural similarities between the NSD test images and the training images, it is crucial to assess whether the predictions made from brain activity truly demonstrate generalization beyond the training set (i.e., zero-shot prediction).
-To evaluate the zero-shot prediction capability of the translators used in the reconstruction methods, we conducted an $N+1$ way identification analysis, where $N$ represents all of the training samples and $1$ refers to the test sample (Figure 5a).
-We calculated the percentage of translated features that successfully identified the corresponding test sample in the NSD dataset.
-The performance of the CLIP features was nearly $0\%$ (Figure 5b).
-This extremely low performance suggests that the CLIP feature translator fails to achieve zero-shot prediction, even within the distribution covered by the training set.
-In other words, the CLIP features translated from brain activity are unable to accurately identify novel test samples, indicating a lack of generalizability beyond the training data.
-In contrast, the VGG19 features (Simonyan and Zisserman, 2015) employed in the iCNN method demonstrated moderate identification performance at the latent features on the intermediate layers.
-This finding suggests that the VGG19 features possess a certain degree of generalizability, enabling them to identify test samples that were not encountered during training.
-Figure 5: Zero-shot sample identification.
-(a) Schematic diagram of novel sample identification: calculating the similarity between predicted features and test features or each of the training features, and counting how many predicted features were more similar to the true features than other candidate training features.
-(b) Results for NSD.
-The left side shows the identification accuracies evaluated on the CLIP features, which were used in the Brain-diffuser methods.
-The right side shows the accuracies evaluated on several VGG19 features used in the iCNN method.
+Failed zero-shot prediction in the feature space
+Given the significant semantic and structural similarities between the NSD test images and the training images, it is crucial to assess whether the predictions derived from brain activity truly demonstrate generalization beyond the training set (i.e., zero-shot prediction).
+To evaluate the zero-shot prediction capability of the translators used in the reconstruction methods, we conducted an $(N+1)$-way identification analysis, where $N$ represents all of the training samples ($8,859$ samples for subject 1 in the NSD) and $1$ refers to the test sample (Fig 5A).
+Specifically, we calculated the similarity (correlation) between the predicted (translated) features and the true features of corresponding test samples, as well as the similarities between the predicted features and the true features corresponding to each of the N training samples.
+We then examined whether the predicted features correctly identified (showed the highest similarity to) the test true features.
+We calculated the percentage of predicted features that successfully identified the corresponding test sample in the NSD dataset.
+Fig 5: Evaluation of zero-shot sample identification performance.
+(A) Schematic diagram of novel sample identification.
+Similarity is calculated between predicted (translated) features and test features, as well as between predicted features and each of the training features.
+Identification accuracy is determined by how often the predicted features show the highest similarity to the true test features compared to all training features.
+(B) Zero-shot identification performance using predicted CLIP features (text and vision) from the Brain-Diffuser method, evaluated on the NSD dataset.
+(C) Zero-shot identification performance using predicted features from several intermediate layers of the VGG19 model employed in the iCNN method, also evaluated on the NSD dataset.
+For both (B) and (C), the identification task involved $8,859$ training samples plus one test sample, resulting in an $8,860$-way identification.
+The chance performance level is thus $1/8860$.
+The performance of the CLIP features was nearly 00% (Fig 5B).
+The CLIP features predicted from brain activity were unable to accurately identify novel test samples, even within the distribution covered by the training set.
+This extremely low performance suggests that the CLIP feature translator fails to achieve zero-shot prediction.
+This finding raises questions about the effectiveness of CLIP features for true zero-shot prediction tasks in brain decoding.
+In contrast, the VGG19 features [56] employed in the iCNN method demonstrated moderate identification performance at the latent features on the intermediate layers (Fig 5C).
+This finding suggests that the VGG19 features possess a certain degree of generalizability, enabling them to identify test samples not encountered during training.
 Furthermore, we investigated whether the CLIP feature translator enables the prediction of novel semantic concepts not in the training.
-We redesigned the dataset split to ensure no semantic concepts were shared between them as in the previous zero-shot prediction study (Mitchell et al., 2008; Brouwer and Heeger, 2009).
-We first applied $k$-means clustering to the UMAP embedding space of the NSD’s CLIP text features (Supplementary Figure S2a).
+We redesigned the dataset split to ensure no semantic concepts were shared between them as in the previous zero-shot prediction study [17, 18].
+We first applied $k$-means clustering to the UMAP embedding space of the NSD’s CLIP text features (S3 FigA).
 We set the number of clusters as $40$ based on visual inspection of the UMAP results.
-Based on these clustering results, we performed a hold-out analysis: when predicting samples within a cluster (e.g., the ski cluster), we excluded samples from that cluster in the training set (Figure 6a; hold-out split conditon).
-As a control, we also prepared a naive data-split condition where the training sample size is the same as in the hold-out split condition but allows overlapping semantic concepts.
-When we visualized the properties of predicted features in hold-out analysis by transforming the predicted features into the previous UMAP embedding space (Figure 4a), we observed that the predicted features tend to diverge largely from their original clusters and move into other clusters (Figure 6b).
-Figure 6: Zero-shot cluster identification.
-(a) Conceptual diagram of hold-out analysis.
-For predictions of test samples in a cluster, training samples belonging to the cluster were excluded.
-The naive split condition serves as a control, using the same number of training samples as in the hold-out split condition but allowing semantic concepts to overlap between training and test sets.
-(b) Representative example of feature translation under the hold-out split condition.
-The green and orange points represent training and test samples, respectively.
-The purple diamonds represent the translated (predicted) latent features.
-The gray lines represent the transition from the test sample to the corresponding predicted feature.
-(c) Two types of evaluation of predicted latent features.
-The left side shows the cluster identification accuracy, which compares predicted features to the average features of training samples within each cluster.
-The right side shows the pairwise sample identification accuracy, comparing predicted features with the test sample feature or with features of other test samples.
-(d) Feature prediction results.
-The four surrounding panels show the semantic clusters observed in the UMAP.
-In each panel, the upper graphs show the cluster identification performances for CLIP text and CLIP vision features.
-The expected chance level is $1/40$.
-The lower graphs show the pairwise sample identification performances for CLIP text and CLIP vision features.
-The expected chance level is $1/2$.
-To quantitatively assess the performance of the feature translator, we employed two metrics: cluster identification accuracy and pairwise sample identification accuracy.
-Cluster identification accuracy focuses on evaluating the translator’s ability to generate features that correctly identify the semantic cluster to which a test sample belongs (Figure 6c; left).
+Based on these clustering results, we performed a hold-out analysis: when predicting samples within a cluster (e.g., the ski cluster), we excluded samples from that cluster in the training set (Fig 6A right; hold-out split condition).
+As a control, we also prepared a naive data-split condition where the training sample size is the same as in the hold-out split condition but allows overlapping semantic concepts (Fig 6A left; naive split condition).
+When we visualized the properties of predicted features in hold-out analysis by transforming them into the previous UMAP embedding space (Fig 4A), we observed that the predicted features tend to diverge largely from their original clusters and move into other clusters (Fig 6B).
+Fig 6: Cluster hold-out analysis.
+(A) Hold-out procedure.
+When predicting a test sample from a specific cluster, all training samples from that cluster are excluded.
+The naive split condition serves as a control, using the same number of training samples but allowing cluster overlap between training and test sets.
+(B) Examples of predictions.
+Under the hold-out split condition, gray and dark orange points represent training and test samples, respectively.
+Light orange points indicate translated (predicted) latent features.
+Black lines connect true features to their corresponding predictions.
+(C) Cluster identification procedure.
+The predicted features are compared with the average features of each cluster, and the most similar cluster is selected.
+(D) Cluster identification results.
+Surrounding panels display CLIP text and vision feature performance in the hold-out and naive split conditions of the four representative semantic clusters.
+The chance level is $1/40$.
+(E) Pairwise sample identification procedure.
+Predicted features are compared with the true test sample and another randomly chosen test sample.
+A sample more similar to the prediction is selected.
+(F) Pairwise sample identification results.
+The accuracy means the average winning rate of the true samples against all of the other test samples.
+Results are presented similarly to cluster identification, with a chance level of $1/2$.
+To quantitatively assess the performance of the feature translator, we employed two identification metrics.
+The first metric is cluster identification accuracy.
+Cluster identification accuracy focuses on evaluating the translator’s ability to generate features that correctly identify the semantic cluster to which a test sample belongs (Fig 6C).
 In this analysis, we compare the predicted features of each test sample to the average features of the training samples within each semantic cluster.
-The accuracy is then calculated as the percentage of predicted features that successfully identify the original semantic cluster of their corresponding test samples in the feature space.
-Pairwise sample identification accuracy is a commonly used metric in the evaluation of feature prediction and reconstruction performance (Beliy et al., 2019; Shen et al., 2019b, a; Mozafari et al., 2020; Qiao et al., 2020; Ren et al., 2021; Gaziv et al., 2022; Takagi and Nishimoto, 2023b; Ozcelik and VanRullen, 2023; Ferrante et al., 2023; Scotti et al., 2023; Denk et al., 2023; Koide-Majima et al., 2024).
-This analysis assesses the translator’s ability to generate features that are more similar to the actual test features than to the features of other samples in the test set (Figure 6c; right).
+The accuracy is then calculated as the percentage of predicted features that successfully identify the original semantic cluster of their corresponding test samples in the latent feature space.
+The cluster identification accuracy in the hold-out split condition exhibited a substantial drop compared to the naive split condition across all semantic clusters (Fig 6D).
+Notably, the cluster identification accuracy was frequently 00% in the hold-out split condition (see also S3 Fig.B for all cluster results).
+These results expose a severe limitation of the translator when dealing with novel semantic concepts absent from the training set.
+This suggests that the CLIP feature translator primarily functions as a “classifier;” its prediction (translation) heavily relies on predefined semantic features used in a training phase rather than generalizing to new concepts.
+The second metric is pairwise sample identification accuracy.
+Pairwise sample identification accuracy is a commonly used metric in the evaluation of feature prediction and reconstruction performance [39, 29, 34, 36, 40, 41, 44, 45, 46, 57, 58].
+This analysis assesses the translator’s ability to predict target features that are more similar to the actual test features than to the features of other samples in the test set (Fig 6E).
 For each test sample, we compare its predicted features with two sets of features: the actual test features and the features of another randomly selected sample from the test set.
-The accuracy is then calculated as the average winning rate of the predicted features against all other test samples, indicating the proportion of cases where the predicted features are more similar to the actual test features than to the features of other samples.
-Figure 6d presents a comparison of the identification performance of the feature translator in each semantic cluster between the hold-out and naive split conditions, as measured by two evaluation metrics: cluster identification accuracy and pairwise sample identification accuracy.
-The cluster identification accuracy in the hold-out split condition exhibited a substantial drop compared to the naive split condition across all semantic clusters.
-Notably, the cluster identification accuracy was frequently found to be $0\%$ in the hold-out split condition, exposing a severe limitation of the translator in novel semantic concepts that were absent from the training set (see also Supplementary Figure S2b).
-This finding strongly suggests that the CLIP feature translator functions primarily as a “classifier,” heavily relying on the semantic features learned during the training phase.
-Intriguingly, despite the feature translator’s complete failure to identify semantic concepts in the hold-out split condition, the pairwise sample identification accuracy often surpassed the chance level across the semantic clusters (Figure 6d).
-The insensitivity of the pairwise sample identification accuracy to the translator’s inability to capture semantic concepts can be attributed to its emphasis on comparing the relative similarity between the predicted features and the actual test features, without taking into account the broader categorical context.
-This can lead to an inflated pairwise sample identification accuracy, potentially misrepresenting the true reconstruction performance and the translator’s ability to generalize to novel clusters.
-This observation raises concerns about the adequacy of using identification accuracy as the sole metric for evaluating reconstruction performance, a practice that has been widely adopted in many studies (Beliy et al., 2019; Shen et al., 2019a, b; Mozafari et al., 2020; Qiao et al., 2020; Ren et al., 2021; Gaziv et al., 2022; Takagi and Nishimoto, 2023b; Ozcelik and VanRullen, 2023; Ferrante et al., 2023; Scotti et al., 2023; Denk et al., 2023; Koide-Majima et al., 2024).
-2.1.4 Failed recovery of a stimulus from its latent features
-Finally, we conducted a rigorous evaluation of the generator component, which typically consists of diffusion models in generative AI-based reconstruction methods.
+The accuracy is then calculated as the average winning rate of the predicted features against all other test samples, indicating the proportion of cases where the predicted features are more similar to the actual test features than those of other samples.
+Intriguingly, despite the feature translator’s complete failure to identify semantic concepts in the hold-out split condition, the pairwise sample identification accuracy often surpassed the chance level across the semantic clusters (Fig 6F).
+The insensitivity of the pairwise sample identification accuracy to the translator’s inability to capture semantic concepts can be explained by its focus on relative similarity.
+Specifically, the similarity compares the predicted features to the actual test features, without taking into account the broader general context.
+This can lead to an inflated pairwise sample identification accuracy, potentially misrepresenting the actual reconstruction performance and the translator’s ability to generalize to novel clusters.
+This observation raises concerns about the adequacy of using identification accuracy as the sole metric for evaluating reconstruction performance, a practice that has been widely adopted in many studies.
+Failed recovery of a stimulus from its latent features
+Finally, we conducted a rigorous evaluation of the generator component, which typically consists of diffusion models in text-guided reconstruction methods.
 To ensure that a visual image reconstruction method has the potential to faithfully reproduce an individual’s perceived visual experiences, it is crucial that the method can recover the original images with a high degree of perceptual similarity when the neural translation from brain activity to latent features is perfect.
-However, it has been unclear whether generative AI-based reconstruction methods meet this fundamental requirement.
+However, it has been unclear whether recent text-guided reconstruction methods meet this fundamental requirement.
 To address this question, we performed a recovery check analysis by reconstructing images using the true latent features of target images.
-Instead of using latent features transformed from brain activity, we directly input the latent features derived from the target images into the generator.
-While generative AI-based reconstruction methods generated images that were semantically similar to the target, the resulting images were not perceptually similar to the original (Figure 7a).
-In contrast, iCNN method yielded results that closely resembled the actual target images, demonstrating their superior ability to capture and replicate the original visual content.
-Figure 7: Recovery check.
+Instead of using latent features translated from brain activity, we directly input the latent features derived from the target images into the generator.
+Text-guided methods produced images semantically similar to targets but not perceptually similar (Fig 7A), while the iCNN method yielded results that closely resembled the actual target images.
+These findings suggest that text-guided reconstruction methods may prioritize semantic similarity over perceptual accuracy.
+In contrast, the iCNN method appears to have a superior ability to capture and replicate original visual content, indicating potential advantages in preserving fine-grained visual details.
+Fig 7: Recovery check.
 (A) Reconstruction from the true latent features.
-The left column shows the true images, followed by each method’s recovery results.
-(b) Reconstruction using each component of the Brain-diffuser method.
-The left column shows the true images, followed by the recovery results of each component of the Brain-diffuser method.
-To further investigate the recovery performance, we conducted a recovery check on each latent feature of the Brain-diffuser method (Figure 7b).
+The leftmost column displays the original images.
+Subsequent columns show reconstruction results from the StableDiffusionReconstruction, Brain Diffuser, and iCNN methods.
+(B) Component-wise reconstruction of the Brain-Diffuser method.
+The leftmost column shows the original images.
+The following columns present reconstruction results using individual components of the Brain-Diffuser method (VDVAE, CLIP text, and CLIP vision features).
+Each row represents a different test image.
+The reconstruction results reveal the upper bound of reconstruction performance for each method and highlight potential limitations in the latent feature representations or generative processes.
+To investigate the recovery performance further, we conducted a recovery check on each latent feature of the Brain-Diffuser method (Fig 7B).
 Interestingly, reconstructions from VDVAE features, which are used for generating initial images in the Brain-Diffuser, exhibited a high degree of similarity to the target images.
@@
 Instead, they tend to create images based on their latent features, such as CLIP features, which can lead to a phenomenon known as “hallucination” in the field of generative AIs.
-Hallucination refers to an output that appears plausible but is actually incorrect or misleading, raising concerns about the reliability and accuracy of the model (Rawte et al., 2023).
-These methods seem to prioritize generating semantically similar images rather than faithfully reconstructing the visual content perceived by the individual.
-These findings may provide an explanation for why the text-guided reconstruction methods performed well only on the NSD dataset.
-The results of the case study demonstrated that the text-guided reconstruction methods struggle to reconstruct (Figure 2) or identify (Figure 6) test samples that lie beyond the distribution of the training set.
+Hallucination refers to an output that appears plausible but is actually incorrect or misleading, raising concerns about the reliability and accuracy of the model [59].
+Text-guided reconstruction methods seem to prioritize generating semantically similar images rather than faithfully reconstructing the visual content perceived by the individual.
+The above findings may provide an explanation for why the text-guided reconstruction methods performed well only on the NSD dataset (Fig 2A).
+The results of the case study demonstrated that the text-guided reconstruction methods struggle to reconstruct (Fig 2B) or identify (Fig 6) test samples that lie beyond the distribution of the training set.
 Such limitations suggest that these methods lack true generalization capabilities and are unable to accurately reconstruct novel visual experiences that differ significantly from the examples they were trained on.
-Moreover, even when the test samples belonged to the same distribution as the training set, the translators of the text-guided reconstruction methods had difficulty correctly identifying those test samples (Figure 5).
+Moreover, even when the test samples belonged to the same distribution as the training set, the translators of the text-guided reconstruction methods had difficulty correctly identifying those test samples (Fig 5).
 This observation indicates that the translators may not have learned a sufficiently robust and generalizable mapping between brain activity patterns and the corresponding latent features, further limiting their ability to faithfully reconstruct the perceived visual experiences.
-The case study revealed that the NSD test stimuli are highly similar to the training set (Figure 4), with a significant overlap in their visual and semantic features.
+This case study also revealed that the NSD test stimuli are highly similar to the training set, with a significant overlap in their visual and semantic features (Fig 4).
 Given this similarity, the impressive reconstruction results achieved by the recent text-guided reconstruction methods on the NSD dataset should not be interpreted as evidence of zero-shot reconstruction capabilities.
 Instead, a more plausible interpretation is that these methods primarily function as a combination of “classification” and “hallucination.”
-In this context, the translator component of the text-guided reconstruction methods functions primarily as a classifier, predicting categorical information present in the training data rather than capturing the fine-grained details of the visual experience.
-This limitation may explain why convincing reconstructions can be obtained even from random brain data through post-hoc selection.
-Due to the limited variety in the outputs generated by the reconstruction models, repeated trials can eventually produce images that are semantically and visually similar to the target stimulus.
-The apparent plausibility and semantic consistency of these generated images can be attributed to the capabilities of the diffusion model, which learns to generate realistic-looking images based on the semantic information.
+In this context, the translator component of the text-guided reconstruction methods functions primarily as a classifier, predicting categorical semantic information present in the training phase rather than capturing the fine-grained details of the visual experience.
+This limitation may explain why convincing reconstructions can be obtained even from random brain data through post-hoc selection (Fig 3B).
+Due to the limited variety in the outputs generated by the reconstruction models, repeated trials and post-hoc selection can eventually find images that are semantically and visually similar to the target stimulus.
+The apparent plausibility and semantic consistency of these generated images can be attributed to the capabilities of the diffusion model, which learns to generate realistic-looking images based on semantic information.
 While these images may seem convincing at first glance, they do not accurately reflect the specific visual experience of the individual.
 This phenomenon of hallucination raises serious concerns about the reliability and validity of the text-guided reconstruction methods when evaluated on the NSD dataset.
-2.2 Formal analysis and simulation
-In the above case study, we have identified several issues with these text-guided reconstruction methods and the dataset, especially the cluster structure of CLIP latent features, the lack of diversity in NSD, and the misspecification of the latent representation for image reconstruction, which results in the inability of diffusion models to faithfully recover the original images from their latent features.
+Although we evaluated only two reconstruction methods in this case study, it is important to recognize that any reconstruction methods evaluated only by NSD [46, 48, 60] can potentially have similar classification and hallucination problems.
+Formal analysis and simulation
+Building on the issues identified in our case study, we examine these challenges in a more general context.
+We have identified several issues with these text-guided reconstruction methods and the dataset, including the cluster structure of CLIP latent features, the lack of diversity in NSD, and the misspecification of the latent representation for image reconstruction.
+These issues resulted in the inability of diffusion models to faithfully recover the original images from their latent features.
 However, it is crucial to recognize that the findings of the case study are not merely specific to CLIP, NSD, or diffusion models.
@@
 Thus in this section, we extend the problems identified in the case study into formal analyses and simulations in generalized settings, aiming to provide a more comprehensive understanding of the factors that contribute to the limitations of current reconstruction methods and explore strategies for mitigating these issues.
-2.2.1 Output dimension collapse
-First, we revisit the properties of the translator and highlight the issues that arise when the prediction targets have insufficient diversity.
-Let us consider the general situation of predicting a feature $\mathbf{y}\in\mathbb{R}^{D}$ from a brain activity pattern $\mathbf{x}\in\mathbb{R}^{D}$ using a linear regression model.
-For the training set, we consider brain activity matrix $X_{\mathrm{tr}}\in\mathbb{R}^{N\times D}$ and feature value matrix $Y_{\mathrm{tr}}\in\mathbb{R}^{N\times D}$, where $X_{\mathrm{tr}}$ consists of $N$ samples of $D$-dimensional brain activity vectors $\mathbf{x}$ and $Y_{\mathrm{tr}}$ consists of $N$ feature value vectors $\mathbf{y}$.
-We then train the linear (ridge) regression model using this training data.
-Given a regularization parameter $\lambda$, the weight of ridge regression model is analytically derived as $W=(X_{\mathrm{tr}}^{\top}X_{\mathrm{tr}}+\lambda I)^{-1}X_{\mathrm{tr}}^{\top}Y_{\mathrm{tr}}$, where $I$ is the $D\times D$ identity matrix.
-The predicted feature value $\hat{\mathbf{y}}_{\mathrm{te}}$ for the test brain activity data $\mathbf{x}_{\mathrm{te}}$ can be represented as:
+Output dimension collapse
+Multivariate linear regression models are widely used in constructing decoding and encoding models of the brain.
+These models are often regarded as capable of independent and compositional predictions for each target, as they create separate regression models for individual targets without sharing the weights between them.
+However, this expectation is not generally true.
+This is particularly evident when input variables are shared [33, 35, 34, 44, 45].
+In this section, we demonstrate this assertion by examining the problem of predicting multiple targets using linear regression models with shared inputs.
+Let us consider predicting a feature vector $\mathbf{y}\in\mathbb{R}^{D}$ from a brain activity pattern $\mathbf{x}\in\mathbb{R}^{D}$ using a linear regression model.
+For the training set, we consider brain activity matrix $X_{\mathrm{tr}}\in\mathbb{R}^{N\times D}$ and feature value matrix $Y_{\mathrm{tr}}\in\mathbb{R}^{N\times D}$, where $X_{\mathrm{tr}}$ consists of $N$ samples of $D$-dimensional brain activity vectors $\mathbf{x}$ and $Y_{\mathrm{tr}}$ consists of $N$ feature vectors $\mathbf{y}$.
+We then train a linear (ridge) regression model using this training data.
+Given a regularization parameter $\lambda$, the weight of the ridge regression model is analytically derived as $W=(X_{\mathrm{tr}}^{\top}X_{\mathrm{tr}}+\lambda I)^{-1}X_{\mathrm{tr}}^{\top}Y_{\mathrm{tr}}$ where $I$ is the $D\times D$ identity matrix.
+The predicted feature vector $\hat{\mathbf{y}}_{\mathrm{te}}$ for the test brain activity data $\mathbf{x}_{\mathrm{te}}$ can be represented as:
 $\hat{\mathbf{y}}_{\mathrm{te}}$ $=W^{\top}\mathbf{x}_{\mathrm{te}}$ (1)
@@
 $=Y_{\mathrm{tr}}^{\top}\mathbf{m}=\sum_{i}^{N}m_{i}\mathbf{y}_{\mathrm{tr}}^{(i)},$ (3)
-where ${\bf m}=X_{\rm tr}(X_{\rm tr}^{\top}X_{\mathrm{tr}}+\lambda I)^{-1}\mathbf{x}_{\mathrm{te}}\in\mathbb{R}^{N\times 1}$ and $\mathbf{y}_{\mathrm{tr}}^{(i)}$ is the $i$th training feature vector.
-This transformation indicates that the predicted value is always represented as a linear combination of the features in the training set.
-This property is not limited to ridge regression but is generally applicable to ordinary ridgeless linear regression and related linear models.
+where ${\bf m}=X_{\rm tr}(X_{\rm tr}^{\top}X_{\mathrm{tr}}+\lambda I)^{-1}\mathbf{x}_{\mathrm{te}}\in\mathbb{R}^{N}$, $m_{i}$ is the $i$th element of $\mathbf{m}$, and $\mathbf{y}_{\mathrm{tr}}^{(i)}$ is the $i$th training feature vector.
+This transformation indicates that the predicted value is always represented as a linear combination of the target features in the training set.
+This property is not limited to ridge regression but generally applies to ordinary ridgeless linear regression and related linear models.
 Next, we consider a scenario where the diversity of the target features is small.
-This situation can arise when the feature space exhibits a clustered structure and the training data lacks sufficient diversity, as observed in the case study with the CLIP text features and the NSD dataset.
-When the training feature values have limited diversity, the predicted values from brain activity, which are represented as linear combinations of these features, also become constrained.
-Consequently, the prediction from brain data to features effectively becomes a projection onto a low-dimensional subspace formed by the training data.
-To illustrate this phenomenon, we conducted a simple simulation to examine the distribution of predicted values from a linear regression model trained on clustered features (Figure 8).
-We generated clustered features by sampling from a Gaussian mixture distributions in a high-dimensional space.
-The corresponding brain activity data was generated from a multivariate Gaussian distribution.
-We then trained a linear regression model to predict the clustered feature values from the randomly generated brain data.
-Figure 8: Output dimension collapse.
-The left panel shows the data distribution of source brain activity in the first two dimensions of high-dimensional space.
-The right panel shows the distribution of target latent features in the first two principal component (PC) dimensions.
-If the target features form a cluster structure in the high-dimensional space (left panel green points), the predictions (gray colors) fall into a subspace created by the training clusters, regardless of the input data (right panel).
-The simulation results clearly demonstrate the impact of clustered features on the predicted values.
-The trained model projects arbitrary brain data onto the subspace defined by the training features, resulting in predicted values that are confined to the vicinity of the training clusters.
-This observation highlights the limitation of training linear regression models with clustered features: the trained model’s predictions are inherently constrained by the diversity and structure of the training data.
-This phenomenon, which we refer to as “output dimension collapse,” has important implications for the generalization capability of linear regression models in the context of brain decoding and visual image reconstruction.
-When the training data lacks diversity and forms distinct clusters in the feature space, the translator overly adapts to the subspace formed by the training data, regardless of the potential of the latent feature space.
-Consequently, the translator’s outputs become similar to patterns in the training set, irrespective of the inputs, severely limiting the model’s ability to predict novel or out-of-distribution samples.
-Note that output dimension collapse is not inherently caused by the limitations of linear regression models but rather stems from a lack of diversity in the outputs of the training data.
-This phenomenon can occur regardless of the type of regression model employed, including nonlinear models.
-In fact, nonlinear models may be even more susceptible to output dimension collapse, as the increased flexibility and complexity of these models can potentially exacerbate the issue.
-Output dimension collapse may explain why plausible reconstructions were obtained even from random brain data by merely generating images several times in the case study.
-The NSD’s lack of semantic diversity causes the translator to adapt only to the feature patterns of the training set, restricting its outputs to the subspace formed by the training data.
+This situation can arise when the feature space exhibits a clustered structure and the training data lacks sufficient diversity, as observed in the case study with the CLIP text features and the NSD dataset (Fig 4A).
+When the training features have limited diversity, the predicted values from brain activity, which are represented as linear combinations of these target features, also become constrained.
+Consequently, the prediction from brain data to target features effectively becomes a projection onto a low-dimensional subspace formed by the training data.
+Fig 8: Demonstration of output dimension collapse in feature prediction.
+The left panel shows the distribution of source brain activity data in the first two dimensions of the original high-dimensional space.
+The right panel displays the distribution of target latent features projected onto the first two principal component (PC) dimensions.
+A linear ridge regression model is trained using the training data (gray points in the left and right panels), in which the output features in the training data are constrained to two clusters.
+When presented with test data from the brain space (left panel), the model generates predictions (orange points) in the latent feature space (right panel).
+This visualization demonstrates how the predicted features are constrained to the subspace defined by the training data, highlighting the limitation in generalizing beyond the training data distribution.
+To illustrate this phenomenon, we simulated teacher-student learning, a framework where a “teacher” model generates data based on certain underlying rules, and a “student” model is trained to learn or approximate those rules by observing the generated data.
+We examined the distribution of predicted values from the student linear regression model trained on clustered features generated by the teacher model.
+We generated clustered features by sampling from a Gaussian mixture distribution in a high-dimensional space.
+The corresponding brain activity samples were generated from the latent feature samples multiplied by teacher weight and adding observation noise (See Materials and Methods “Simulation with clustered data”).
+Student weights were obtained by training a linear regression model to predict the clustered feature values from the corresponding brain data.
+We then projected the randomly generated brain samples into latent feature space by learned regression weight.
+To visualize the high-dimensional predicted patterns effectively, we projected the predicted features onto the PCA spaces derived from the training features and showed the first two dimensions.
+The simulation results clearly demonstrate the impact of clustered features on the predicted values (Fig 8).
+The trained linear regression model projects arbitrary brain data onto the subspace defined by the latent features in the training set, resulting in predicted values that are confined to the vicinity of the training clusters.
+This observation highlights the limitation of training linear regression models with clustered features as prediction targets: the trained model’s predictions are inherently constrained by the diversity and structure of the training data.
+We refer to this phenomenon as “output dimension collapse,” where the model’s predictions become confined to a limited subspace of the output feature space.
+It has important implications for the generalization capability of linear regression models in the context of brain decoding and visual image reconstruction.
+When training data lack diversity and form distinct clusters in the feature space, the translator overly adapts to the subspace formed by the training data, regardless of the potential of the latent feature space.
+Consequently, the translator’s outputs become confined to patterns similar to those in the training set, severely limiting the model’s ability to predict novel or out-of-distribution samples.
+Output dimension collapse may explain why plausible reconstructions were obtained even from random brain data by merely generating images several times in the case study (Fig 3).
+The lack of semantic diversity in NSD causes the translator to adapt only to the feature patterns of the training set, restricting its outputs to the subspace formed by the training data.
 As a result, convincing images could be found even from random data through questionable post-hoc selection.
-2.2.2 Simulation with clustered features: What makes prediction compositional
-The case study revealed that the NSD exhibits limited diversity (Figure 4) and poses difficulties for zero-shot prediction (Figure 5 and 6).
-These observations suggest that the translator of CLIP features suffers from output dimension collapse due to the lack of semantic diversity in NSD.
-To explore potential strategies for mitigating output dimension collapse and achieving flexible predictions, we conducted simulation analyses using clustered data to assess generalization performance beyond the training set.
-Our simulation involved teacher-student learning with Gaussian mixture models (see Materials and Methods: “Simulation with clustered data”).
-To imitate the feature translation situation from brain activity, we first generated feature data and assumed a scenario where observation noise affects the brain data.
-To simulate a situation where the dataset has cluster structures and to control diversity effectively, the training feature $\mathbf{y}\in\mathbb{R}^{D}$ was generated from a $D$-dimensional mixture of Gaussians (Figure 9a).
-The brain activity data $\mathbf{x}\in\mathbb{R}^{D}$ was generated from $\mathbf{y}$ multiplied by teacher weight and adding observation noise.
-Figure 9: Simulation analysis of predicting features with cluster structure.
-(a) Illustration of target latent features setting.
+It is crucial to understand that output dimension collapse may not inherently be a limitation of linear regression models.
+This phenomenon can manifest across various regression models, including nonlinear approaches such as multi-layer perceptrons (MLPs) [46] and transformers [48].
+Nonlinear models may be more susceptible to output dimension collapse due to their enhanced flexibility in adapting to training data.
+It is also important to recognize that the mathematical formulation in Eqs 1–3 assumes that all input variables are shared across target variables.
+If each target is predicted from a distinct set of input variables through feature (voxel) selection, the predictions can become more independent, potentially mitigating output dimension collapse.
+This approach has been utilized in the field since its early days, with techniques such as sparse voxel selection and modular modeling [28, 29, 61, 37].
+Simulation with clustered features: What makes prediction compositional?
+The case study revealed that the NSD exhibits limited diversity (Fig 4) and poses difficulties for zero-shot prediction (Fig 5 and Fig 6D).
+These observations suggest that the translator of CLIP features suffers from output dimension collapse due to the lack of semantic diversity in the NSD.
+To explore potential strategies for mitigating output dimension collapse and achieving flexible predictions, we conducted simulation analyses using clustered features to assess generalization performance beyond the training set.
+As in the previous section, our simulation involved teacher-student learning.
+We first generated feature data ${\mathbf{y}}\in\mathbb{R}^{D}$ then made the input brain data ${\mathbf{x}}\in\mathbb{R}^{D}$ by translating ${\mathbf{y}}$ with the teacher weight and added observation noise.
+To simulate a situation where the dataset has cluster structures and to control diversity effectively, the training feature vector $\mathbf{y}\in\mathbb{R}^{D}$ was generated from a $D$-dimensional Gaussian mixture (Fig 9A).
+We trained a ridge regression model on large training data samples and obtained the student weight.
+To simulate the situation where the trained model encounters clusters that are not available at the training phase, we used two types of test samples, in-distribution and out-of-distribution test samples: in-distribution test samples were generated from one of the clusters used in the training set, whereas out-of-distribution (OOD) samples were generated from the novel cluster that is not included in the training set.
+For these two types of predicted features, we calculated the cluster identification accuracy (Fig 6C) by using $C+1$ cluster centers: $C$ centers from the training set and one cluster center of the OOD test set.
+Fig 9: Simulation analysis of predicting features with a cluster structure (A) Illustration of target latent features setting.
 The latent features were generated from Gaussian mixture distributions.
-$\sigma_{r}^{2}$ is a parameter of the sample variance of each cluster.
-$\sigma_{s}^{2}$ controls the distance across clusters.
-In-distribution test samples are generated from the same clusters as in the training set.
+$\sigma_{\mathrm{intra}}$ is a parameter that controls the scale of the samples within a cluster.
+$\sigma_{\mathrm{inter}}$ controls the scale between the cluster centers.
+In-distribution test samples are generated from one of the clusters in the training set.
 Out-of-distribution (OOD) test samples are generated from a novel cluster not included in the training set.
 (B) Cluster identification accuracy.
-The $x$ axis represents the number of clusters used in the training and the $y$ axis represents the cluster identification performance.
-Green line shows the results of In-distribution test samples and orange line shows those for OOD samples.
-Dashed curve indicates the chance level.
-(c) Illustration of a possible scenario where sufficient data diversity is needed to achieve generalizability.
-The upper illustration shows that the training data should cover the whole latent feature space, requiring an exponential order relative to the feature dimension.
-The lower one shows that the training data is covering only the effective axes of the latent feature space, leading to a linear order relative to the feature dimension.
-(d) Relation between latent feature dimension and sufficient number of clusters for generalization.
-The $x$ axis represents the dimension of target latent features.
-The $y$ axis represents the number of clusters that achieve above $0.5$ cluster identification accuracy.
-Three lines indicate the results from different $\sigma_{s}^{2}/\sigma_{r}^{2}$.
-(e) Representative example of generalizability beyond the training set.
-Light green points indicate the in-distribution test samples while light orange points indicate OOD test samples.
-Green and orange points represent the predictions.
-We trained a ridge regression model on large training data samples and obtained the student weight.
-To simulate the situation where the trained model encounters clusters that are not available at the training phase, we generated two types of test samples, in-distribution and out-of-distribution.
-In-distribution test samples were generated using the same probability distribution as the training data, while out-of-distribution (OOD) samples were generated from a different distribution.
-For these two types of predicted features, we calculated the cluster identification accuracy (Figure 6c), using $C+1$ cluster centers, which included $C$ centers from the training set and one additional center for the OOD test set.
-We first examined zero-shot prediction performance as the number of clusters in the training data was increased while fixing the feature dimension $D$ and the cluster variance ratio between each cluster constant (Figure 9b).
-While the cluster identification performance of in-distribution samples were perfect, that of OOD samples showed different patterns depending on the number of clusters in the training data.
-When the number of training clusters was low, cluster identification accuracy was $0\%$.
-This indicates that the behavior of the translator became more similar to a classifier, difficult to generalize beyond the training set, as observed in NSD cases (see also Figure 6d).
+The $x$-axis represents the number of clusters used in the training, and the $y$-axis shows the cluster identification performance.
+The green line represents the results of in-distribution test samples, and the orange line shows the results for OOD samples.
+The dashed curve indicates the chance level.
+(C) Scenarios for achieving generalizability with sufficient data diversity.
+The upper illustration shows the training data covering the whole latent feature space, requiring an exponential order relative to the feature dimension.
+The lower one shows that the training data covers only the effective axes of the latent feature space, leading to a linear order relative to the feature dimension.
+(D) Relation between latent feature dimension and the sufficient number of clusters for generalization.
+The $x$-axis represents the dimension of the target latent features.
+The $y$-axis shows the number of clusters that achieve above $0.5$ cluster identification accuracy.
+The three lines indicate the results from different ratios of $\sigma_{\mathrm{intra}}^{2}/\sigma_{\mathrm{inter}}^{2}$.
+(E) Example of successful prediction beyond the training distribution in a 2D output feature space.
+Dark green points represent the in-distribution target features, with their corresponding light green points showing the model’s predictions.
+Dark orange points denote the OOD target features, with light orange points indicating their predictions.
+Despite the OOD target features (dark orange) not being included in the training clusters, they are accurately predicted (light orange), demonstrating the model’s generalization ability.
+This successful generalization highlights the power of compositional prediction, where the model effectively combines individual dimensions to predict novel feature combinations outside the training distribution.
+We first examined zero-shot prediction performance for different numbers of training clusters (Fig 9B).
+We fixed the feature dimension $D$ and the cluster variance ratio constant.
+While the cluster identification performances of in-distribution test samples were perfect, the performances of OOD test samples showed different patterns depending on the number of clusters in the training data.
+When the number of training clusters was small, the cluster identification accuracy was 00%.
+This result indicates that the behavior of the translator became more similar to a classifier, making it difficult to generalize beyond the training set, as observed in NSD cases (see also Fig 6D).
 On the other hand, as the number of clusters in the training data increased, it became possible to identify the novel clusters, achieving the same performance as in-distribution test samples.
 This observation indicates the importance of the diversity of the training dataset.
-We also emphasize that large numbers of training data do not necessarily address the problem.
-All of these results were obtained with a sufficiently large number of training data, and the number of training clusters was varied while keeping the number of training data fixed.
-Also, we observed qualitatively similar results in increasing the data diversity by controlling the cluster variance ratio (Supplementary Figure S3a).
+We also emphasize that large numbers of training samples do not necessarily address the problem.
+All of these results were obtained with a sufficiently large amount of training data, and the number of training clusters was varied while keeping the amount of training data fixed.
+Also, we observed qualitatively similar results in increasing the data diversity by controlling the cluster variance ratio and keeping the number of dimension and training clusters constant (S4 Fig.A).
 Next, we investigate how diverse the training data needs to be to ensure sufficient generalization.
-There are two possible scenarios for diversifying training samples: either by densely sampling the entire space so that there are no remaining gaps or by uniformly sampling to the extent that it covers the entire dimension of the feature space (Figure 9c).
-The former scenario requires an exponentially larger number of samples relative to the dimension while the latter scenario only requires up to a linear order.
-We sought to reveal which of these two scenarios was more likely to be true by varying the dimensions of the feature space and identifying the number of clusters required for generalization (see also Supplementary Figure S3b).
-Here, we defined the number of clusters required for generalization as the point at which the prediction accuracy of OOD samples exceeds $50\%$.
-The relationship between the dimension of the feature space and the number of clusters necessary for generalization appears to be linear (Figure 9d).
+There are two possible scenarios for diversifying training samples: either by densely sampling the entire target feature space so that there are no remaining gaps (Fig 9C; top) or by uniformly sampling to the extent that it covers the entire dimension of the target feature space (Fig 9C; bottom).
+The former scenario requires an exponentially larger number of samples/clusters relative to the dimension, whereas the latter only requires up to a linear order.
+We sought to reveal which of these two scenarios was more likely to be true by varying the dimensions of the feature space and identifying the number of clusters required for generalization (S4 Fig.
+B for identification accuracy in each condition).
+Here, we defined the number of clusters required for generalization as the point at which the identification accuracy of OOD samples exceeds $50$%.
+The relationship between the dimension of the feature space and the number of clusters necessary for generalization appears to be linear (Fig 9D).
 This finding suggests that achieving generalization does not necessarily require an exponentially large diversity that fills the entire feature space.
-Instead, it suffices to have a number of clusters that cover the effective dimensions within the feature space.
+Instead, it suffices to have a number of clusters that cover the adequate dimensions within the target feature space.
 Although obtaining a large amount of brain data is hard work, it is important for a dataset to contain sufficiently diverse stimuli covering the effective dimensions of the target feature space to achieve zero-shot prediction.
-We also confirmed this phenomenon by a simple and transparent example ($D=2$, Figure 9e).
-The training data covers sufficient axes in the feature space, enabling the prediction of locations not present in the training set.
-Based on this low dimensional intuition, we argue that successful zero-shot prediction requires that training data leads to representations that can serve as a basis spanning the feature space.
+We also confirmed this phenomenon with a simple and transparent example ($D=2$, Fig 9E).
+The training data covers sufficient axes in the target feature space, enabling the prediction of locations not present in the training set.
+Based on this low-dimensional intuition, we argue that successful zero-shot prediction requires that training data leads to representations that can serve as a basis for spanning the target feature space.
 Leveraging such bases effectively enables the model to predict novel samples by predicting each basis and combining them.
-This compositional representation, spanning the target feature space is crucial for zero-shot prediction (Schug et al., 2024) and reconstructing arbitrary visual images from limited brain data.
-2.2.3 Preserved image information across hierarchical DNN layers
-To enable the reconstruction for arbitrary visual images, compositional latent features that can be appropriately mapped into the image space are essential.
-Although it is often assumed that the hierarchical feature representations in DNNs, especially convolutional neural network (CNN), discard pixel-level information through their hierarchical processing with progressively expanding receptive fields, this is not necessarily the case.
-For example, the latent features of auto-encoder models (Hinton and Salakhutdinov, 2006; Kingma and Welling, 2014; van den Oord et al., 2017) can represent images in a low-dimensional space while preserving their reversibility, which is reasonable considering that the output is trained to match the input.
-Furthermore, Mahendran and Vedaldi (2015) showed that input images can be reconstructed with reasonable accuracy even from relatively high-level layers of a DNN designed for an object recognition task.
-It has also been argued that large receptive field sizes do not necessarily impair neural coding capacity as long as the number and density of units remain constant (Zhang and Sejnowski, 1999; Majima et al., 2017).
-These results challenge the notion that higher-level layers in DNNs discard all pixel-level information and highlight the potential for utilizing intermediate representations in visual image reconstruction.
-To further illustrate this point, we performed a recovery check on each intermediate layer of the VGG19 network used in the iCNN methods (Figure 10; see also Figure 7).
-Given the DNN features of a target image, we optimized input pixel values to make the image’s latent features similar to the targets (See Materials and Methods “Recovery check of a single layer by iCNN”).
-We observed that input images can be recovered with reasonable accuracy from relatively high-level layers (around the 11th layers, the 19th layers as a total).
-Furthermore, by introducing image generator networks to add constraints on image statistics, reasonable recovery can be achieved from even higher layers.
-By utilizing a weak image prior (Ulyanov et al., 2017), which contains only information about the structure of images without any prior information on natural images, input images can be recovered from the 13th layer.
-When using an image generator that has learned natural image information (Shen et al., 2019b; Dosovitskiy and Brox, 2016), input images can be recovered even from the 15th layer.
-Figure 10: Recovery check: single layer used in iCNN method.
-For three images, the different optimization algorithm is applied to recover the visual input.
-The left one is pixel optimization, which directly optimizes the pixel values so that the loss between that generated image feature and target image features is minimized.
-The middle one is optimization with week image prior, which optimizes the parameters of image prior models and their latent features simultaneously (see Material and methods).
-The right one is optimization with pretrained image prior, which optimizes the latent features of the fixed image generator model so that the loss between the output image features and target image features is minimized.
-These findings underscore the significance of intermediate DNN representations that can effectively recover the input image in visual image reconstruction studies.
-Conversely, utilizing high-level image features or features from other modalities, such as text annotations, for which the corresponding images are challenging to recover, is not a rational choice for visual image reconstruction tasks.
-Although recent text-to-image models and decoded text features can easily generate images, the outputs should not be interpreted as reconstruction results.
-Instead, it is more reasonable to interpret them as visualizations of decoded semantic information.
-2.2.4 Caveat with evaluation by identification
-Figure 11: Pairwise sample identification with structured data.
-The $x$ axis represents the number of categories in the test set.
-The $y$ axis represents the pairwise sample identification accuracy.
-The lines show the different accuracy $q$ in the situations where the candidate sample is from a different category than those in the test samples.
-The representative similarity matrices corresponding to these conditions are shown in the figure above.
-Pairwise identification has been a standard metric for evaluating latent feature decoding (Horikawa and Kamitani, 2017) or reconstruction performance (Beliy et al., 2019; Shen et al., 2019a, b; Mozafari et al., 2020; Qiao et al., 2020; Ren et al., 2021; Gaziv et al., 2022; Takagi and Nishimoto, 2023b; Ozcelik and VanRullen, 2023; Ferrante et al., 2023; Scotti et al., 2023; Denk et al., 2023; Koide-Majima et al., 2024).
-However, our analysis revealed that even with difficulties in accurately identifying semantic clusters the test samples belong to, pairwise identification performance still surpassed chance levels (Figure 6d).
+This compositional representation, spanning the target feature space, is crucial for zero-shot prediction [62] and reconstructing arbitrary visual images from limited brain data.
+Caveat with evaluation by pairwise identification
+Pairwise identification has been a standard metric for evaluating latent feature decoding [19] or reconstruction performance [39, 29, 38, 34, 36, 40, 41, 44, 45, 46, 57, 58].
+However, our analysis revealed that even with difficulties in accurately identifying semantic clusters the test samples belong to, the pairwise identification performance still surpassed its chance level (Fig 6F).
 Here we critically examine this metric and demonstrate that significant results can be easily obtained when the target or predicted features exhibit certain structures.
-Pairwise identification is calculated as the accuracy with which the predicted features (reconstructed image or its features) can correctly identify the true ones, in pairs consisting of a true sample and one of the remaining samples in the test set.
-If the candidate feature belongs to the same category as the true feature, we can expect the identification to be difficult.
-Conversely, if the candidate feature belongs to a different category from the true feature, identification would become easier.
-Here we assume the test set comprises $k$ categories and all test samples are equally distributed across each category for simplicity.
-We model the situation mentioned above as the expected identification accuracy is set to chance level (i.e., 0.5) when a candidate belongs to the same category as the true feature.
-In contrast, the expected identification accuracy when a candidate belongs to a different category from the true feature is set to $q$, ranging between 0.5 to 1 to mimic the ease of identification across categories.
-Assuming the number of test samples is large enough, the pairwise identification becomes
+Pairwise identification is calculated as the accuracy with which the predicted features (either the output of a translator or features extracted from the output of a generator) can correctly identify the corresponding true ones, in pairs consisting of a true sample and one of the remaining samples in the test set.
+If the candidate sample belongs to the same category as the true sample, the identification is expected to be difficult.
+Conversely, if the candidate sample belongs to a different category from the true one, identification becomes easier.
+Fig 10: Expected pairwise sample identification performance in categorically structured data.
+The $x$-axis represents the number of categories in the test set.
+The $y$-axis represents the pairwise sample identification accuracy.
+Different lines represent various levels of accuracy ($q$) in distinguishing samples from different categories.
+The chance level of $0.5$ is represented by the bottom line.
+Above the graph, hypothetical similarity matrices are shown to illustrate the categorical structures of the samples, where the color (white/black) indicates similarity between samples.
+The block diagonal structure reflects the categorical nature of the data.
+Samples within the same category are assumed to be indistinguishable, resulting in a pairwise identification accuracy of $0.5$ (chance level).
+Samples from different categories can be distinguished with a pairwise identification accuracy of $q$, where q varies between $0.5$ and $1$.
+In this model, we assume the test set comprises $k$ categories with test samples equally distributed across each category for simplicity.
+We model the situation mentioned above by setting the expected identification accuracy at the chance level (i.e., $0.5$) when a candidate sample belongs to the same category as the true one.
+Conversely, when the candidate sample belongs to a different category than the true sample, we set the expected identification accuracy to a parameter $q$, ranging from 0.5 to 1.
+This parameter $q$ reflects the ease of identification across categories.
+Assuming the number of test samples is sufficiently large, the pairwise identification accuracy $\mathrm{Acc}$ becomes
 $\mathrm{Acc}=\frac{1}{k}\cdot 0.5+\left(1-\frac{1}{k}\right)\cdot q$ (4)
 (See Materials and Methods “Expected identification accuracy in imprecise reconstructions” for the derivation).
-Figure 11 shows the line plots of expected values of pairwise identification accuracy as a function of the number of categories in the test set, while presenting the possible similarity structure of the test set.
-Even with failure to identify within categories completely and success only between two categories, pairwise identification accuracy can still reach a high value up to $75\%$.
-This finding highlights a potential limitation of the pairwise identification metric, as judging reconstruction performance solely based on it can be misleading.
-2.2.5 How are we fooled by hallucinations of generative AIs
-Figure 12: The dangerous interpretation of photo-like appearances as accurately reflecting the perceived visual image.
-The left Venn diagram represents the interpretation of photo-like reconstructions.
-$T$ denotes the event where the generator’s output accurately reflects the visual image and $R$ denotes the event where the generator’s output has a photo-like appearance.
-Introducing generative AIs (diffusion models) might increase the overlap between $T$ and $R$; however, it may also significantly increase the areas unrelated to $T$.
-The right equation represents the probability that the generator’s output accurately reflects the visual image given that the generator’s output has a photo-like appearance as inferred from the left Venn diagram.
-Generative AIs have recently made remarkable progress, with models now capable of producing high-resolution and hyper-realistic images from text input (Ramesh et al., 2021) or generating text of a quality indistinguishable from human-written content (Brown et al., 2020).
+Fig 10 illustrates the relationship between pairwise identification accuracy and the number of categories in the test set through line plots of expected values.
+Inset figures depict the underlying similarity structure of the test set.
+Notably, even when identification within categories fails completely and succeeds only between two categories, the pairwise identification accuracy can still reach a high value of up to $75$%.
+This finding underscores a significant limitation of the pairwise identification metric: evaluating reconstruction performance based solely on whether this metric is above the chance level may lead to misleading conclusions.
+Preserved image information across hierarchical DNN layers
+The reconstruction of arbitrary visual images requires compositional latent features that can be effectively mapped into the image space.
+Hierarchical DNN features are often employed as such representations.
+However, do these features truly map to the image space?
+A common narrative suggests that the hierarchical processing discard pixel-level information through progressively expanding receptive fields.
+This view, however, is not entirely accurate.
+Fig 11: Image recovery check for hierarchical latent features.
+The images show the ability to recover visual inputs from different layers of the VGG19 network.
+For each of the three sample images, three different optimization methods were applied.
+The left column shows results from pixel optimization, which directly optimizes pixel values to minimize the loss between the generated image features and the target image features [65].
+The middle column displays optimization with a weak image prior, simultaneously optimizing both the weight parameters of the image prior models and their latent features [68].
+The right column presents optimization with a pre-trained image prior, which optimizes the latent features of a parameter-fixed image generator model to minimize the loss between the output image features and the target image features [69].
+Rows correspond to different layers of the VGG19 network, progressing from earlier layers at the top to later layers at the bottom.
+This progression illustrates how image information is preserved or lost at different stages of the network.
+Reasonable image recovery is possible even from relatively high-level layers of the VGG19 network, challenging the notion that higher layers in deep neural networks discard image-level visual information.
+For example, the latent features of auto-encoder models [63, 51, 64] can represent images in a low-dimensional space while preserving their reversibility, which is reasonable considering that the model’s output is trained to match the input.
+Furthermore, Mahendran and Vedaldi [65] showed that input images can be reconstructed with reasonable accuracy even from relatively high-level layers of a DNN designed for an object recognition task.
+It has also been argued in the neuroscience field that large receptive field sizes do not necessarily impair neural coding capacity as long as the number and density of units remain constant [66, 67].
+These results challenge the notion that higher-level layers in DNNs discard all pixel-level information.
+To further illustrate this point, we performed a recovery check on each intermediate layer of the VGG19 network used in the iCNN methods (Fig 11; see also Fig 7).
+Given the DNN features of a target image, we optimized input pixel values to make the image’s latent features similar to the targets (See Materials and Methods “Recovery check of a single layer by iCNN”).
+We observed that input images can be recovered with reasonable accuracy from relatively high-level layers (around the 11th layer out of the total 19th layers).
+Furthermore, by introducing image generator networks to add constraints on image statistics, reasonable recovery can be achieved from even higher layers.
+By utilizing a weak image prior [68], which contains only information about the structure of images without any prior information on natural images, input images can be recovered from the 13th layer.
+When using an image generator that has learned natural image information [29, 69], input images can be recovered even from the 15th layer.
+These observations suggest that, even when feature representations shift from lower to higher levels through hierarchical processing, pixel-level information is not largely discarded; rather, much of the input information is preserved across almost the entire level.
+This perspective highlights the potential for utilizing intermediate DNN representations in the visual image reconstruction study as the generator should recover the original stimulus from the true latent features (see also “Failed recovery of a stimulus feature from its latent features” in the case study section).
+With this insight in mind, exploring which representations have compositional representation and are more predictable from brain activity will be a critical step in advancing visual image reconstruction.
+Conversely, utilizing high-level image features, such as the output of DNNs, or features from other modalities, such as text annotations, is not a rational choice for visual image reconstruction tasks.
+These latent features make it challenging to recover the corresponding input image (Figure 7B) and are insufficient as surrogate representations of perceived visual images.
+Although recent text-to-image models and predicted text features can easily generate images, the outputs should not be interpreted as reconstruction results.
+Instead, it is more appropriate to view them as visualizations of decoded semantic information.
+Such visualizations are valuable for illustrating purposes.
+However, it is crucial to recognize the significant distinction between semantic visualization and reconstruction.
+How are we fooled by hallucinations of generative AIs?
+Generative AIs have recently made remarkable progress, with models now capable of producing high-resolution and realistic images from text input [23] or generating text of a quality indistinguishable from human-written content [24].
 However, due to the complex internal structure of these models and the vast amounts of data they are trained on, we are often fooled by the outputs of generative AIs.
-For instance, when searching for an unfamiliar topic using a large language model (LLM) in our daily life, we may not realize that the model is creating false concepts.
-Someone may have mistaken generative AI show real-time response in the promotion, despite the fact that these promotions were edited (Coldewey, 2023).
-We might also believe that these models are unbiased or can represent all possible viewpoints, even though they inherently contain biases from their training data and developers (Messeri and Crockett, 2024).
-As we have observed that the generative AI-based reconstruction methods exhibit photo-like appearances but poor generalizability (Figure 2), could similar issues be occurring in visual image reconstruction studies as well?
+For instance, when searching for an unfamiliar topic using a large language model (LLM) in our daily lives, we may not realize that the model is creating false concepts.
+This is likely a result of generative AIs being trained on a large amount of data and producing highly coherent and contextual responses.
+We may also believe these models are unbiased and can represent all possible data points, even though they inherently contain biases from their training data and developers [70].
+As we have observed, the generative AI-based reconstruction methods exhibit realistic appearances but poor generalizability (Fig 2).
+Could similar issues occur in visual image reconstruction studies as well?
 The goal of visual image reconstruction is to generate images from brain activity that precisely mirror visual experience.
-However, there appears to be a prevalent focus among the general public, reviewers, and even researchers on achieving outputs that are as photorealistic as possible, rather than emphasizing the accuracy of these reconstructions.
-This shift in focus raises questions about the extent to which these photorealistic reconstructions truly represent the actual visual experiences.
-Traditionally, we have held two beliefs: 1) generating photo-like images from brain activity is challenging, and 2) if the reconstruction pipeline effectively captures the brain representation under natural image perception, the model’s output should also appear naturalistic or photo-like.
-Based on these beliefs, we tend to consider photo-like reconstructions as an indication of accurately reflecting the actual visual experience.
-This heuristic can be formalized as follows:
-Assumption 1: $\Pr[R]\ll\Pr[\bar{R}]\approx 1,$ (5)
-Assumption 2: $\Pr[R\mid T]\approx 1,$ (6)
-Conclusion: $\Pr[T\mid R]\text{ is high},$ (7)
-where $T$ represents the event of the model’s output truthfully reflects the visual image and $R$ represents the event of the model’s output has a realistic appearance.
-$\bar{R}$ represents a complementary event of $R$.
-$\Pr[T\mid R]$ is the probability that how likely we should conclude that the model’s output actually reflects the subject’s visual experience based solely on the fact that the output is photo-like.
-In fact, this heuristic is reasonable when the above two assumptions are true and $\Pr[T]$ is not extremely low.
-However, recent developments in generative AIs, such as diffusion models, have made it easily to produce convincing photo-like outputs, subverting the first assumption i.e, $\Pr[R]\gg\Pr[\bar{R}]$.
-Consequently, it becomes invalid to infer that the visual images are accurately reflected in the generator outputs solely because they appear photo-like.
-Rather, as shown in Figure 12, the probability $\Pr[T\mid R]$ may become smaller as the generative AIs produce more convincing outputs (see also Figure 7).
+However, there appears to be a prevalent focus among the general public, reviewers, and even researchers on achieving as realistic outputs as possible, rather than emphasizing the accuracy of these reconstructions.
+This shift in focus raises questions about the extent to which these realistic reconstructions truly represent the actual visual experiences.
+Traditionally, we have held two beliefs: (1) generating realistic images from brain activity is challenging, and (2) if the reconstruction pipeline effectively captures the brain’s representation under natural image perception, the model’s output should also appear realistic.
+Based on these beliefs, realistic reconstruction is often considered as an indication of accurately reflecting the actual visual experience.
+Fig 12: The illustration of how realistic appearances can be misleading in visual reconstruction.
+On the left, Venn diagrams depict the relationship between two events: $R$, where the output has realistic appearances, and $T$, where the output truthfully reflects the visual experience.
+$T$ and a small dashed circle of $R$ depict a common pre-generative AI heuristic: the assumption that realistic appearance implies truthful reconstruction.
+Dashed arrows illustrate the effect of generative AI models (e.g., diffusion models) on this relationship.
+These models can potentially increase the overlap between $T$ and $R$, but may also expand areas of $R$ unrelated to $T$.
+On the right, a probabilistic interpretation of the relationship is presented.
+The equation represents the probability that the generator’s output truthfully reflects visual experience ($T$), given that the generator’s output has a realistic appearance ($R$).
+This conditional probability is derived from the areas in the Venn diagram.
+The expansion of $R$ without a proportional increase in $T$ can lead to a decrease in $\Pr[T\mid R]$, emphasizing the need for careful evaluation of reconstruction results beyond just assessing their realistic appearance.
+To formalize this heuristic reasoning, let us first define the two events $R$ and $T$, where $R$ represents the event that the output of reconstruction models has a realistic appearance, and $T$ represents the event that the model’s output truthfully reflects the visual image.
+The first belief, concern about the difficulty of generating realistic images from brain activity, is expressed as $\Pr[R]\ll\Pr[\bar{R}]\approx 1$ where $\bar{R}$ represents the complementary event of $R$.
+The second belief, that reconstruction achieves a realistic appearance if the pipeline effectively captures the brain’s representation under natural image perception, is expressed as $\Pr[R\mid T]\approx 1$.
+This conditional probability implies that the likelihood of the model’s output being realistic is high, given that the model truthfully captures the subject’s visual experience.
+The heuristic reasoning that realistic reconstructions indicate an accurate reflection of the actual visual experience can thus be represented as “$\Pr[T\mid R]\text{ is high}$.” This heuristic is, in fact, reasonable as it can be derived from Bayes’ rule, assuming the above two beliefs hold true, and that $\Pr[T]$ is not extremely low.
+However, recent developments in generative AIs, such as diffusion models, have made it easy to produce convincing, realistic outputs, subverting the first assumption, i.e., $\Pr[R]\gg\Pr[\bar{R}]$.
+Consequently, it becomes invalid to infer that the visual images are accurately reflected in the generator outputs solely because they appear realistic.
+Rather, as shown in Fig 12, the probability $\Pr[T|R]$ may become smaller as the generative AIs produce more convincing outputs (see also Fig 7).
 This perspective emphasizes the need for careful evaluation of reconstruction performance, considering the possibility of hallucinations by generators.
-While pursuing photo-like reconstructions to improve reconstruction fidelity is undoubtedly important, it would be counterproductive to obsess over naturalistic appearance to the point of neglecting the original goal of reconstructing perceived visual images.
-3 Discussion
-In this study, we critically examined recent generative AI-based visual image reconstruction methods to assess their true capabilities and limitations.
+While pursuing realistic reconstructions to improve reconstruction fidelity is undoubtedly important, it would be counterproductive to obsess over naturalistic appearance to the point of neglecting the original goal of reconstructing perceived visual images.
+Discussion
+In this study, we critically examined generative AI-based visual image reconstruction methods to assess their true capabilities and limitations.
 Our primary goals were to (1) investigate the performance of these methods on different datasets, (2) identify potential issues and pitfalls in their methodology and evaluation, and (3) provide insights and recommendations for future research in this field.
 We conducted a case study focusing on text-guided reconstruction methods and their validation on the Natural Scene Dataset (NSD).
-Our findings revealed several concerns, including the failure to replicate the reconstruction performance on a different dataset, the use of problematic post-hoc image selection procedures, the lack of diversity and limited number of clusters in the NSD stimulus set, the failure of zero-shot prediction by the translator component, and the inability to accurately recover original stimuli by the generator component.
+Our findings revealed several concerns, including the failure to replicate the reconstruction performance on a different dataset, the use of problematic post-hoc image selection procedures, the lack of diversity and a limited number of clusters in the NSD stimulus set, the failure of zero-shot prediction by the translator component, and the inability to recover original stimuli by the generator component accurately.
 Formal analysis and simulations further demonstrated the phenomenon of output dimension collapse, the importance of compositional representations for achieving zero-shot prediction, and the potential pitfalls of relying solely on identification metrics to evaluate reconstruction performance.
-Moreover, we highlighted that photo-like appearance does not necessarily imply accurate reflection of the perceived visual images.
+Moreover, we highlighted that a realistic appearance does not necessarily imply an accurate reflection of the perceived visual images.
 Based on these findings, we argue that the reconstructions from the recent text-guided reconstruction methods are, in large part, the result of a combination of classification and hallucination.
@@
 While our study critically examined the limitations of recent text-guided, diffusion-based methods for visual image reconstruction, it is important to acknowledge that these approaches provide new and potentially promising directions for brain decoding research.
-Although we emphasized the importance of achieving zero-shot prediction, it is crucial to recognize that most brain decoding studies focus on classification tasks, which, while not zero-shot, have provided insights into neural representations (Kamitani and Tong, 2005; Haxby et al., 2001).
+Although we emphasized the importance of achieving zero-shot prediction, it is crucial to recognize that most brain decoding studies focus on classification tasks, which, while not zero-shot, have provided insights into neural representations [1, 2].
 Moreover, the visualization of semantic contents (e.g., the supplementary movies in Horikawa et al.
-(2013)), can have significant utility in visual communication, even if it does not constitute zero-shot prediction.
+[4]), can have significant utility in visual communication, even if it does not constitute zero-shot prediction.
 It is also worth noting that the individual components of these text-guided, diffusion-based methods are already being utilized in various brain decoding applications.
-For instance, text latent features derived from deep neural networks and large language models (LLMs) have shown promise in analyzing semantic information from brain activity (Tang et al., 2023; Caucheteux et al., 2023; Zhou et al., 2024).
-Furthermore, diffusion models are not limited to text-to-image generation; they can be trained to generate images from visual latent features, as demonstrated by Cheng et al.
-(2023), enabling the successful reconstruction of subjective experiences.
-By leveraging these components and exploring their potential synergies, researchers can continue to push the boundaries of brain decoding and visual image reconstruction, while being mindful of the challenges and limitations highlighted in our study.
+For instance, text latent features derived from deep neural networks and large language models (LLMs) have shown promise in analyzing semantic information from brain activity [71, 72, 73].
+Furthermore, diffusion models are not limited to text-to-image generation; they can be trained to generate images only from visual latent features, as Cheng et al.
+[32] demonstrated in successful reconstruction of subjective experiences.
+By leveraging these components and exploring their potential synergies, researchers can continue to push the boundaries of brain decoding and visual image reconstruction, while keeping in mind the challenges and limitations highlighted in our study.
 The recent trend of collecting and sharing large-scale visual neural datasets, such as those by Hebart et al.
-(2023) and Xu et al.
-(2024), is a welcome development in the field of neuroscience.
+[14] and Xu et al.
+[15], is a welcome development in the field of neuroscience.
 These datasets provide valuable resources for researchers to investigate brain function and advance our understanding of visual processing.
-The NSD is a particularly notable example, as it was created with the goal of extensively sampling brain responses to a wide range of natural visual stimuli (Allen et al., 2022; Naselaris et al., 2021).
-The NSD has been widely utilized in various studies (Prince et al., 2022; Conwell et al., 2023; Gifford et al., 2023), demonstrating its value to the research community.
+The NSD is a particularly notable example, as it was created with the goal of extensively sampling brain responses to a wide range of natural visual stimuli [12, 13].
+The NSD has been widely utilized in various studies [74, 75, 76], demonstrating its value to the research community.
 While our results suggest that the semantic and visual diversity of the NSD stimuli may not be as high as initially thought, and there is substantial overlap between the training and test sets provided by the NSD authors, this does not diminish the overall importance and usefulness of the dataset.
-However, to fully leverage the NSD and other publicly available large-scale datasets for developing generalizable and zero-shot prediction models, it is crucial to carefully consider the data split between training and test sets.
-When aiming for generalizable predictions, such as in visual image reconstruction, researchers should verify whether there are significantly similar stimuli included in both the training and test sets.
+However, to fully leverage the NSD and other publicly available large-scale datasets for developing generalizable and zero-shot prediction models, it is crucial to consider the data split between training and test sets carefully.
+When aiming for generalizable predictions, such as in visual image reconstruction, researchers should verify whether significantly similar stimuli are included in both the training and test sets.
 If necessary, redesigning the training and test split can ensure that the models are tested on their ability to predict novel stimuli.
-Moreover, recent advancements in functional alignment and inter-site neural code conversion methods (Haxby et al., 2011; Yamada et al., 2015; Wang et al., 2024) hold promise for combining datasets from different sources, enabling truly larger-scale data analysis in neuroscience.
+Moreover, recent advancements in functional alignment and inter-site neural code conversion methods [77, 78, 79] hold promise for combining datasets from different sources, enabling truly larger-scale data analysis in neuroscience.
 These techniques allow researchers to align brain activity patterns across individuals and measurement sites even when stimuli are not shared across datasets.
 By leveraging these methods, researchers can pool data from various sources, increasing the sample size and diversity of the combined dataset, mitigating the limitations of individual datasets, and enhancing the development of generalizable and zero-shot prediction models.
-Investigating neural responses to natural stimuli is a highly valuable approach to understand brain function and representation (Nastase et al., 2020; Hasson et al., 2020).
-As our brains have developed while being exposed to natural scenes, it is crucial to use natural stimuli especially in model training.
+Investigating neural responses to natural stimuli is a highly valuable approach to understanding brain function and representation [80, 81].
+As our brains have developed while being exposed to natural scenes, it is crucial to use natural stimuli, especially in model training.
 However, we should not forget that we are also capable of perceiving non-natural stimuli like artificial images.
-We would like to emphasize that there are the potential pitfalls when relying too heavily on evaluations based solely on natural stimuli.
-With the increasing scale of neural data and the growing complexity of analysis pipelines, there is a risk that the learned mappings may produce unexpected shortcuts, just as we have demonstrated that the generative AI-based reconstruction methods exploited the semantic and visual overlap between training and test sets.
-Drawing inspiration from comparative and developmental psychology, where researchers often employ simple stimuli and tasks to measure cognitive abilities in infants for better experimental control and precise inferences (Kominsky et al., 2022; Messeri and Crockett, 2024) , the evaluation of visual image reconstruction should not be limited to complex natural stimuli alone.
+We would like to emphasize that there are potential pitfalls when relying too heavily on evaluations based solely on natural stimuli.
+With the increasing scale of neural data and the growing complexity of analysis pipelines, there is a risk that the learned mappings may produce unexpected shortcuts, just as we have demonstrated that the text-guided reconstruction methods exploited the semantic and visual overlap between training and test sets.
+In the field of comparative and developmental psychology, researchers often prioritize using not natural but simple stimuli for better experimental controls and more precise inferences about infants’ cognitive abilities [82, 83].
+Drawing inspiration from this approach, we argue that the evaluation of visual image reconstruction should not be limited to complex natural stimuli alone.
 While natural stimuli are essential for ensuring ecological validity and understanding how the brain processes real-world information, it is equally important to assess the performance in controllable and transparent manners.
-We discussed output dimension collapse and related issues with clustered data found in natural image features.
-It is important to note that these issues can arise in various situations.
-Several studies have used brain data where the stimuli shared the category information between the training and test sets.
+We have observed that an inappropriate split between training and test stimulus sets can lead to spurious reconstruction, invalidating zero-shot predictions.
+It is crucial to recognize that these problems can arise in various research contexts.
+Several studies have used data where the stimuli shared the category information between the training and test sets, potentially compromising the validity of their results.
 Kavasidis et al.
-(2017) collected a dataset of EEG signals recorded during natural image perception.
-Their visual stimuli consisted of 2,000 images selected from 40 object categories in ImageNet (50 images per category) and the test set contains the same categories that are included in the training set.
+[84] collected a dataset of EEG signals recorded during natural image perception.
+Their visual stimuli consisted of $2,000$ images selected from $40$ object categories in ImageNet ($50$ images per category) and the test set contained the same categories that are included in the training set (See also [85, 15] for other issues with the data set).
 Denk et al.
-(2023) attempted to develop music reconstruction methods from fMRI activity patterns.
-Their music stimuli consisted of 540 music pieces selected from 10 music genres and the test set contains the same genres as the training set (Nakai et al., 2021).
+[57] attempted to develop music reconstruction methods from fMRI activity patterns.
+Their music stimuli consisted of $540$ music pieces selected from $10$ music genres, and the test set contained the same genres as the training set [86].
 Orima et al.
-(2024) attempted to reconstruct perceived texture images from EEG signals.
-Their texture stimuli consisted of 191 images of 21 natural textures and they performed a reconstruction analysis with a leave-one-out manner.
-It should be carefully examined whether these studies may suffer from output dimension collaplse, merely decoding category information from brain activity and generating stimuli based on this classified category information, similar to our interpretation of the text-guided reconstruction methods.
-Note that for other issues in the data from Kavasidis et al.
-(2017), see Li et al.
-(2018); Xu et al.
-(2024).
-Our inspection of the data from Nishimoto et al.
-(2011) revealed that many frames in the test movie stimuli were almost identical to those in the training set (Supplementary Figure S4).
-This is presumably because temporally adjacent video frames were split between the training and test stimuli.
-While our preliminary analysis of their motion energy features did not exhibit peculiar clustering (note also that their model was an encoding model where the output is brain activity), caution should be exercised if the dataset is used for models that extract other features of the videos.
+[87] attempted reconstructing perceived texture images from EEG signals.
+Their texture stimuli consisted of $191$ image patches extracted from $21$ natural textures, and they performed a reconstruction analysis in a leave-one-out manner.
+It should be carefully examined whether these studies may suffer from output dimension collapse, merely decoding the information from trained categories from brain activity.
+Experimental design for train-test stimuli setups requires careful consideration.
+Dado et al.
+[88] conducted an image reconstruction analysis from the multi-unit activity of a macaque using images generated from latent features of generative models.
+Importantly, all of the test stimuli were generated from the averaged latent features of the categories used in their training phase.
+While the authors addressed potential biases by redesigning the training and test split, researchers should exercise prudence when utilizing the dataset.
+Our inspection of the movie stimuli from Nishimoto et al.
+[89] revealed that many frames in the test movie stimuli were nearly identical to those in the training set (S5 Fig).
+This similarity likely results from temporally adjacent video frames being split between the training and test stimuli.
+While our preliminary analysis of their latent features (motion energy features) did not show unusual clustering (note also that the study presents a way of retrieving movie instances via a brain encoding model, rather than reconstruction in the current sense), caution is required when using the stimulus set to extract other types of features [90, 91].
+Claims of predicting arbitrary instances or achieving zero-shot prediction warrant thorough scrutiny.
+There is a widespread practice of using test data for fine-tuning, which can also be a questionable procedure.
+Some studies have proposed methods that involve fine-tuning models using the entire test brain data, but not the test stimuli, following initial training with the training stimuli and brain data [39, 43].
+While the absence of test labels (stimuli) during fine-tuning may help avoid obvious overfitting, these approaches still treat test brain data as previously observed information, rather than as a proxy for novel data potentially encountered in real-world situations.
+Consequently, this practice violates test data independence, making it difficult to evaluate the model’s actual generalization ability.
+Although such procedures may effectively improve performance on existing benchmarks or competitions, it is crucial to recognize that incorporating test data information during any stage of training can undermine the validity of neuroscientific claims and limit the real-world applicability of the methods.
+The issue of double dipping, which refers to the use of the same dataset for both data/variable selection and selective analysis (inference and prediction), has been widely recognized in neuroscience [92, 93].
+In classification tasks, while selecting input variables using test data is problematic, using the same output labels (target variables) for training and test sets is not inherently flawed, given that the nature of classification assumes consistent categories across datasets.
+However, the challenges we address in this study, although related to double dipping, present distinct concerns.
+In evaluating zero-shot prediction performance, the mere similarity of test labels (features) can lead to overestimating model performance.
+It is crucial to distinguish between conventional double dipping and the current issues we identify, such as output dimension collapse in zero-shot prediction scenarios.
+These emerging challenges necessitate careful consideration not only of input data independence but also of structural similarities between training and test sets in the output space.
+Additionally, the independence of test stimuli from the model training processes requires careful examination, particularly when using pre-trained deep neural network (DNN) models and foundation models like CLIP or diffusion models.
+These models are typically trained on vast amounts of data available on the internet [22, 24, 50], which likely includes public datasets such as MS-COCO [94] that are used in the NSD.
+This overlap raises potential concerns about the true independence of test stimuli, as we cannot rule out the possibility that these pre-trained models have acquired representations specifically tailored to the stimuli used for the model training.
+To address these issues and ensure a more rigorous evaluation of model generalization capabilities, researchers may consider using test stimuli that are not publicly available on the internet.
+This could include self-created stimuli or carefully curated datasets that have not been used in the training of widely used AI models [29, 32].
+Such an approach would provide a more stringent test of a model’s ability to generalize to truly novel inputs.
 One of the remaining challenges in visual image reconstruction is the development of metrics for evaluating the quality and accuracy of the reconstructed images.
@@
 Following this, quantitative metrics should be employed for a more objective, high-throughput evaluation.
-However, as our analysis has suggested, it can be misleading to evaluate reconstruction by heavily relying on identification performance based on the relative similarity among alternatives (Koide-Majima et al., 2024).
+However, as our analysis has suggested, it can be misleading to evaluate reconstruction by heavily relying on identification performance based on the relative similarity among alternatives [58].
 Even in cases where the reconstructed images only capture superficial information, such as categories or overall brightness, identification metrics can still be high.
@@
 It is crucial to develop more appropriate similarity metrics that can accurately measure the perceptual similarity between the reconstructed and original images.
-One promising approach is to leverage image quality assessment (IQA) techniques from the computer vision field (Fu et al., 2023; Ding et al., 2020).
+One promising approach is to leverage image quality assessment (IQA) techniques from the computer vision field [55, 95].
 These techniques are designed to quantify the perceptual quality of images and can be adapted to the specific requirements of visual image reconstruction.
-Visual image reconstruction methods have gained attention not only from neuroscientists but also from the general public and policymakers, sparking discussions about their potential applications and risks (UNESCO, 2023).
+Our findings have significant implications for research integrity and responsible dissemination of scientific results.
+Visual image reconstruction methods have gained attention not only from neuroscientists but also from the general public and policymakers, sparking discussions about their potential applications and risks [8].
 These stakeholders often contemplate the possibilities of seamless information communication through the brain, such as in brain-machine interfaces (BMIs), or the dangers of unauthorized access to private information from brain activity.
@@
 However, current technology and analysis methods fall short of these expectations.
-For instance, most reconstruction methods analyze previously acquired brain data offline.
-Additionally, the brain data used for reconstructing images are often averaged over multiple presentations of the test image, with only a few studies demonstrating single-trial reconstruction results (Miyawaki et al., 2008; Cheng et al., 2023).
-Further, It has been argued that subject cooperation is essential for reliably training and testing decoding models (Tang et al., 2023).
-Public expectations are thus set too high, and it is challenging to meet these demands quickly.
-It is essential to explicitly state these limitations to avoid disappointment and prevent governments or companies from making misguided decisions.
-Key limitations include the lack of real-time analysis, reliance on averaged brain data, the necessity of subject cooperation, and the need for further model fine-tuning using the test brain dataset (Beliy et al., 2019; Chen et al., 2023a).
-It is crucial to avoid making overly optimistic claims about the ability to reconstruct arbitrary images, as the applicability can be highly dependent on specific training data and conditions.
-4 Materials and Methods
-4.1 Datasets
-We utilized two datasets: the natural scene dataset (NSD) (Allen et al., 2022) and Deeprecon dataset (Shen et al., 2019b).
+Beyond the limitations discussed in this study, there are additional challenges in the field.
+Most reconstruction methods analyze previously acquired brain data offline.
+The brain data used for reconstructing images are often averaged over multiple presentations of the test image, with only a few studies demonstrating single-trial reconstruction results [28, 32].
+Further, it has been argued that subject cooperation is essential for reliably training and testing decoding models [71].
+Given these realities, public expectations often exceed current capabilities, and meeting these high demands in the short term is challenging.
+By clearly articulating these constraints, we can help manage expectations, prevent disappointment, and guide governments and companies away from misguided decisions.
+It is crucial to resist making overly optimistic claims about the ability to reconstruct arbitrary images.
+In light of these challenges, while the field of visual image reconstruction from brain activity holds great promise, it is our responsibility as researchers to ensure that its current capabilities and limitations are accurately communicated to all stakeholders.
+Materials and methods
+Datasets
+We utilized two datasets: the Natural Scene Dataset (NSD) [12] and the Deeprecon dataset [29].
 Both datasets comprise visual stimuli and corresponding fMRI activity collected when subjects perceived the stimuli.
-In the NSD dataset, eight subjects were presented with MSCOCO images (Lin et al., 2014), yielding 30,000 brain activity samples per subject, which is three times the amount provided by the Deeprecon dataset.
-The Deeprecon dataset includes fMRI activity data from subjects presented with both ImageNet images (Deng et al., 2009) and artificial images.
+In the NSD dataset, eight subjects were presented with MSCOCO images [94], yielding $30,000$ brain activity samples per subject, which is three times the amount provided by the Deeprecon dataset.
+The Deeprecon dataset includes fMRI activity data from subjects presented with both ImageNet images [9] and artificial images.
 It contains roughly $8,000$ brain samples per subject.
-Since this dataset is designed for evaluating reconstruction performance, the test data is carefully selected.
+Since this dataset is designed to evaluate reconstruction performance, the test stimuli were carefully selected.
 The test natural images were selected from ImageNet, which were in categories different from those used in the training.
 The artificial images were only used as test data to check the generalizability performance of the proposed reconstruction methods.
-In both datasets, we adopted the train/test split used in previous studies and utilized data from the first subject (S1 in NSD, Subject 1 in Deeprecon).
-Text-guided reconstruction methods require text annotations of images..
+In both datasets, we adopted the train-test split used in previous studies and utilized data from the first subject (S1 in the NSD and Subject 1 in the Deeprecon).
+Text-guided reconstruction methods require text annotations of images.
 For the NSD, text annotations accompanying the MSCOCO database were used.
-For the Deeprecon dataset, we collected captions for each experimental stimulus via crowd workers on Amazon Mechanical Turk, yielding five captions per image.
-The captions of training stimuli are publicly available at the GitHub repository.
-4.2 Reconstruction methods
-We utilized three image reconstruction methods: StableDiffusionReconstruction (Takagi and Nishimoto, 2023b), Brain-diffuser (Ozcelik and VanRullen, 2023) and iCNN (Shen et al., 2019b).
-Each method employs two common steps: first, translating brain activity patterns into latent features of the stimuli, and second, generating images from these latent features using an image generator (see also Figure 1).
-In the StableDiffusionReconstruction method, the latent features are the VAE (Kingma and Welling, 2014) features calculated from stimulus images and CLIP text features (Radford et al., 2021) from the image annotations.
-The generator is StableDiffusion (Rombach et al., 2022).
+For the Deeprecon dataset, we collected text annotations for each experimental stimulus via crowd workers on Amazon Mechanical Turk, yielding five annotations per image.
+The text annotations of training stimuli are publicly available in the GitHub repository (https://github.com/KamitaniLab/GOD_stimuli_annotations).
+Reconstruction methods
+We utilized three image reconstruction methods: StableDiffusionReconstruction [44], Brain-Diffuser [45], and iCNN [29].
+Each method employs two common steps: first, translating brain activity patterns into latent features of the stimuli, and second, generating images from these latent features using an image generator (see also Fig 1).
+In the StableDiffusionReconstruction method, the latent features are the VAE [51] features calculated from stimulus images and the CLIP text features [22] from the image annotations.
+The generator is StableDiffusion [50].
 They first generate low-resolution images from the translated VAE features, and those images are further fed into the StableDiffusion model with translated text features to generate images.
 The generated images are regarded as reconstructed images from brain activity.
-In the Brain-diffuser method, the latent features are the VDVAE (Child, 2021) and CLIP vision features from stimulus images and CLIP text features from the image annotations.
-The generator is Versatile diffusion (Xu et al., 2022).
+In the Brain-Diffuser method, the latent features are the VDVAE [52], CLIP vision features from stimulus images and CLIP text features from the text annotations of the image.
+The generator is Versatile diffusion [49].
 Similar to StableDiffusionReconstruction, low-resolution images are first generated from the translated VDVAE features, and these images are further used for the input of the versatile diffusion model with the translated vision and text features.
 The generated images are regarded as reconstructed images from brain activity.
-In the iCNN method, the latent features are the intermediate output of VGG19 (Simonyan and Zisserman, 2015) layer from stimulus images.
-As a generator, they used the pretrained image generator (Dosovitskiy and Brox, 2016) and they solved the optimization problem to minimize the discrepancy between the VGG19 features calculated from the generated images and the transformed VGG19 features.
+In the iCNN method, the latent features are the intermediate output of the VGG19 layer [56].
+As a generator, they used the pre-trained image generator [69], and they solved the optimization problem to minimize the discrepancy between the VGG19 features calculated from the generated images and the translated VGG19 features.
 Well-optimized images are regarded as reconstructed images.
-4.3 UMAP visualization
-To investigate dataset diversity, we employed Uniform Manifold Approximation and Projection (UMAP), a non-linear dimensionality reduction technique (McInnes et al., 2018) to learn a projection from a latent features space to a lower dimension (UMAP embedding space).
-We used both the training and test CLIP text features for learning the UMAP projection.
+UMAP visualization
+To investigate dataset diversity, we employed uniform manifold approximation and projection (UMAP), a non-linear dimensionality reduction technique [54], to learn a projection from a latent features space to a lower dimension (UMAP embedding space).
+We used both the training and test CLIP text features to learn the UMAP projection.
 These features were combined and standardized beforehand.
-The UMAP hyperparameters follow the official UMAP guide for clustering usage with cosine distance as a distance metric.
-The learned UMAP was also used to project the features predicted from brain activity (Figure 4).
-After standardizing the predicted features using the same mean and standard deviation parameters used in UMAP projection learning, we projected the decoded features into the UMAP embedding space.
-4.4 Simulation with clustered data
-We conducted a simulation analysis to examine the generalization performance beyond the training data (Figure 6).
-This analysis involves a teacher-student learning task using Gaussian mixture distributions.
-We generated both input and target pairs from teachers, then examined the prediction performance of regression models of students.
-To imitate the feature translation situation from brain activity, we first generated feature data and assumed a scenario where observation noise affects the brain data.
-The training sample of latent feature data $\mathbf{y}\in\mathbb{R}^{D}$ was generated from a $D$-dimensional space using Gaussian mixture distributions, formulated as:
-$p_{\mathrm{tr}}(\mathbf{y})=\frac{1}{C}\sum_{c=1}^{C}\mathcal{N}(\boldsymbol{\mu}_{c}^{\mathrm{tr}},\sigma_{r}^{2}I),\quad\boldsymbol{\mu}_{c}^{\mathrm{tr}}\sim\mathcal{N}(\mathbf{0},\sigma_{s}^{2}),$ (8)
+The UMAP hyperparameters followed the official UMAP guide for clustering usage (https://umap-learn.readthedocs.io/en/latest/clustering.html#umap-enhanced-clustering) with cosine distance as a distance metric (the visualization result with the default UMAP parameters is shown in S1 Fig.).
+The learned UMAP was also used to project the features predicted from brain activity (Fig 6B).
+After standardizing the predicted features using the same mean and standard deviation parameters used in UMAP projection learning, we projected the predicted features onto the UMAP embedding space.
+Simulation with clustered data
+We conducted a simulation analysis to illustrate output dimension collapse (Fig 8) and to examine the generalization performance beyond the training data (Fig 9).
+These analyses involve a teacher-student learning task.
+The teacher model generated the pairs of target features and brain activity as training samples, and the student model learned a mapping from the training samples.
+To imitate the feature translation situation from brain activity, observation noise was added to the brain data.
+To imitate the feature translation situation from brain activity, observation noise was added to the brain data.
+The training sample of latent feature data $\mathbf{y}\in\mathbb{R}^{D}$ was generated from a Gaussian mixture distribution, formulated as:
+$p_{\mathrm{tr}}(\mathbf{y})=\frac{1}{C}\sum_{c=1}^{C}\mathcal{N}(\boldsymbol{\mu}_{c}^{\mathrm{tr}},\sigma_{\mathrm{intra}}^{2}I),\quad\boldsymbol{\mu}_{c}^{\mathrm{tr}}\sim\mathcal{N}(\mathbf{0},\sigma_{\mathrm{inter}}^{2}I),$ (5)
 where $C$ is the number of clusters in the training set.
-$\sigma_{r}^{2}$ is the scalar value representing the variance of the Gaussian distribution corresponding to each cluster.
-$\sigma_{s}^{2}$ is the scalar value representing the variance of the distribution of cluster centers ($\boldsymbol{\mu}_{c}^{\mathrm{tr}}$).
-Brain activity data, $\mathbf{x}\in\mathbb{R}^{D}$, were generated using teacher weights $\bar{A}\in\mathbb{R}^{D\times D}$ and incorporating observation noise $\boldsymbol{\xi}$ with $\mathbf{x}=\bar{A}^{\top}\mathbf{y}+\boldsymbol{\xi}$, where $\boldsymbol{\xi}\sim\mathcal{N}(\mathbf{0},\sigma_{n}^{2}I)$.
-$\sigma^{2}_{n}$ is the scalar value representing the variance of observation noise.
-Using $N$ training samples, the ridge regression model was trained.
-With the hyperparameter $\lambda$, the weight of ridge regression model $W$ can be calculated analytically:
-$W=(X_{\mathrm{tr}}^{\top}X_{\mathrm{tr}}+\lambda I)^{-1}X_{\mathrm{tr}}^{\top}Y_{\mathrm{tr}}.$ (9)
-For testing the model performance, we prepared two types of test samples, in-distribution test samples and out-of-distribution test samples.
-In-distribution test samples are generated using the same parameter as $p_{\mathrm{tr}}(\mathbf{y})$.
-Out-of-distribution (OOD) test samples were generated in two steps.
-First, $C_{\mathrm{ood}}$ different cluster centers $\boldsymbol{\mu}_{c}^{\mathrm{ood}}$ were obtained by sampling from a Gaussian distribution: $\boldsymbol{\mu}_{c}^{\mathrm{ood}}\sim\mathcal{N}(\mathbf{0},\sigma_{s}^{2})$.
-Then OOD test samples were generated from these new cluster centers:
-$p_{\mathrm{ood}}(\mathbf{y})=\frac{1}{C_{\mathrm{ood}}}\sum_{c=1}^{C_{\mathrm{ood}}}\mathcal{N}(\boldsymbol{\mu}_{c}^{\mathrm{ood}},\sigma_{r}^{2}I).$ (10)
-To evaluate the model’s zero-shot performance, we evaluate $C+1$ way cluster identification analyis on the $C$ training clusters and the one cluster to which each OOD test sample belongs.
-Similar to Figure 6c, we caluculated the proportion of predicted features from test brain data that were correctly classified as belonging to the original cluster centers compared with the other candidate clusters.
-The similarity between predicted features and each cluster center was calculated using the correlation coefficient, with the highest similarity determining the cluster assignment, where the chance level is $1/(C+1)$.
-We evaluated the cluster identification performance for each of the OOD clusters and reported the median performance.
-Note that, for in-distribution test samples, we also conducted a C + 1 way identification analysis on the C training clusters and one selected OOD cluster for fair comparison.
-In the simulation analysis, we mainly focused on the cluster identification performance for the dimension $D$, for the number of clusters in the training set $C$ or the distance among cluster centers for large sample size $N=500000$.
-Other parameters fixed in this simulation analysis as follows: $\sigma_{r}^{2}=10/D$, $\sigma_{s}^{2}=100/D$, $\bar{A}_{ij}\sim\mathcal{N}(0,D^{-1/2})$, $\sigma_{n}^{2}=0.1$, $\lambda=1.0$, and $C_{\mathrm{ood}}=32$.
-We parameterized the scale of the variances ($\sigma_{r}^{2}$ and $\sigma_{s}^{2}$) and the teacher weights $\bar{A}$ using the dimension $D$ so that the order of magnitude of the output invariant from the dimension.
-4.5 Recovery check of a single layer by iCNN
-We performed a recovery check analysis using a single layer from the iCNN method at Figure 10.
-Briefly, the iCNN method generates an image by optimizing pixel values to make the image’s latent features similar to the target latent features (Shen et al., 2019b).
-In the pixel optimization condition (the left columns of each recovery image in Figure 10), we directly optimized the pixel values of images to minimize the mean squared loss between the latent feature of the image as well as the total-variance (TV) loss of pixel values (Mahendran and Vedaldi, 2015).
-Additionally, the iCNN method can incorporate image generator networks (middle and right columns of each recovery image in Figure 10) to add constraints on image statistics.
+$\sigma_{\mathrm{intra}}^{2}$ is the scalar value representing the variance of the Gaussian distribution corresponding to each cluster.
+$\sigma_{\mathrm{inter}}^{2}$ is the scalar value representing the variance of the distribution of cluster centers ($\boldsymbol{\mu}_{c}^{\mathrm{tr}}$).
+$I$ is the $D\times D$ identity matrix.
+Brain activity data, $\mathbf{x}\in\mathbb{R}^{D}$, were created using teacher weights $\bar{A}\in\mathbb{R}^{D\times D}$ and incorporating observation noise $\boldsymbol{\xi}$ with $\mathbf{x}=\bar{A}^{\top}\mathbf{y}+\boldsymbol{\xi}$, where $\boldsymbol{\xi}\sim\mathcal{N}(\mathbf{0},\sigma_{\mathrm{noise}}^{2}I)$.
+$\sigma^{2}_{\mathrm{noise}}$ is the scalar value representing the variance of observation noise.
+Using $N$ training samples, $X_{\mathrm{tr}}=[\mathbf{x}^{(1)},\ldots,\mathbf{x}^{(N)}]^{\top}$ and $Y_{\mathrm{tr}}=[\mathbf{y}^{(1)},\ldots,\mathbf{y}^{(N)}]^{\top}$, we trained the student model.
+by the ridge regression algorithm.
+The trained weight of ridge regression model $W$ can be calculated analytically:
+$W=(X_{\mathrm{tr}}^{\top}X_{\mathrm{tr}}+\lambda I)^{-1}X_{\mathrm{tr}}^{\top}Y_{\mathrm{tr}},$ (6)
+where $\lambda$ is the regularization parameter.
+After training the student weight $W$, we illustrated the phenomenon of output dimension collapse as in Fig 8 by predicting randomly generated data.
+First, we generated random test latent features from the Gaussian distribution with a mean of $\mathbf{0}$ and a variance equal to the training set variance scaled by a single scalar value ($9.0$).
+The corresponding test brain samples were obtained by translating the latent features by the teacher’s weight $\bar{A}$ and adding observation noise $\boldsymbol{\xi}$.
+The predicted latent features were derived by projecting the test brain samples by the learned student weight $W$.
+We set $D=512$, $\sigma_{\mathrm{intra}}^{2}=10/512$, $\sigma_{\mathrm{inter}}^{2}=100/512$, and $\sigma_{\mathrm{noise}}^{2}=10$.
+We evaluated the zero-shot performance of the student model as in Fig 9.
+We prepared two types of test samples: in-distribution test samples and out-of-distribution test samples.
+In-distribution test samples are generated from one of the (Gaussian) clusters used in training data:
+$p_{\mathrm{{te}}}(\mathbf{y})=\mathcal{N}(\boldsymbol{\mu}^{\mathrm{tr}}_{c},\sigma_{\mathrm{intra}}^{2}I),$ (7)
+where we randomly chose $\boldsymbol{\mu}_{c}^{\mathrm{tr}}$ over $C$ cluster centers.
+Out-of-distribution (OOD) test samples were generated from a novel cluster center $\boldsymbol{\mu}^{\mathrm{ood}}$ that is not included in the training set.
+The novel cluster center was obtained by sampling from a Gaussian distribution $\boldsymbol{\mu}^{\mathrm{ood}}\sim\mathcal{N}(\mathbf{0},\sigma_{\mathrm{inter}}^{2}I)$.
+The OOD test samples were then generated from the novel cluster center:
+$p_{\mathrm{{ood}}}(\mathbf{y})=\mathcal{N}(\boldsymbol{\mu}^{\mathrm{ood}},\sigma_{\mathrm{intra}}^{2}I).$ (8)
+To evaluate the model’s zero-shot performance, we conducted a cluster identification analysis.
+For each test sample, we calculated the similarity between its predicted features and the center of its original cluster, as well as the similarity between its predicted features and the centers of the other candidate clusters.
+Each test sample was assigned to the cluster whose center had the highest similarity with its predicted features, and the proportion of samples correctly assigned to their true cluster centers was calculated across all $n$ test samples.
+We used the correlation coefficient as the similarity measure.
+To reduce the variability associated with the cluster selection, we repeated this process $t$ times by randomly selecting both in-distribution and out-of-distribution cluster centers and reported the median value of the results.
+For in-distribution test samples, we chose the cluster center randomly without replacement.
+We mainly focused on the typical cluster identification performance for dimension $D$, the number of clusters in the training set $C$, and the ratio of variance about the cluster structure $\sigma_{\mathrm{intra}}^{2}/\sigma_{\mathrm{inter}}^{2}$.
+We used a sufficiently large number of training samples $N=500,000$.
+This size was constant when the above parameters were varied.
+Other parameters were also fixed in this simulation as follows: $\sigma_{\mathrm{intra}}^{2}+\sigma_{\mathrm{inter}}^{2}=110/D$, ${\bar{A}_{ij}}\sim\mathcal{N}(0,D^{-1/2})$, $\sigma_{\mathrm{noise}}^{2}=0.25$, $\lambda=1.0$, $n=100$ and $t=32$.
+We parameterized the scale of the variances ($\sigma_{\mathrm{intra}}^{2}$ and $\sigma_{\mathrm{inter}}^{2}$) and the teacher weights $\bar{A}_{ij}$ using the dimension $D$ so that the order of scale of the output samples were invariant from the dimension.
+When the number of clusters in the training set $C$ was less than $t=32$, we set $t=C$ instead.
+Expected identification accuracy in imprecise reconstructions
+A pairwise identification accuracy is a metric defined on three types of samples: the test sample, the predicted sample, and the candidate sample selected from a test set as $\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-}\in\mathcal{Y}$, respectively.
+We define a function $S\colon\mathcal{Y}\times\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$ that takes the triplet above as the input and output whether the predicted sample was much closer to the test sample than the candidate sample as
+$S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})=\begin{cases}1&\left(\operatorname{sim}(\hat{\mathbf{y}},\mathbf{y})\geq\operatorname{sim}(\hat{\mathbf{y}},\mathbf{y}_{-})\right)\\ 0&\left(\text{otherwise}\right)\end{cases}$ (9)
+where $\operatorname{sim}(\cdot,\cdot)$ is an arbitrary function that evaluates a similarity between two samples.
+The pairwise identification accuracy $\mathrm{Acc}$ over $n$ test samples is defined as
+$\mathrm{Acc}=\frac{1}{n(n-1)}\sum_{i}^{n}\sum_{j\neq i}^{n}S(\hat{\mathbf{y}}^{(i)},\mathbf{y}^{(i)},\mathbf{y}_{-}^{(j)}).$ (10)
+Now, we consider a scenario where the translator only decodes semantic information (e.g., category) and cannot decode information about its precise visual appearance.
+Suppose the test set contains a categorical structure like NSD stimuli, we model such a scenario as
+$\mathbb{E}_{p(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-}\mid(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in Z)}[S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})]$ $=0.5,$ (11)
+$\mathbb{E}_{p(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-}\mid(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in\bar{Z})}[S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})]$ $=q\text{ where }q\in[0.5,1].$ (12)
+$Z$ is a set of triplets in which the test sample and the candidate sample belong to the same category.
+$\bar{Z}$ is a complementary set of $Z$.
+$\mathbb{E}_{p(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-}\mid\cdot)}[S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})]$ represents the pairwise identification accuracy in the conditional expectation form.
+If the candidate sample belongs to the same category as the test sample, pairwise identification is challenging because of the poor prediction of the translator.
+On the other hand, if the candidate samples belong to a different category than the test sample, the test sample is easily identified only from the semantic information.
+Here, we assume that the test set contains $k$ categories in total and that all samples are equally distributed across each category for simplicity.
+If we have a sufficiently large number of test samples, the above identification accuracy can be approximated as
+$\mathrm{Acc}$ $\underset{\phantom{n\to\infty}}{=}\frac{1}{n(n-1)}\left(\sum_{(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in Z}S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})\right)+\frac{1}{n(n-1)}\left(\sum_{(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in\bar{Z}}S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})\right)$ (13)
+$\underset{\phantom{n\to\infty}}{=}\frac{|Z|}{n(n-1)}\left(\frac{1}{|Z|}\sum_{(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in Z}S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})\right)+\frac{|\bar{Z}|}{n(n-1)}\left(\frac{1}{|\bar{Z}|}\sum_{(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in\bar{Z}}S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})\right)$ (14)
+=$\underset{n\to\infty}{=}\frac{1}{k}\cdot\mathbb{E}_{p(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-}\mid(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in Z)}[S]+\left(1-\frac{1}{k}\right)\cdot\mathbb{E}_{p(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-}\mid(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in\bar{Z})}[S]$ (15)
+$\underset{\phantom{n\to\infty}}{=}\frac{1}{k}\cdot 0.5+\left(1-\frac{1}{k}\right)\cdot q,$ (16)
+where $|Z|=n\left(n/k-1\right)$, and $|\bar{Z}|=n\left(n-n/k\right)$. We used the assumption of a large sample size at the third equality.
+Recovery check of a single layer by iCNN
+We performed a recovery check analysis using a single layer from the iCNN method in Fig 11.
+The iCNN method generates an image by optimizing pixel values to make the image’s latent features similar to the target latent features [29].
+In the pixel optimization condition (the left columns of each recovery image in Fig 11), we directly optimized the pixel values of images to minimize the mean squared loss between the latent features of the image as well as the total-variance (TV) loss of pixel values [65].
+Additionally, the iCNN method can incorporate image generator networks (middle and right columns of each recovery image in Fig 11) to add constraints on image statistics.
 Instead of optimizing the pixel values, we optimized the parameters related to the generator networks to minimize the mean squared loss between the latent features obtained through the generator networks and the target latent features.
-As a weak image prior, we used Deep image prior (DIP) (Ulyanov et al., 2017).
+As a weak image prior, we used Deep Image Prior (DIP) [68].
 DIP utilizes a hierarchical U-Net architecture as an inherent prior for image tasks, capturing the statistical regularities of images without relying on a specific dataset.
@@
 In our analysis, DIP started with a U-Net initialized with random noise.
-Subsequently, the latent features and parameters of DIP were optimized to minimize the difference between the network’s output and target DNN features.
-For the pretrained image prior, we used the same generator model as in Shen et al.
-(2019b) (Dosovitskiy and Brox, 2016), optimizing the latent features of the pretrained networks.
-4.6 Expected identification accuracy in imprecise reconstructions
-A pairwise identification accuracy is a metric defined on three types of samples: the test sample, the predicted sample, and the candidate sample selected from a test set as $\mathbf{y},\mathbf{\hat{y}},\mathbf{y}_{-}\in\mathcal{Y}$, respectively.
-We define a function $S\colon\mathcal{Y}\times\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$ that takes the aforementioned triplet as the input and output whether the predicted sample was much close to the test sample than the candidate sample as
-$S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})=\begin{cases}1&\left(\operatorname{sim}(\hat{\mathbf{y}},\mathbf{y})\geq\operatorname{sim}(\hat{\mathbf{y}},\mathbf{y}_{-})\right)\\ 0&\left(\text{otherwise}\right)\end{cases}$ (11)
-where $\operatorname{sim}(\cdot,\cdot)$ is an arbitrary function that evaluates a similarity between two samples.
-The pairwise identification accuracy $\mathrm{Acc}$ over $n$ test samples is defined as
-$\mathrm{Acc}=\frac{1}{n(n-1)}\sum_{i}^{n}\sum_{j\neq i}^{n}S(\hat{\mathbf{y}}^{(i)},\mathbf{y}^{(i)},\mathbf{y}_{-}^{(i)}).$ (12)
-Now we consider a scenario where the translator only decodes semantic information (e.g., category) and cannot decode information about its precise visual appearance.
-Suppose the test set contains a categorical structure like NSD stimuli, we model such a scenario as
-$\mathbb{E}_{p(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-}\mid(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in Z)}[S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})]$ $=0.5,$ (13)
-$\mathbb{E}_{p(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-}\mid(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in\bar{Z})}[S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})]$ $=q\text{ where }q\in[0.5,1].$ (14)
-$Z$ is a set of triplet of which the test sample and the candidate sample belong to the same category.
-$\bar{Z}$ is a complementary set of $Z$.
-$\mathbb{E}_{p(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-}\mid\cdot)}[S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})]$ represents the pairwise identification accuracy in the conditional expectation form.
-If the candidate sample belongs to the same category as the test sample, pairwise identification is challenging because of the poor prediction of the translator.
-On the other hand, if the candidate samples belongs to a different category than the test sample, the test sample is easily identified only from the semantic information.
-Here, we assume that the test set contains $k$ categories in total and all samples are equally distributed across each category for simplicity.
-If we have sufficiently large number of test samples, the above identification accuracy can be approximated as
-$\mathrm{Acc}$ $\underset{\phantom{n\to\infty}}{=}\frac{1}{n(n-1)}\left(\sum_{(\mathbf{y},\mathbf{\hat{y}},\mathbf{y}_{-})\in Z}S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})\right)+\frac{1}{n(n-1)}\left(\sum_{(\mathbf{y},\mathbf{\hat{y}},\mathbf{y}_{-})\in\bar{Z}}S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})\right)$ (15)
-$\underset{\phantom{n\to\infty}}{=}\frac{|Z|}{n(n-1)}\left(\frac{1}{|Z|}\sum_{(\mathbf{y},\mathbf{\hat{y}},\mathbf{y}_{-})\in Z}S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})\right)+\frac{|\bar{Z}|}{n(n-1)}\left(\frac{1}{|\bar{Z}|}\sum_{(\mathbf{y},\mathbf{\hat{y}},\mathbf{y}_{-})\in\bar{Z}}S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})\right)$ (16)
-=$\underset{n\to\infty}{=}\frac{1}{k}\cdot\mathbb{E}_{p(\mathbf{\hat{y}},\mathbf{y},\mathbf{y}_{-}\mid(\mathbf{y},\mathbf{\hat{y}},\mathbf{y}_{-})\in Z)}[S]+\left(1-\frac{1}{k}\right)\cdot\mathbb{E}_{p(\mathbf{\hat{y}},\mathbf{y},\mathbf{y}_{-}\mid(\mathbf{y},\mathbf{\hat{y}},\mathbf{y}_{-})\in\bar{Z})}[S]$ (17)
-$\underset{\phantom{n\to\infty}}{=}\frac{1}{k}\cdot 0.5+\left(1-\frac{1}{k}\right)\cdot q,$ (18)
-where $|Z|=n\left(n/k-1\right)$, and $|\bar{Z}|=n\left(n-n/k\right)$. We used the assumption on the large sample size at the third equality.
-5 Acknowledgements
-We thank our laboratory team, especially Eizaburo Doi, Hideki Izumi, Matthias Mildenberger and for their invaluable comments and suggestions on the manuscript.
-This work was supported by the Japan Society for the Promotion of Science (JSPS: KAKENHI grants JP20H05954, JP20H05705 to Y.K, JP21K17821 to Y.N, and 22KJ1801 to K.S.), Japan Science and Technology Agency (JST: CREST grants JPMJCR18A5, and JPMJCR22P3 to Y.K.), and New Energy and Industrial Technology Development Oprganization (NEDO: commissioned project, JPNP20006 to Y.K.).
-Appendix A Supplementary figures
-Figure S1: UMAP visualization of CLIP text features in Deeprecon. Blue points represent training samples, while red points represent test samples.
-Figure S2: Supplementary information of hold out analysis.
-(a) The clustering results, applying the $k$-Means clustering to the UMAP embedding of CLIP text features.
-Based on these clustering results, the assignment of training and test samples under the hold-out split condition was determined.
-(b) Confusion probability matrix of cluster identification results.
-The top row shows the cluster identification results under the hold-out split conditions.
-The left side is for CLIP text features and the right side is for CLIP vision features.
-The bottom row shows the cluster identification results under the naive split conditions.
-In each matrix, the $(i,j)$ cell indicates the ratio of samples assigned by cluster $i$ are predicted to cluster $j$.
-The empty cell means the ratio is 00.
-Figure S3: Additional results about simulation with clustered data.
-(a) Cluster identification accuracy as a function of the cluster variance ratio.
-The $x$ axis indicates the cluster variance ratio $\sigma_{r}^{2}/\sigma_{s}^{2}$ and the $y$ axis represents the cluster identification performance.
-Green lines show the results of in-distribution test samples and orange lines show those for OOD samples.
-The number of clusters used in training is increasing as the color lightens.
-Similar to Figure 9b in the main figure, the OOD identification performance become closer to the in distribution samples as the cluster variance ratio is increasing.
-(b).
-Heatmap plot of cluster identification accuracy varying the number of clusters in training and dimension.
-The left matrix shows the cluster identification results for the in-distribution test samples.
-The right matrix shows the cluster identification results for the OOD test samples.
-In each matrix, the $(i.j)$ cell indicates the cluster identification accuracy when the dimension is $i$ and the number of clusters used in the training is $j$.
-Figure S4: Frame simirality in Nishimoto et al.
-(2011).
-We first detected scenes from the training and test movie stimuli.
-Then, for the first and last frames of each test scene, we searched for the frame from the training movies with the closest Euclidean distance.
+Subsequently, the latent features and weight parameters of DIP were optimized to minimize the difference between the network’s output and target DNN features.
+For the pre-trained image prior [69], we used the same generator model as in Shen et al.
+[29] optimizing the latent features of the pre-trained networks.
+Acknowledgments
+We thank our laboratory team, especially Eizaburo Doi, Haibao Wang, Kenya Otsuka, Hideki Izumi, and Matthias Mildenberger for their invaluable comments and suggestions on the manuscript.
+This work was supported by the Japan Society for the Promotion of Science (JSPS: KAKENHI grants JP20H05954, JP20H05705 to Y.K, JP21K17821 to Y.N, and 22KJ1801 to K.S.), Japan Science and Technology Agency (JST: CREST grants JPMJCR18A5, and JPMJCR22P3 to Y.K.), and New Energy and Industrial Technology Development Organization (NEDO: commissioned project, JPNP20006 to Y.K.).
+Supporting information
+S1 Fig.
+UMAP visualization of CLIP text features in the Natural Scene Dataset (NSD) using default parameters.
+Similar to Fig 4A in the main text, this scatter plot visualizes the distribution of semantic concepts within the NSD.
+The gray points represent training samples, and the orange points represent test samples.
+Unlike Fig 4A, which used parameters optimized for clustering visualization, this figure employs default UMAP settings.
+Despite the difference in parameters, the visualization still reveals distinct clusters and substantial overlap between training and test samples.
+S2 Fig.
+UMAP visualization of CLIP text features in the Deeprecon dataset.
+This scatter plot visualizes the distribution of semantic concepts within the Deeprecon data, with the gray points representing training samples and the orange points indicating test samples.
+Unlike the NSD dataset (Fig 4A in the main text), this visualization demonstrates a clearer separation between training and test samples.
+This distinction highlights the Deeprecon dataset’s intentional design to differentiate object categories between training and test sets.
+S3 Fig.
+Confusion matrices from the hold-out analyses.
+(A) Clustering results of CLIP text features.
+This panel shows the outcome of applying $k$-means clustering to the UMAP embedding of CLIP text features from the NSD dataset (Fig 4A in the main text).
+These clustering results were used to determine the assignment of training and test samples under the hold-out split condition (Fig 6A in the main text).
+(B) Confusion probability matrices for cluster identification.
+These matrices illustrate the performance of cluster identification under different conditions.
+The top row shows results under the hold-out split conditions, while the bottom row presents results under the naive split conditions.
+For each row, the left matrix represents CLIP text features, and the right matrix represents CLIP vision features.
+In each matrix, the cell at position $(i,j)$ indicates the proportion of samples originally from cluster $i$ that were predicted to belong to cluster $j$.
+As visualized in the confusion matrices, the hold-out split condition, where entire semantic clusters are excluded from training, reveals the challenges in zero-shot prediction of novel concepts.
+In contrast, the naive split condition, where samples from all clusters are included in the training, demonstrates the model’s performance when semantic concepts overlap between training and test sets.
+S4 Fig.
+Extended simulation results for clustered data analysis.
+(A) Relationship between cluster identification accuracy and cluster variance ratio.
+Each subplot represents a different number of clusters used in model training.
+The $x$-axis represents the cluster variance ratio $\sigma_{\mathrm{intra}}^{2}/\sigma_{\mathrm{inter}}^{2}$, and the $y$-axis represents the cluster identification performance.
+The green lines represent results for in-distribution test samples, and the orange lines show outcomes for out-of-distribution (OOD) samples.
+As the cluster variance ratio increases, clusters tend to overlap more, leading to difficulty with identification in the in-distribution condition.
+However, larger cluster variance also allows for more extensive coverage of the feature space, potentially improving OOD prediction.
+Consequently, OOD curves typically show a peak at intermediate variance ratios, reflecting a balance between cluster distinctiveness and feature space coverage.
+(B) Heatmap of cluster identification accuracy.
+These matrices visualize how accuracy changes with varying numbers of training clusters and feature space dimensions.
+The cluster variance ratio ($\sigma_{\mathrm{intra}}^{2}/\sigma_{\mathrm{inter}}^{2}$) is fixed at $0.1$ for all simulations.
+The left matrix shows results for in-distribution samples, while the right matrix displays outcomes for out-of-distribution (OOD) samples.
+In each matrix, cell $(i,j)$ indicates the cluster identification accuracy when the feature dimension is $i$ and the number of clusters used in training is $j$.
+This view illustrates the interplay between feature dimensionality and the number of training clusters in determining the model’s generalization capabilities.
+Notably, in the OOD condition, as the feature dimensionality increases, a greater number of training clusters becomes necessary to achieve a robust zero-shot prediction.
+S5 Fig.
+Frame similarity analysis in the dataset of Nishimoto et al.
+[89].
+Movie scenes were detected from the movie stimuli.
+Then, for the first and last frames of each test scene, we identified the frames from the training movies with the closest Euclidean distance.
 Our analysis revealed that $37$ out of the $48$ scenes in the test set contained frames that were nearly identical to those in the training set.

### v5

@@
 Abstract
-Advances in brain decoding, particularly visual image reconstruction, have sparked discussions about the societal implications and ethical considerations of neurotechnology.
-As these methods aim to recover visual experiences from brain activity and achieve prediction beyond training samples (zero-shot prediction), it is crucial to assess their capabilities and limitations to inform public expectations and regulations.
-Our case study of recent text-guided reconstruction methods, which leverage a large-scale dataset (Natural Scene Dataset, NSD) and text-to-image diffusion models, reveals limitations in their generalizability.
-We found poor performance when applying these methods to a different dataset designed to prevent category overlaps between training and test sets.
-UMAP visualization of the text features with NSD images showed a limited diversity of semantic and visual clusters, with overlap between training and test sets.
-Formal analysis and simulations demonstrated that clustered training samples can lead to “output dimension collapse,” restricting predictable output feature dimensions.
-Simulations further showed that diversifying the training set improved generalizability.
-However, text features alone are insufficient for mapping to the visual space.
-We argue that recent realistic reconstructions may primarily be a blend of classification into trained categories and generation of inauthentic images through text-to-image diffusion (hallucination).
-Diverse datasets and compositional representations spanning the image space are essential for genuine zero-shot prediction.
-Interdisciplinary discussions grounded in understanding the current capabilities and limitations, as well as ethical considerations, of the technology are crucial for its responsible development.
-Author summary
-Brain decoding has gained attention not only in neuroscience but also in fields like computer science, and even policy-making.
-However, claims about its predictive performance often tend to be overly optimistic.
-Here, we critically examine recent brain decoding methods, particularly visual image reconstruction methods that use generative AIs to generate seemingly realistic images from brain activity.
-Our analyses reveal that the apparent success of these methods largely stems from exploiting similarities between training and test stimuli in the experiment and from the hallucination of recent generative AIs, failing to achieve the genuine zero-shot reconstruction of novel images.
-We introduce the concept of “output dimension collapse,” a fundamental limitation in predicting within a target space, and discuss conditions necessary for true zero-shot prediction in terms of compositional representations and image recovery from latent features.
-We also highlight the limitations of current evaluation metrics and provide a framework for interpreting seemingly realistic reconstructions.
-By examining current limitations and proposing improvement strategies, we aim to encourage rigorous development and evaluation of brain decoding techniques.
-Introduction
-Brain decoding has been widely used in the neuroscience field, revealing specific contents of the mind [1, 2, 3, 4].
-As brain decoding is sometimes referred to as “mind-reading” in popular media [5, 6, 7], it has attracted significant attention beyond the scientific community due to its potential for real-world applications in medicine and industry.
-Such neurotechnology has also started to affect future ethical discussions and legal regulations [8].
+Advances in brain decoding, particularly in visual image reconstruction, have sparked discussions about the societal implications and ethical considerations of neurotechnology.
+As reconstruction methods aim to recover visual experiences from brain activity and achieve prediction beyond training samples (zero-shot prediction), it is crucial to assess their capabilities and limitations to inform public expectations and regulations.
+Our case study of recent text-guided reconstruction methods, which leverage a large-scale dataset (Natural Scene Dataset, NSD) and text-to-image diffusion models, reveals critical limitations in their generalizability, demonstrated by poor reconstructions on a different dataset.
+UMAP visualization of the text features from NSD images shows limited diversity with overlapping semantic and visual clusters between training and test sets.
+We identify that clustered training samples can lead to “output dimension collapse,” restricting predictable output feature dimensions.
+While diverse training data improves generalization over the entire feature space without requiring exponential scaling, text features alone prove insufficient for mapping to the visual space.
+Our findings suggest that the apparent realism in current text-guided reconstructions stems from a combination of classification into trained categories and inauthentic image generation (hallucination) through diffusion models, rather than genuine visual reconstruction.
+We argue that careful selection of datasets and target features, coupled with rigorous evaluation methods, is essential for achieving authentic visual image reconstruction.
+These insights underscore the importance of grounding interdisciplinary discussions in a thorough understanding of the technology’s current capabilities and limitations to ensure responsible development.
+keywords:
+Brain decoding , Visual image reconstruction , Naturalistic approach , NeuroAI
+journal: Neural Networks
+[…]
+[inst1]organization=Graduate School of Informatics, Kyoto University,addressline=Yoshida-honmachi, Sakyo-ku, city=Kyoto, postcode=606-8501, country=Japan
+[…]
+[inst2]organization=Computational Neuroscience Laboratories, Advanced Telecommunications Research Institute International,addressline=Seika-cho, city=Sorakugun, postcode=619-0288, country=Japan
+[…]
+[inst3]organization=Institute for Quantum Life Science, National Institutes for Quantum Science and Technology,city=Chiba, postcode=263-8555, country=Japan
+[…]
+[inst4]organization=Guardian Robot Project, RIKEN,addressline=Seika-cho, city=Sorakugun, postcode=619-0288, country=Japan
+1 Introduction
+Brain decoding has been widely used in the neuroscience field, revealing specific contents of the mind (Haxby et al., 2001; Kamitani and Tong, 2005; Soon et al., 2008; Horikawa et al., 2013).
+As brain decoding is sometimes referred to as “mind-reading” in popular media (Somers, 2021; Whang, 2023; Raasch, 2023), it has attracted significant attention beyond the scientific community due to its potential for real-world applications in medicine and industry.
+Such neurotechnology has also started to affect future ethical discussions and legal regulations (UNESCO, 2023).
 To prevent misleading public expectations and policies, scientists need to carefully assess the current status of brain decoding techniques and clarify the possibilities and limitations.
 One of the major challenges in brain decoding is the limited amount of brain data we can collect.
-The current brain measurement devices are costly, yielding far less brain data than the amounts typically used in image or text processing within the field of computer science and AI [9, 10].
-Although we have gradually increased the amount of brain data per subject [11, 12, 13, 14, 15], it remains impractical to collect brain data covering the full range of cognitive states and perceptual experiences.
-Consequently, classification-based decoding approaches, primarily developed in the early stages of this field, are insufficient to uncover the neural code in general or natural conditions since decodable information is confined to the same stimuli or predefined categories used in a training phase.
+The current brain measurement devices are costly, yielding far less brain data than the amounts typically used in image or text processing within the field of computer science and AI (Deng et al., 2009; Schuhmann et al., 2022).
+Although we have gradually increased the amount of brain data per subject (Van Essen et al., 2012; Allen et al., 2022; Naselaris et al., 2021; Hebart et al., 2023; Xu et al., 2024), it remains impractical to collect brain data covering the full range of cognitive states and perceptual experiences.
+The scarcity of brain data limits the applicability and scalability of classification-based decoding approaches, which are primarily developed in the early stage of this field.
+Such approaches can only decode information confined to the same stimuli or predefined categories used in the training phase, rendering them insufficient for uncovering the neural representation under general or natural conditions.
 To overcome this limitation, several decoding methods have been developed to enable the prediction of novel contents from brain activities that are not encountered during the training phase.
 Kay et al.
-[16] proposed a general visual decoding approach via a statistical encoding model that predicted fMRI voxel values from image features.
+(2008) proposed a general visual decoding approach via a statistical encoding model that predicted fMRI voxel values from image features.
 It successfully identified novel test images from a set of $1,000$ candidates.
 Mitchell et al.
-[17] utilized co-occurrence rates of specific verb sets for nouns and built a computational model to predict fMRI voxel values while thinking about nouns in presented line drawing images.
+(2008) utilized co-occurrence rates of specific verb sets for nouns and built a computational model to predict fMRI voxel values while thinking about nouns in presented line drawing images.
 Their model demonstrated the ability to predict voxel values for novel nouns not seen during the training phase.
-Brouwer and Heeger [18] constructed a color-tuning model and predicted brain activity while the subjects were presented with color stimuli.
+Brouwer and Heeger (2009) constructed a color-tuning model and predicted brain activity while the subjects were presented with color stimuli.
 As their training stimuli covered most of the color space, their methods successfully identified novel colors not included in the training dataset.
-Horikawa and Kamitani [19] utilized deep neural network (DNN) features to decode brain activity measured while subjects perceived natural images.
+Horikawa and Kamitani (2017) utilized deep neural network (DNN) features to decode brain activity measured while subjects perceived natural images.
 They showed successful prediction of novel object categories not encountered during the training phase.
-In the field of machine learning, “zero-shot” prediction refers to the ability of a model to accurately predict or classify novel contents not encountered during the training phase [20, 21].
-This capability has emerged in various applications across different domains, including image classification [22], image generation [23], and natural language processing [24].
+In the field of machine learning, “zero-shot” prediction refers to the ability of a model to accurately predict or classify novel contents not encountered during the training phase (Larochelle et al., 2008; Palatucci et al., 2009).
+This ability has emerged in various applications across different domains, including image classification (Radford et al., 2021), image generation (Ramesh et al., 2021), and natural language processing (Brown et al., 2020).
 The concept of zero-shot prediction can be considered analogous to brain decoding techniques that aim to interpret brain activity patterns associated with previously unseen stimuli or experiences.
 Both approaches seek to generalize knowledge gained from a limited set of training data to novel situations, enabling the interpretation of new information without explicit prior exposure.
-To achieve effective zero-shot prediction, the model often utilizes a compositional representation of the output [25, 26].
-Compositional representation refers to the ability to understand and generate novel combinations of previously learned concepts or features.
-By learning the underlying structure and relationships between different elements, the model can generalize its knowledge to new, unseen instances.
+To achieve effective zero-shot prediction, the model often utilizes a compositional representation of the output (Lake et al., 2017; Higgins et al., 2018).
+Compositional representation enables the understanding and generating of novel features through the combination of previously learned ones.
+By learning the underlying structure and relationships between different features, the model can generalize its knowledge to new, unseen instances.
 Visual image reconstruction is another prominent example of zero-shot prediction in brain decoding.
-This task aims to recover perceived novel images that were not encountered during the training phase, effectively reconstructing visual experiences from brain activity patterns [27, 28].
+This task aims to recover perceived novel images that were not encountered during the training phase, effectively reconstructing visual experiences from brain activity patterns (Stanley et al., 1999; Miyawaki et al., 2008).
 As our perceptual visual experiences cannot be fully covered by limited brain data, reconstruction methods require strong generalizability.
 Miyawaki et al.
-[28] conducted a study demonstrating the reconstruction of perceived arbitrary $10\times 10$ binary-contrast images from brain activity.
+(2008) conducted a study demonstrating the reconstruction of perceived arbitrary $10\times 10$ binary-contrast images from brain activity.
 They built multiple modular decoders to predict the local contrasts of each location and combined their predictions.
 This approach leverages the compositional representation of the visual field, which is organized retinotopically in the early visual cortex.
-Incorporating cortical organization into the model’s architecture can significantly improve its ability to perform zero-shot prediction and reconstruct novel visual experiences from brain activity.
+Incorporating cortical organization into the model’s architecture can improve its ability to perform zero-shot prediction and reconstruct novel visual experiences from brain activity.
 Although the training stimuli were only $400$ random images, it was possible to reconstruct an arbitrary image from a set of possible $2^{100}$ instances, including geometric shapes such as crosses and alphabets.
 Similarly, Shen et al.
-[29] replaced local decoders with DNN feature decoders.
+(2019b) replaced local decoders with DNN feature decoders.
 Although their training stimuli were $1,200$ natural images, they demonstrated reconstructing novel images, including artificial images, which were not part of the training set.
-These successes suggest that the proposed reconstruction models capture rich and comprehensive information about the general aspects of the neural code, beyond merely the information defined by the training data [30].
-Developing reliable reconstruction methods also enables further analysis of subjective visual experiences, such as visual imagery [29], attention [31], and illusion [32].
-Enabling the decoding of novel brain states that were never encountered during the training phase can be a promising approach to neural mind-reading [2].
-Visual image reconstruction pipelines typically comprise three main components: translator, latent features, and generator (Fig 1).
-The translator converts brain activity patterns into a latent feature space, employing either linear regression [29, 33, 34, 35] or nonlinear transformation [36].
-Latent features serve as surrogate representations of perceived visual images, evolving from primitive forms like local contrasts [28] to more sophisticated DNN features, such as intermediate outputs of recognition models [19, 29].
+These successes suggest that the proposed reconstruction models capture rich and comprehensive information about the general aspects of the neural representation, beyond merely the information defined by the training data (Kriegeskorte and Douglas, 2019).
+Developing reliable reconstruction methods also enables further analysis of subjective visual experiences, such as visual imagery (Shen et al., 2019b), attention (Horikawa and Kamitani, 2022), and illusion (Cheng et al., 2023).
+Decoding novel brain states that were never encountered during the training phase can be a promising approach to neural mind-reading (Kamitani and Tong, 2005).
+Visual image reconstruction pipelines typically comprise three main components: translator, latent features, and generator (Fig.
+1).
+The translator converts brain activity patterns into a latent feature space, employing either linear regression (Shen et al., 2019b; Seeliger et al., 2018; Mozafari et al., 2020; Ozcelik et al., 2022) or nonlinear transformation (Qiao et al., 2020).
+Latent features serve as surrogate representations of perceived visual images, evolving from primitive forms like local contrasts (Miyawaki et al., 2008) to more sophisticated DNN features, such as intermediate outputs of recognition models (Horikawa and Kamitani, 2017; Shen et al., 2019b).
 In our current work, we reframe this process as “translation” rather than “feature decoding,” a term we used in previous studies.
-This terminology acknowledges two important points: first, brain activity itself can be considered a latent representation of an image or perception of it, and second, it helps avoid potential ambiguity between image encoding/decoding and brain encoding/decoding processes.
+This terminology acknowledges two important points: first, brain activity itself can be considered a latent representation of an image or the perception formed by it and second, it helps avoid potential ambiguity between image encoding/decoding and brain encoding/decoding processes.
 This new perspective conceptualizes the process as a translation between two latent spaces: from neural representation to machine representation.
-Fig 1: Visual image reconstruction pipeline.
+Fig.
+1: Visual image reconstruction pipeline.
 The first step involves translating brain activity patterns into machine/AI latent representations.
@@
 The generator visualizes translated latent features into images.
-Some studies have used pretraining image generative models for the generator module [34, 36, 35].
-Image optimization can also be regarded as a generator [29].
-End-to-end mapping from brain activity to images using DNNs can also be considered to contain these components implicitly or as a generator-only method [37, 38, 39, 40, 41, 42, 43].
+Some studies have used pretraining image generative models for the generator module (Mozafari et al., 2020; Qiao et al., 2020; Ozcelik et al., 2022).
+Image optimization can also be regarded as a generator (Shen et al., 2019b).
+End-to-end mapping from brain activity to images using DNNs can also be considered to contain these components implicitly or as a generator-only method (Fujiwara et al., 2013; Shen et al., 2019a; Beliy et al., 2019; Ren et al., 2021; Gaziv et al., 2022; Lin et al., 2022; Chen et al., 2023).
 Recent advances in generative AI, particularly in text-to-image generation, have naturally given rise to expectations that these techniques could provide a valuable tool for visual image reconstruction by leveraging semantic representations.
 In addition, there has been a growing trend towards collecting neural datasets using a wide range of diverse visual and semantic content.
-This shift aims to capture a more comprehensive and ecologically valid representation of the human experience [13].
-Researchers have started to collect large-scale fMRI datasets, such as the Natural Scene Dataset (NSD) [12] and the THINGS-fMRI dataset [14], which incorporate a broader range of brain data induced by diverse visual stimuli with text or category annotations.
-Notably, recent studies leveraging the latest generative AI techniques with semantic representation and large-scale datasets have reported realistic reconstructions from brain activity [44, 45, 46, 47, 48].
-They utilized contrastive language-image pretraining (CLIP) features [22] as latent features and text-to-image diffusion models [23, 49] as generators.
+This shift aims to capture a more comprehensive and ecologically valid representation of the human experience (Naselaris et al., 2021).
+Researchers have started to collect large-scale fMRI datasets, such as the Natural Scene Dataset (NSD; Allen et al., 2022) and the THINGS-fMRI dataset (Hebart et al., 2023), which include more than $10,000$ brain samples per subject.
+These datasets incorporate a broader range of brain data induced by diverse visual stimuli with text or category annotations.
+Importantly, recent studies have demonstrated that combining large-scale datasets with generative AI techniques can lead to more realistic reconstructions from brain activity (Takagi and Nishimoto, 2023a; Ozcelik and VanRullen, 2023; Scotti et al., 2023; Bai et al., 2024; Benchetrit et al., 2024; Scotti et al., 2024).
+These approaches commonly utilize linear regression models as translators, contrastive language-image pretraining (CLIP) text features (Radford et al., 2021) as part of the latent features, and text-to-image diffusion models (Ramesh et al., 2021; Xu et al., 2022) as generators.
+MindEye2 (Scotti et al., 2024) has recently shown improved reconstruction performance on the NSD, using a nonlinear translator, latent features of a CLIP’s image model, and a fine-tuned generator.
+This approach also includes a refinement step that enhances the realism of reconstructed images.
 While these recent approaches show promising results, it remains uncertain whether these methods truly achieve zero-shot reconstruction due to several factors.
@@
 It is unclear whether the recently proposed datasets, such as the NSD, are optimally designed to capture the full range of human visual experiences and to support the development of truly generalizable prediction models.
-In the following, we begin with a case study that critically tests the methods proposed by Takagi and Nishimoto [44] and Ozcelik and VanRullen [45], which originally used the NSD.
-Our findings reveal several issues, including the failure to replicate the same level of reconstruction with another dataset specifically designed to avoid overlap between training and test sets [29], and the questionable post-hoc selection procedures of Takagi and Nishimoto that can produce convincing reconstructions even with random brain data.
-We identify the lack of diversity represented by the limited number of semantic clusters in the NSD dataset as a potential factor contributing to these issues.
-We also demonstrate the failure of zero-shot prediction in the latent feature space and the inability to recover a stimulus from its latent features, suggesting a misspecification of the model.
-Based on these findings, we conclude that the apparent realistic reconstructions are essentially a result of classification to clusters shared between training and test sets, combined with hallucinations by the generative model.
-In the formal analysis and simulation section, we aim to uncover the general factors underlying the issues and observations highlighted in the case study.
-We first describe the phenomenon of “output dimension collapse” in the translation from brain activity to the latent feature space.
-We show that the regression model trained on the clustered targets becomes specialized to the training examples, and its outputs collapse into a subspace of the training set.
-Our simulations with clustered data demonstrate that out-of-sample prediction can be achieved with an increasing number of training clusters, suggesting that zero-shot prediction is possible if stimulus diversity and a compositional representation are available.
+In the following, we begin with a case study that critically tests text-guided reconstruction methods.
+We evaluate the approaches of Takagi and Nishimoto (2023a) and Ozcelik and VanRullen (2023), which were originally evaluated using the NSD (additionally, Scotti et al.
+(2024)’s method is examined).
+Our analysis reveals limitations in these methods.
+First, reconstruction quality substantially degrades when tested on a dataset specially designed to avoid object category overlaps between training and test sets (Shen et al., 2019b).
+Second, the post-hoc selection procedure used by Takagi and Nishimoto (2023a) can produce seemingly convincing reconstructions even from random brain data when applied to the NSD.
+Further investigation reveals limited semantic and visual diversity in the NSD stimulus set, with few distinct semantic clusters, potentially explaining these issues.
+We also demonstrate the failure of zero-shot prediction in the latent feature space and the inability to recover a stimulus from its latent features, suggesting fundamental constraints in the latent feature representation.
+These findings indicate that the apparent realism of reconstructions likely results from classification into clusters shared between training and test sets, combined with hallucinations by the generative model.
+In the formal analysis and simulation section, we investigate the general factors underlying the limitations observed in our case study.
+We introduce the phenomenon of “output dimension collapse” that occurs when translating brain activity into latent feature space.
+Our analysis shows that regression models trained on clustered targets become overly specialized to training examples, causing their outputs to collapse into a restricted subspace of the training set.
+Through systematic simulations with clustered data, we demonstrate that successful out-of-sample prediction requires the number of training clusters to scale linearly with feature dimensionality, suggesting that zero-shot prediction becomes feasible given sufficient stimulus diversity and compositional representations.
 We also discuss the caveats associated with evaluating reconstructions using identification metrics alone and explore the preservation of image information at hierarchical layers of DNNs.
-Finally, we provide speculation on how we are fooled by the seemingly realistic reconstructions generated by AI models.
-We conclude with recommendations for critically testing reconstruction models.
-Our findings highlight the importance of rigorous evaluation, the use of diverse datasets, and the need for careful specification and characterization of model components to advance the field of visual image reconstruction and develop truly generalizable prediction capabilities.
-Results
-Case study
-We investigated two types of recent generative AI-based reconstruction methods, StableDiffusionReconstruction [44] and Brain-Diffuser [45], as well as their validation dataset, the Natural Scene Dataset (NSD) [12].
-Both reconstruction methods utilize CLIP features [22] to effectively apply recent text-to-image diffusion models in visual image reconstruction analysis.
+Finally, we provide general accounts on how we could be fooled by seemingly realistic reconstructions generated by AI models.
+Based on these analyses, we conclude with recommendations for developing more reliable reconstruction methods and establishing rigorous evaluation protocols.
+2 Results
+2.1 Case study
+We primarily investigated two recent generative AI-based reconstruction methods, StableDiffusionReconstruction (Takagi and Nishimoto, 2023a) and Brain-Diffuser (Ozcelik and VanRullen, 2023), as well as their validation dataset, the Natural Scene Dataset (NSD; Allen et al., 2022).
+We selected these two methods for three key reasons.
+First, these methods represent reconstruction approaches that have gained significant public attention by leveraging recent advances in generative AI.
+Second, both utilize the NSD, which currently serves as a widely adopted benchmark for predictive modeling in the field.
+Third, these methods employ straightforward linear translators through ridge regression, an approach that has become standard practice.
+Both reconstruction methods utilize CLIP features (Radford et al., 2021) to effectively apply recent text-to-image diffusion models in visual image reconstruction analysis.
 CLIP text features are obtained from the average of five text annotations corresponding to the stimulus image.
 This text annotation information is only used during training to map the brain activity into the CLIP text features.
-In the test phase, they directly translate text features from brain activity during image perception.
-Hence, these two reconstruction methods will be referred to together as text-guided reconstruction methods.
-The StableDiffusionReconstruction method [44] uses the components of the Stable Diffusion model [50], the VAE features [51], and CLIP text features as latent features.
-Similarly, the Brain-Diffuser method [45] utilizes the component of another type of diffusion model [49], CLIP text features, CLIP vision features, and VDVAE features [52] as latent features.
-Both methods first generate initial images from translated VAE/VDVAE features.
+In the test phase, they directly translate CLIP text features from brain activity during image perception.
+Hereafter, these two reconstruction methods will be referred to together as text-guided reconstruction methods.
+We also replicated the MindEye2 reconstruction method (Scotti et al., 2024), a more recent study evaluated using the NSD.
+Unlike the other two methods, MindEye2 uses a variant of the CLIP model’s image embeddings as latent features, implements a nonlinear translator, and utilizes specially designed diffusion-based generators rather than relying on text features as direct decoding targets.
+The StableDiffusionReconstruction method (Takagi and Nishimoto, 2023a) uses components of the Stable Diffusion model (Rombach et al., 2022), the VAE features (Kingma and Welling, 2014), and CLIP text features as latent features.
+Similarly, the Brain-Diffuser method (Ozcelik and VanRullen, 2023) utilizes components of another type of diffusion model (Xu et al., 2022), CLIP text features, CLIP vision features, and VDVAE features (Child, 2021) as latent features.
+Both methods translate brain activity into their latent features using a linear translator, and initial images are first generated from translated VAE/VDVAE features.
 These initial images are then passed through the image-to-image pipeline of the diffusion model conditioned on the translated CLIP features, producing the final reconstructed images.
 They validated the reconstruction performance using the NSD dataset, preparing training and test data based on the data split provided by the NSD study.
-Thanks to the authors of these studies who made efforts to make their datasets and scripts publicly available, we were able to conduct our replication analysis effectively.
-We compared the reconstructed results of these two text-guided reconstruction methods with those from a previous image reconstruction method, iCNN [29].
-For more information on datasets and reconstruction methods, see Material and Methods (“Datasets” or “Reconstruction methods”) or the original studies [12, 29, 44, 45].
-Observations: Failed replication and convincing reconstruction from random data
-We first confirmed the reproduction of the findings of the text-guided reconstruction methods, particularly those of the Brain-Diffuser study [45] when applied to the NSD dataset (Fig 2A).
-The reconstructed images generated by the Brain-Diffuser method effectively captured most of the layout and semantics of the test images.
-Although the reconstructed images produced by the StableDiffusionReconstruction method [44] showed slightly degraded performance compared to the original paper, they still successfully captured the semantics of the test images.
-Notably, the iCNN method [29] also performed well on the NSD dataset, with reconstructed images capturing the dominant structures of the objects within the images, consistent with the findings reported in the original study.
-Fig 2: Comparison of image reconstruction results across datasets and methods.
+Thanks to the authors’ efforts in making the datasets and scripts publicly available, we were able to conduct our replication analysis effectively.
+We compared the reconstructed results of these two text-guided reconstruction methods (additionally, MindEye2) with those from a previous image reconstruction method, iCNN (Shen et al., 2019b).
+For more information on datasets and reconstruction methods, refer to Methods (“Datasets” and “Reconstruction methods”) or the original studies (Allen et al., 2022; Shen et al., 2019b; Takagi and Nishimoto, 2023a; Ozcelik and VanRullen, 2023).
+2.1.1 Observations: Failed replication and convincing reconstruction from random data
+We first confirmed the reproduction of the findings of the original methods (Fig.
+2A).
+The reconstructed images produced by the StableDiffusionReconstruction method (Takagi and Nishimoto, 2023a) showed slightly degraded performance compared to the original paper, but still successfully captured the semantics of the test images.
+The Brain-Diffuser method (Ozcelik and VanRullen, 2023) effectively captured most of the layout and semantics of the test images when applied to the NSD dataset.
+Similarly, MindEye2 (Scotti et al., 2024) generated reconstructions that preserved key visual elements of the original stimuli.
+Notably, despite originally being validated on a different dataset (Deeprecon), the iCNN method (Shen et al., 2019b) also performed well on the NSD dataset, capturing the dominant structures of the objects.
+This performance is consistent with the findings reported in the original study.
+Fig.
+2: Comparison of image reconstruction results across datasets and methods.
 (A) Reconstruction results for the Natural Scene Dataset (NSD).
-The first row shows the original test images, followed by reconstructions using StableDiffusionReconstruction, Brain-Diffuser, and iCNN methods.
+The first row shows the original test images, followed by reconstructions using StableDiffusionReconstruction, Brain-Diffuser, MindEye2, and iCNN methods.
 (B) Reconstruction results for the Deeprecon dataset, presented in the same format as (A).
 To further investigate the generalizability of the text-guided reconstruction methods, we attempted to replicate their performance using a different dataset, Deeprecon, which was originally collected for the study by Shen et al.
-[29].
-The Deeprecon dataset was explicitly designed to avoid overlap between training and test sets, making it an ideal benchmark for evaluating the zero-shot prediction capabilities of reconstruction methods.
+(2019b).
+The Deeprecon dataset was explicitly designed to avoid overlap between training and test sets, making it a suitable benchmark for evaluating the zero-shot prediction capabilities of reconstruction methods.
 However, the original Deeprecon dataset lacked the text annotations required by the text-guided reconstruction methods.
-To facilitate a fair comparison, we collected five text annotations for each training stimulus in the Deeprecon dataset through crowd-sourcing and used them to generate CLIP text features.
-Despite including these text annotations, the text-guided reconstruction methods failed to achieve the same level of performance on the Deeprecon dataset as they did on the NSD dataset.
-The reconstructed images produced by the text-guided methods exhibited realistic appearances but suffered from largely degraded quality compared to their performance on the NSD dataset (Fig 2B).
-Notably, the text-guided reconstruction methods generated realistic reconstructions even for simple geometric shapes in the Deeprecon dataset, which deviated significantly from the original stimuli.
-In contrast, the iCNN method consistently provided faithful reconstructions for both the NSD and Deeprecon datasets, despite its simplicity compared to the text-guided methods.
-These results suggest that the text-guided methods may not generalize well across different datasets.
-The ability to generate realistic but inaccurate reconstructions, especially for simple shapes, indicates that these methods might rely more on training stimuli than on actual brain activity information.
+To enable a fair comparison, we collected five text annotations for each training stimulus in the Deeprecon dataset through crowd-sourcing and used them to generate CLIP text features.
+Despite our careful replication efforts, including preparing text annotations, both text-guided reconstruction methods and MindEye2 failed to achieve the same level of performance on the Deeprecon dataset as they did on the NSD dataset (Fig.
+2B).
+The reconstructed images produced by the text-guided methods exhibited realistic appearances but suffered from largely degraded quality compared to their performance on the NSD.
+Notably, the text-guided reconstruction methods generated realistic reconstructions even for simple geometric shapes in the Deeprecon dataset, which deviated strikingly from the original stimuli.
+Similarly, the reconstructions from MindEye2 did not resemble the test images but tended to exhibit object categories in the Deeprecon training set (frog, fighter aircraft, baby buggy, or baseball glove).
+Sample size matching between NSD and Deeprecon yielded similar reconstruction quality (Fig.
+A1), suggesting that sample size alone does not account for the poorer results on the Deeprecon.
+In contrast, the iCNN method consistently provided faithful reconstructions for both the NSD and Deeprecon datasets, despite its simplicity compared to these methods.
+These results suggest that the text-guided reconstruction methods and MindEye2 struggle to generalize across different datasets.
+The tendency to generate realistic yet inaccurate reconstructions, especially for simple shapes, indicates that these methods might rely more on learned training stimuli than on actual brain activity information.
 Upon further investigation, we noted a questionable post-hoc image selection procedure.
-In the StableDiffusionReconstruction study [44], they presented the reconstruction results by the following procedure: “We generated five images for each test image and selected the generated images with highest PSM” (PSM means perceptual similarity metrics, and it was calculated from early/middle/late layers of several image recognition deep neural networks [DNNs]), as illustrated in Fig 3A.
+In the StableDiffusionReconstruction study (Takagi and Nishimoto, 2023a), they presented the reconstruction results by the following procedure: “We generated five images for each test image and selected the generated images with highest PSM.” In their paper, PSM refers to perceptual similarity metric, which was calculated from early, middle, and late layers of several image recognition DNNs.
+This procedure is illustrated in Fig.
+3A.
 This selection might lead readers or peer reviewers, particularly those not specialized in the brain decoding field, to overestimate the effectiveness of the methods and potentially lead to a distorted understanding of the actual reconstruction performance.
-Note that the Brain-Diffuser study [45] did not execute such procedures, and in their subsequent report [53], they updated the image presentation procedure more fairly as: “we generated five images with different stochastic noise and selected three images randomly.”
-Fig 3: Analysis of post-hoc selection in the StableDiffusionReconstruction method [44].
+Note that the Brain-Diffuser and MindEye2 studies did not execute such procedures, and in their subsequent report (Takagi and Nishimoto, 2023b), they updated the image presentation procedure more fairly as: “we generated five images with different stochastic noise and selected three images randomly.”
+Fig.
+3: Analysis of post-hoc selection in the StableDiffusionReconstruction method (Takagi and Nishimoto, 2023a).
 (A) Selection procedure.
@@
 Examples of selected images are shown with the text images.
-Examples are shown as in (B).
+Examples are shown in (B).
 To examine the impact of this post-hoc selection procedure, we conducted an experiment using random brain data.
-Instead of feeding the test brain data into trained translators, we shuffled the activities within each voxel of the NSD test set independently to create random brain data.
+Instead of feeding the test brain data to trained translators, we shuffled the activities within each voxel of the NSD test set independently to create random brain data.
 Specifically, the brain data of the NSD test set is in a matrix shape, with rows representing stimulus samples and columns representing voxels.
@@
 This process was repeated for all voxels independently.
-Surprisingly, when the random brain data were input into the VAE feature translator, which contributes to producing initial images, plausible images were obtained by generating five images and selecting the best one (Fig 3B).
-Even more strikingly, when the random brain data were input into both VAE and CLIP text feature translators, we still obtained convincing results by simply generating images five times and conducting the selection mentioned above (Fig 3C).
+Surprisingly, when the random brain data were input into the VAE feature translator, which contributes to producing initial images, plausible images were obtained by generating five images and selecting the best one (Fig.
+3B).
+Even more strikingly, when the random brain data were input into both VAE and CLIP text feature translators, we still obtained convincing results by simply generating images five times and conducting the selection mentioned above (Fig.
+3C).
 These observations are inexplicable because the artificially created brain data should completely lack any information related to the original visual stimuli.
-These observations raise perplexing questions about the performance and generalizability of recent text-guided reconstruction methods.
-The performance significantly deteriorated when we changed from the NSD to Deeprecon dataset.
-This highlights the need to better understand why these methods succeed with the NSD.
+These observations raise perplexing questions about the performance and generalizability of recent reconstruction methods.
+Their performances largely deteriorated when we switched the evaluated dataset from the NSD to Deeprecon, highlighting the need to better understand why these methods succeed with the NSD.
 Moreover, the ability to obtain plausible reconstructions from random brain data by merely generating multiple images and selecting the best ones suggests that there may be fundamental issues with both the evaluation dataset and the components of the reconstruction methods themselves.
 In the following sections, we will thoroughly investigate the potential problems associated with the NSD dataset and each component of the text-guided reconstruction pipeline.
-Lack of diversity in the stimulus set
+2.1.2 Lack of diversity in the stimulus set
 First, we examined the characteristics and limitations of the NSD dataset itself.
 To characterize the diversity of stimuli in the datasets and their latent representations, we focused on the CLIP text features, which are used as latent features in text-guided reconstruction methods.
-We employed uniform manifold approximation and projection (UMAP) [54] to visualize the CLIP text features of the NSD stimuli (see Materials and Methods “UMAP visualization”).
-The visualization revealed approximately $40$ distinct clusters, with considerable overlap between the training and test sets (Fig 4A).
-Interestingly, we were able to describe the stimulus images in each cluster using a single semantic word, such as airplane, giraffe, or tennis.
-Despite the NSD containing around $30,000$ brain samples per subject, the diversity of the presented stimuli was quite limited to just around $40$ semantic concepts.
-Note that we performed UMAP visualization using the parameters recommended in the official guide for clustering, but a similar cluster structure was also observed with the default parameters (S1 Fig).
-In contrast, the Deeprecon dataset, which was specifically designed to differentiate object categories between training and test data, exhibited less overlap between the two sets (S2 Fig).
-Fig 4: Dataset diversity and similarity between training and test stimuli.
-(A) UMAP visualization of CLIP text features in the NSD.
+We employed uniform manifold approximation and projection (UMAP) (McInnes et al., 2018) to visualize the CLIP text features of the NSD stimuli (see Methods “UMAP visualization”).
+The visualization revealed approximately $40$ distinct clusters, with considerable overlap between the training and test sets (Fig.
+4A).
+Interestingly, we were able to describe the stimulus images in each cluster using a single semantic label, such as airplane, giraffe, or tennis.
+Despite the NSD containing around $30,000$ brain samples per subject, the diversity of the presented stimuli was quite limited to just around $40$ semantic categories.
+Here, we performed UMAP visualization using the parameters recommended in the official guide for clustering.
+Even with the default UMAP parameters, a similar cluster structure was observed (Fig.
+A2).
+In contrast, the Deeprecon dataset, which was specifically designed to differentiate object categories between training and test data, exhibited less overlap between the two sets (Fig.
+A3; see Fig.
+A4 for the latent features used in MindEye2).
+Fig.
+4: Dataset diversity and similarity between training and test stimuli.
+(A) UMAP visualization of CLIP text features of the NSD stimuli.
 The center figure shows the scatter plot of the UMAP embedding of CLIP text features.
@@
 This analysis was performed for the NSD (B) and Deeprecon (C) datasets, using the same procedure.
-To further investigate the similarity between the training and test stimuli, we analyzed using a state-of-the-art perceptual similarity metric called DreamSim [55].
-For each test image in the NSD, we used DreamSim to identify the most similar training images based on their perceptual similarity scores.
-We found that the training images extracted using DreamSim were highly similar to the test images in the NSD, not only in terms of semantic concepts but also in terms of overall layout and visual composition (Fig 4B).
-These findings suggest that the distribution of NSD test images is heavily biased towards that of the training images, with a significant overlap in the visual and semantic features present in both sets.
+To further investigate the similarity between the training and test stimuli, we analyzed stimulus images using DreamSim, a state-of-the-art perceptual similarity metric (Fu et al., 2024).
+DreamSim was used to identify the most perceptually similar training images for each test image.
+We found that the training images identified by DreamSim metric were highly similar to the test images in the NSD, not only in semantic labels but also in overall layout and visual composition (Fig.
+4B).
+In contrast, the same analysis on the Deeprecon dataset revealed that its training images were substantially different from the test images (Fig.
+4C).
+To further assess whether the similarity between NSD training and test images is excessively high, we measured the similarity between the two sets.
+As a reference, we also measured the similarity between NSD training images and a large-scale independent dataset (CC3M; Sharma et al., 2018).
+This analysis revealed that the NSD test set contained stimuli that were highly similar to the training set, compared to those in the independent dataset (Fig.
+A5).
+In contrast, the Deeprecon test set, due to its carefully designed training–test split, exhibited a similar level of similarity to the independent dataset.
+These findings suggest that the distribution of NSD test images is heavily biased toward that of the training images, with a significant overlap in the visual and semantic features present in both sets.
 Such a strong bias raises concerns about the actual reconstruction performance of methods evaluated on this dataset.
-The impressive results may be largely attributable to the methods’ ability to memorize and reproduce the characteristics of the training stimuli rather than generalizing them to novel stimuli.
-In contrast, when we applied the same DreamSim-based analysis to the Deeprecon dataset, we found that the extracted training images were substantially different from the test images (Fig 4C).
-This observation supports the careful design of the train-test split in the Deeprecon dataset, ensuring that the test images contain novel visual and semantic features that are not well-represented in the training set.
-The distinct differences between the NSD and Deeprecon datasets in terms of stimulus similarity highlight the importance of carefully designing evaluation benchmarks that can effectively assess the generalization capabilities of visual image reconstruction methods.
-Failed zero-shot prediction in the feature space
-Given the significant semantic and structural similarities between the NSD test images and the training images, it is crucial to assess whether the predictions derived from brain activity truly demonstrate generalization beyond the training set (i.e., zero-shot prediction).
-To evaluate the zero-shot prediction capability of the translators used in the reconstruction methods, we conducted an $(N+1)$-way identification analysis, where $N$ represents all of the training samples ($8,859$ samples for subject 1 in the NSD) and $1$ refers to the test sample (Fig 5A).
-Specifically, we calculated the similarity (correlation) between the predicted (translated) features and the true features of corresponding test samples, as well as the similarities between the predicted features and the true features corresponding to each of the N training samples.
-We then examined whether the predicted features correctly identified (showed the highest similarity to) the test true features.
-We calculated the percentage of predicted features that successfully identified the corresponding test sample in the NSD dataset.
-Fig 5: Evaluation of zero-shot sample identification performance.
+The convincing reconstruction results with the NSD may be largely attributed to the methods’ tendency to replicate specific characteristics observed in the training set, potentially at the expense of generalizing to novel stimuli.
+The distinct differences between the NSD and Deeprecon datasets in terms of stimulus similarity highlight the importance of carefully designing evaluation benchmarks to rigorously evaluate the generalization capabilities of visual image reconstruction studies.
+2.1.3 Failed zero-shot prediction in the feature space
+Given that many of the NSD test images closely resemble those in the training set, it is uncertain whether the translator’s predictions genuinely capture new, unseen stimuli (i.e., zero-shot generalization) or simply replicate features of training images that share similar semantics and visual structure.
+To probe this, we evaluated the zero-shot prediction capability of the CLIP feature translators by conducting an $(N+1)$-way identification analysis, where $N$ represents the entire training set ($8,859$ samples for subject 1 in the NSD) and $1$ is the target test sample (Fig.
+5A).
+Concretely, we measured the correlation between each translator’s predicted features and the true features of (1) the correct test sample and (2) every training sample.
+We then asked whether the predicted features were most similar to the true test features, thereby correctly identifying them.
+Identification performance above chance suggests that the translator captures information specific to the test sample beyond the learned training patterns, supporting zero-shot prediction.
+Conversely, poor performance indicates that the translator does not predict the unique properties of the test sample.
+Fig.
+5: Evaluation of zero-shot sample identification performance.
 (A) Schematic diagram of novel sample identification.
@@
 The chance performance level is thus $1/8860$.
-The performance of the CLIP features was nearly 00% (Fig 5B).
-The CLIP features predicted from brain activity were unable to accurately identify novel test samples, even within the distribution covered by the training set.
-This extremely low performance suggests that the CLIP feature translator fails to achieve zero-shot prediction.
-This finding raises questions about the effectiveness of CLIP features for true zero-shot prediction tasks in brain decoding.
-In contrast, the VGG19 features [56] employed in the iCNN method demonstrated moderate identification performance at the latent features on the intermediate layers (Fig 5C).
-This finding suggests that the VGG19 features possess a certain degree of generalizability, enabling them to identify test samples not encountered during training.
-Furthermore, we investigated whether the CLIP feature translator enables the prediction of novel semantic concepts not in the training.
-We redesigned the dataset split to ensure no semantic concepts were shared between them as in the previous zero-shot prediction study [17, 18].
-We first applied $k$-means clustering to the UMAP embedding space of the NSD’s CLIP text features (S3 FigA).
-We set the number of clusters as $40$ based on visual inspection of the UMAP results.
-Based on these clustering results, we performed a hold-out analysis: when predicting samples within a cluster (e.g., the ski cluster), we excluded samples from that cluster in the training set (Fig 6A right; hold-out split condition).
-As a control, we also prepared a naive data-split condition where the training sample size is the same as in the hold-out split condition but allows overlapping semantic concepts (Fig 6A left; naive split condition).
-When we visualized the properties of predicted features in hold-out analysis by transforming them into the previous UMAP embedding space (Fig 4A), we observed that the predicted features tend to diverge largely from their original clusters and move into other clusters (Fig 6B).
-Fig 6: Cluster hold-out analysis.
+The results show that the identification performance of CLIP feature translators was nearly 00% (Fig.
+5B).
+This poor performance suggests that the CLIP feature translator captures only rough semantic categories encountered in the training set rather than fine-grained, instance-level details.
+This finding raises a question about the effectiveness of CLIP features for zero-shot prediction tasks in brain decoding.
+By contrast, the VGG19 features (Simonyan and Zisserman, 2015), as used in the iCNN method, demonstrated moderate identification performance at the intermediate layers (Fig.
+5C).
+This success could be attributed to the compositional representation of VGG19’s intermediate features.
+Unlike CLIP features, which are optimized for semantic alignment across vision and language, VGG19’s intermediate features, extracted through convolutional layers, contain primarily visual representations that retain sufficient local structure.
+This characteristic allows the translator to generalize to novel images by combining learned local spatial features rather than relying solely on semantic similarity.
+Consequently, VGG19’s compositional features may help achieve a certain degree of zero-shot prediction, distinguishing new images from closely resembling training examples.
+Furthermore, we investigated whether the CLIP feature translator enables the prediction of novel semantic clusters not in the training.
+We redesigned the dataset split to ensure no semantic clusters were shared between them as in previous zero-shot prediction studies (Mitchell et al., 2008; Brouwer and Heeger, 2009).
+We first applied $k$-means clustering to the UMAP embedding space of the NSD’s CLIP text features (Fig.
+A6A).
+We set the number of clusters as $40$ based on our visual inspection of the UMAP results.
+Based on these clustering results, we performed a hold-out analysis: when predicting samples within a cluster (e.g., ski cluster), we excluded samples of that cluster from the training set (Fig.
+6A right; hold-out split condition).
+As a control, we also prepared a naive data split condition where the training sample size is the same as in the hold-out split condition but allows overlapping semantic clusters (Fig.
+6A left; naive split condition).
+When we visualized the properties of predicted features in hold-out analysis by transforming them into the previous UMAP embedding space (Fig.
+4A), we observed that the predicted features tend to diverge largely from their original clusters and move into other clusters (Fig.
+6B).
+Fig.
+6: Cluster hold-out analysis.
 (A) Hold-out procedure.
-When predicting a test sample from a specific cluster, all training samples from that cluster are excluded.
-The naive split condition serves as a control, using the same number of training samples but allowing cluster overlap between training and test sets.
-(B) Examples of predictions.
-Under the hold-out split condition, gray and dark orange points represent training and test samples, respectively.
-Light orange points indicate translated (predicted) latent features.
-Black lines connect true features to their corresponding predictions.
+For each test sample from a given cluster, all training samples in that cluster are excluded.
+The naive split (control) uses an equal number of training samples but allows overlap between training and test clusters.
+(B) Prediction examples.
+Gray and dark orange points denote training and test samples in the hold-out condition, respectively; light orange points are the predicted latent features, with black lines connecting true features to their predictions.
 (C) Cluster identification procedure.
-The predicted features are compared with the average features of each cluster, and the most similar cluster is selected.
+The predicted features are compared with the cluster centers (average features in each cluster), and the most similar cluster is selected.
 (D) Cluster identification results.
@@
 (E) Pairwise sample identification procedure.
-Predicted features are compared with the true test sample and another randomly chosen test sample.
-A sample more similar to the prediction is selected.
+Predicted features are compared with the true test features and with one of the other test features.
+The sample with features more similar to the prediction is selected.
+This procedure is repeated for all other test samples, and the proportion of correctly identified true test features is calculated.
 (F) Pairwise sample identification results.
-The accuracy means the average winning rate of the true samples against all of the other test samples.
 Results are presented similarly to cluster identification, with a chance level of $1/2$.
@@
 The first metric is cluster identification accuracy.
-Cluster identification accuracy focuses on evaluating the translator’s ability to generate features that correctly identify the semantic cluster to which a test sample belongs (Fig 6C).
-In this analysis, we compare the predicted features of each test sample to the average features of the training samples within each semantic cluster.
+Cluster identification accuracy focuses on evaluating the translator’s ability to predict features that correctly identify the semantic cluster to which a test sample belongs (Fig.
+6C).
+In this analysis, we calculated the similarity between the predicted features of test samples and the average features of the training samples within each semantic cluster.
 The accuracy is then calculated as the percentage of predicted features that successfully identify the original semantic cluster of their corresponding test samples in the latent feature space.
-The cluster identification accuracy in the hold-out split condition exhibited a substantial drop compared to the naive split condition across all semantic clusters (Fig 6D).
-Notably, the cluster identification accuracy was frequently 00% in the hold-out split condition (see also S3 Fig.B for all cluster results).
-These results expose a severe limitation of the translator when dealing with novel semantic concepts absent from the training set.
-This suggests that the CLIP feature translator primarily functions as a “classifier;” its prediction (translation) heavily relies on predefined semantic features used in a training phase rather than generalizing to new concepts.
-The second metric is pairwise sample identification accuracy.
-Pairwise sample identification accuracy is a commonly used metric in the evaluation of feature prediction and reconstruction performance [39, 29, 34, 36, 40, 41, 44, 45, 46, 57, 58].
-This analysis assesses the translator’s ability to predict target features that are more similar to the actual test features than to the features of other samples in the test set (Fig 6E).
-For each test sample, we compare its predicted features with two sets of features: the actual test features and the features of another randomly selected sample from the test set.
-The accuracy is then calculated as the average winning rate of the predicted features against all other test samples, indicating the proportion of cases where the predicted features are more similar to the actual test features than those of other samples.
-Intriguingly, despite the feature translator’s complete failure to identify semantic concepts in the hold-out split condition, the pairwise sample identification accuracy often surpassed the chance level across the semantic clusters (Fig 6F).
-The insensitivity of the pairwise sample identification accuracy to the translator’s inability to capture semantic concepts can be explained by its focus on relative similarity.
-Specifically, the similarity compares the predicted features to the actual test features, without taking into account the broader general context.
-This can lead to an inflated pairwise sample identification accuracy, potentially misrepresenting the actual reconstruction performance and the translator’s ability to generalize to novel clusters.
-This observation raises concerns about the adequacy of using identification accuracy as the sole metric for evaluating reconstruction performance, a practice that has been widely adopted in many studies.
-Failed recovery of a stimulus from its latent features
+The cluster identification accuracy in the hold-out split condition exhibited a substantial drop compared to the naive split condition across all semantic clusters (Fig.
+6D).
+Notably, the cluster identification accuracy was frequently 00% in the hold-out split condition (see Fig.
+A6B for all cluster results).
+These results expose a severe limitation of the translator when dealing with novel semantic clusters absent from the training set.
+This suggests that the CLIP feature translator primarily functions as a “classifier;” its prediction (translation) heavily relies on predefined semantic features used in a training phase rather than generalizing to new semantic categories.
+The second metric is pairwise sample identification accuracy, a commonly used metric in the evaluation of feature prediction and reconstruction performance (Beliy et al., 2019; Shen et al., 2019a, b; Mozafari et al., 2020; Qiao et al., 2020; Ren et al., 2021; Gaziv et al., 2022; Takagi and Nishimoto, 2023a; Ozcelik and VanRullen, 2023; Scotti et al., 2023; Denk et al., 2023; Koide-Majima et al., 2024).
+This analysis assesses whether the translated features for a given test sample are more similar to its actual features than to those of a randomly chosen sample in the test set (Fig.
+6E).
+The accuracy is calculated as the average winning rate of the predicted features against all the test samples, reflecting how often the predicted features are closer to the correct sample than to any alternative.
+Intriguingly, even though the feature translator completely failed to identify the true cluster in the hold-out split condition, pairwise sample identification accuracy often exceeded chance across clusters even in the hold-out split condition (Fig.
+6F).
+This discrepancy arises because pairwise identification is based on relative similarity: if the translated features are “less wrong” for the true sample than for another test sample, they will still be deemed a match.
+As a result, merely success by a smaller margin than the alternatives can inflate the overall identification score, giving a misleading impression of the translator’s ability to capture new clusters.
+This observation suggests that relying on pairwise identification accuracy alone, a common practice in many studies, may overestimate reconstruction performance.
+This issue is further discussed in “Caveat with evaluation by pairwise identification”.
+When applying the hold-out split procedure to the full reconstruction pipeline, StableDiffusionReconstruction showed noticeable degradation, whereas Brain-Diffuser remained more robust (Fig.
+A7).
+This robustness arises because, even in the hold-out split, the training and test stimuli remain highly similar in the NSD (Fig.
+A7A).
+As a result, Brain-Diffuser reconstructions still capture the general visual layout of the original images, with only minor semantic differences (Fig.
+A7C: 1-cluster hold-out).
+However, when removing additional clusters such that the similarity between training and test samples approximates that of a completely independent dataset (Fig.
+A7A: similarity-matched hold-out), the reconstructions, though still realistic, deviate considerably from the original images, both semantically and visually (Fig.
+A7C: similarity-matched hold-out).
+This observation suggests that training–test similarity in the NSD played a crucial role in driving text-guided reconstruction performance.
+When this similarity is reduced to the level of an independent dataset, the ability to reconstruct meaningful details of unseen images diminishes, confirming that the performance of these methods relies on such overlap.
+2.1.4 Failed recovery of a stimulus from its latent features
 Finally, we conducted a rigorous evaluation of the generator component, which typically consists of diffusion models in text-guided reconstruction methods.
 To ensure that a visual image reconstruction method has the potential to faithfully reproduce an individual’s perceived visual experiences, it is crucial that the method can recover the original images with a high degree of perceptual similarity when the neural translation from brain activity to latent features is perfect.
-However, it has been unclear whether recent text-guided reconstruction methods meet this fundamental requirement.
+However, it has remained unclear whether recent text-guided reconstruction methods meet this fundamental requirement.
 To address this question, we performed a recovery check analysis by reconstructing images using the true latent features of target images.
 Instead of using latent features translated from brain activity, we directly input the latent features derived from the target images into the generator.
-Text-guided methods produced images semantically similar to targets but not perceptually similar (Fig 7A), while the iCNN method yielded results that closely resembled the actual target images.
+Text-guided methods produced images semantically similar to targets but not perceptually similar (Fig.
+7A), while the iCNN method yielded results that closely resembled the actual target images.
 These findings suggest that text-guided reconstruction methods may prioritize semantic similarity over perceptual accuracy.
 In contrast, the iCNN method appears to have a superior ability to capture and replicate original visual content, indicating potential advantages in preserving fine-grained visual details.
-Fig 7: Recovery check.
+Fig.
+7: Recovery check.
 (A) Reconstruction from the true latent features.
@@
 Each row represents a different test image.
-The reconstruction results reveal the upper bound of reconstruction performance for each method and highlight potential limitations in the latent feature representations or generative processes.
-To investigate the recovery performance further, we conducted a recovery check on each latent feature of the Brain-Diffuser method (Fig 7B).
+The reconstruction results indicate the upper bound of reconstruction performance for each method and highlight potential limitations in the latent feature representations or generative processes.
+To investigate the recovery performance further, we conducted a recovery check on each latent feature of the Brain-Diffuser method (Fig.
+7B).
 Interestingly, reconstructions from VDVAE features, which are used for generating initial images in the Brain-Diffuser, exhibited a high degree of similarity to the target images.
@@
 These findings suggest that text-guided reconstruction methods may not be well-suited for visual image reconstruction tasks, as they fail to faithfully recover the original visual images.
-Instead, they tend to create images based on their latent features, such as CLIP features, which can lead to a phenomenon known as “hallucination” in the field of generative AIs.
-Hallucination refers to an output that appears plausible but is actually incorrect or misleading, raising concerns about the reliability and accuracy of the model [59].
-Text-guided reconstruction methods seem to prioritize generating semantically similar images rather than faithfully reconstructing the visual content perceived by the individual.
-The above findings may provide an explanation for why the text-guided reconstruction methods performed well only on the NSD dataset (Fig 2A).
-The results of the case study demonstrated that the text-guided reconstruction methods struggle to reconstruct (Fig 2B) or identify (Fig 6) test samples that lie beyond the distribution of the training set.
+Instead, they tend to create images based on their semantic features, such as CLIP features, which can lead to a phenomenon known as “hallucination” in the field of generative AIs.
+Hallucination refers to an output that appears plausible but is actually incorrect or misleading, raising concerns about the reliability and accuracy of the model (Rawte et al., 2023).
+Text-guided reconstruction methods seem to prioritize generating semantically similar images rather than faithfully reconstructing the visual content perceived by the individual (see Fig.
+A8 for the recovery with MindEye2’s generators).
+The above findings may provide an explanation for why the text-guided reconstruction methods performed well only on the NSD dataset (Fig.
+2A).
+The results of the case study demonstrated that the text-guided reconstruction methods struggle to reconstruct (Fig.
+2B) or identify (Fig.
+6D) test samples that lie beyond the distribution of the training set.
 Such limitations suggest that these methods lack true generalization capabilities and are unable to accurately reconstruct novel visual experiences that differ significantly from the examples they were trained on.
-Moreover, even when the test samples belonged to the same distribution as the training set, the translators of the text-guided reconstruction methods had difficulty correctly identifying those test samples (Fig 5).
+Moreover, even when the test samples belonged to the same distribution as the training set, the translators had difficulty correctly identifying those test samples (Fig.
+5).
 This observation indicates that the translators may not have learned a sufficiently robust and generalizable mapping between brain activity patterns and the corresponding latent features, further limiting their ability to faithfully reconstruct the perceived visual experiences.
-This case study also revealed that the NSD test stimuli are highly similar to the training set, with a significant overlap in their visual and semantic features (Fig 4).
+This case study also revealed that the NSD test stimuli are highly similar to the training set, with a significant overlap in their visual and semantic features (Fig.
+4).
 Given this similarity, the impressive reconstruction results achieved by the recent text-guided reconstruction methods on the NSD dataset should not be interpreted as evidence of zero-shot reconstruction capabilities.
 Instead, a more plausible interpretation is that these methods primarily function as a combination of “classification” and “hallucination.”
-In this context, the translator component of the text-guided reconstruction methods functions primarily as a classifier, predicting categorical semantic information present in the training phase rather than capturing the fine-grained details of the visual experience.
-This limitation may explain why convincing reconstructions can be obtained even from random brain data through post-hoc selection (Fig 3B).
+In this context, the CLIP feature translator in the text-guided reconstruction methods functions primarily as a classifier, predicting categorical semantic information present in the training phase rather than capturing the fine-grained details of the visual experience.
+This limitation may explain why convincing reconstructions can be obtained even from random brain data through post-hoc selection (Fig.
+3BC).
 Due to the limited variety in the outputs generated by the reconstruction models, repeated trials and post-hoc selection can eventually find images that are semantically and visually similar to the target stimulus.
@@
 This phenomenon of hallucination raises serious concerns about the reliability and validity of the text-guided reconstruction methods when evaluated on the NSD dataset.
-Although we evaluated only two reconstruction methods in this case study, it is important to recognize that any reconstruction methods evaluated only by NSD [46, 48, 60] can potentially have similar classification and hallucination problems.
-Formal analysis and simulation
+Although we particularly evaluated three reconstruction methods in this case study, it is important to recognize that any reconstruction methods evaluated only by NSD (Scotti et al., 2023; Quan et al., 2024) can potentially have similar classification and hallucination problems due to the limited diversity of the NSD.
+Furthermore, recent highly realistic reconstruction methods leveraging generative AI models (Chen et al., 2023; Bai et al., 2024; Benchetrit et al., 2024) should also be carefully validated to ensure their performance is not overly dependent on dataset biases (Fig.
+A5).
+2.2 Formal analysis and simulation
 Building on the issues identified in our case study, we examine these challenges in a more general context.
-We have identified several issues with these text-guided reconstruction methods and the dataset, including the cluster structure of CLIP latent features, the lack of diversity in NSD, and the misspecification of the latent representation for image reconstruction.
+We have identified several issues with these text-guided reconstruction methods and the dataset, including the cluster structure of CLIP latent features, the lack of diversity in the NSD, and the misspecification of the latent representation for image reconstruction.
 These issues resulted in the inability of diffusion models to faithfully recover the original images from their latent features.
@@
 Instead, these issues likely reflect more fundamental problems that can arise in the development and evaluation of brain decoding and visual image reconstruction methods.
-Thus in this section, we extend the problems identified in the case study into formal analyses and simulations in generalized settings, aiming to provide a more comprehensive understanding of the factors that contribute to the limitations of current reconstruction methods and explore strategies for mitigating these issues.
-Output dimension collapse
+Thus, in this section, we extend the problems identified in the case study into formal analyses and simulations in generalized settings, aiming to provide a more comprehensive understanding of the factors that contribute to the limitations of current reconstruction methods and explore strategies for mitigating these issues.
+2.2.1 Output dimension collapse
 Multivariate linear regression models are widely used in constructing decoding and encoding models of the brain.
@@
 However, this expectation is not generally true.
-This is particularly evident when input variables are shared [33, 35, 34, 44, 45].
+This is particularly evident when input variables are shared (Seeliger et al., 2018; Ozcelik et al., 2022; Mozafari et al., 2020; Takagi and Nishimoto, 2023a; Ozcelik and VanRullen, 2023).
 In this section, we demonstrate this assertion by examining the problem of predicting multiple targets using linear regression models with shared inputs.
@@
 Next, we consider a scenario where the diversity of the target features is small.
-This situation can arise when the feature space exhibits a clustered structure and the training data lacks sufficient diversity, as observed in the case study with the CLIP text features and the NSD dataset (Fig 4A).
+This situation can arise when the feature space exhibits a clustered structure and the training data lacks sufficient diversity, as observed in the case study with the CLIP text features and the NSD dataset (Fig.
+4A).
 When the training features have limited diversity, the predicted values from brain activity, which are represented as linear combinations of these target features, also become constrained.
 Consequently, the prediction from brain data to target features effectively becomes a projection onto a low-dimensional subspace formed by the training data.
-Fig 8: Demonstration of output dimension collapse in feature prediction.
+Fig.
+8: Demonstration of output dimension collapse in feature prediction.
 The left panel shows the distribution of source brain activity data in the first two dimensions of the original high-dimensional space.
@@
 We generated clustered features by sampling from a Gaussian mixture distribution in a high-dimensional space.
-The corresponding brain activity samples were generated from the latent feature samples multiplied by teacher weight and adding observation noise (See Materials and Methods “Simulation with clustered data”).
+The corresponding brain activity samples were generated from the latent feature samples multiplied by teacher weight and adding observation noise (see Methods “Simulation with clustered data”).
 Student weights were obtained by training a linear regression model to predict the clustered feature values from the corresponding brain data.
-We then projected the randomly generated brain samples into latent feature space by learned regression weight.
+We then projected the randomly generated brain samples into the latent feature space using the learned regression weights.
 To visualize the high-dimensional predicted patterns effectively, we projected the predicted features onto the PCA spaces derived from the training features and showed the first two dimensions.
-The simulation results clearly demonstrate the impact of clustered features on the predicted values (Fig 8).
+The simulation results clearly demonstrate the impact of clustered features on the predicted values (Fig.
+8).
 The trained linear regression model projects arbitrary brain data onto the subspace defined by the latent features in the training set, resulting in predicted values that are confined to the vicinity of the training clusters.
 This observation highlights the limitation of training linear regression models with clustered features as prediction targets: the trained model’s predictions are inherently constrained by the diversity and structure of the training data.
-We refer to this phenomenon as “output dimension collapse,” where the model’s predictions become confined to a limited subspace of the output feature space.
+We refer to this phenomenon as “output dimension (or domain) collapse,” where the model’s predictions become confined to a limited subspace (or subdomain) of the output feature space.
 It has important implications for the generalization capability of linear regression models in the context of brain decoding and visual image reconstruction.
@@
 Consequently, the translator’s outputs become confined to patterns similar to those in the training set, severely limiting the model’s ability to predict novel or out-of-distribution samples.
-Output dimension collapse may explain why plausible reconstructions were obtained even from random brain data by merely generating images several times in the case study (Fig 3).
-The lack of semantic diversity in NSD causes the translator to adapt only to the feature patterns of the training set, restricting its outputs to the subspace formed by the training data.
+Output dimension collapse may explain why plausible reconstructions were obtained even from random brain data by merely generating images several times in the case study (Fig.
+3).
+The lack of semantic diversity in the NSD causes the translator to adapt only to the feature patterns of the training set, restricting its outputs to the subspace formed by the training data.
 As a result, convincing images could be found even from random data through questionable post-hoc selection.
-It is crucial to understand that output dimension collapse may not inherently be a limitation of linear regression models.
-This phenomenon can manifest across various regression models, including nonlinear approaches such as multi-layer perceptrons (MLPs) [46] and transformers [48].
-Nonlinear models may be more susceptible to output dimension collapse due to their enhanced flexibility in adapting to training data.
-It is also important to recognize that the mathematical formulation in Eqs 1–3 assumes that all input variables are shared across target variables.
+It should be noted that this phenomenon is not inherently limited to linear regression models; it can occur in various multivariate regression models, including multi-layer neural networks.
+In fact, when we replaced the nonlinear translator in the MindEye2 reconstruction with a linear translator, the reconstructions’ bias toward object categories in the training set was substantially reduced (Fig.
+A9).
+This observation suggests that nonlinear models are more susceptible to the collapse of the output domain due to their greater flexibility in fitting to training data.
+It is also important to recognize that the mathematical formulation in Eqs.
+1–3 assumes that all input variables are shared across target variables.
 If each target is predicted from a distinct set of input variables through feature (voxel) selection, the predictions can become more independent, potentially mitigating output dimension collapse.
-This approach has been utilized in the field since its early days, with techniques such as sparse voxel selection and modular modeling [28, 29, 61, 37].
-Simulation with clustered features: What makes prediction compositional?
-The case study revealed that the NSD exhibits limited diversity (Fig 4) and poses difficulties for zero-shot prediction (Fig 5 and Fig 6D).
+This approach has been utilized in the field since its early days, with techniques such as sparse voxel selection and modular modeling (Miyawaki et al., 2008; Yamashita et al., 2008; Fujiwara et al., 2013; Shen et al., 2019b).
+2.2.2 Simulation with clustered features: What makes prediction compositional?
+The case study revealed that the NSD exhibits limited diversity (Fig.
+4) and poses difficulties for zero-shot prediction (Fig.
+5 and Fig.
+6D).
 These observations suggest that the translator of CLIP features suffers from output dimension collapse due to the lack of semantic diversity in the NSD.
@@
 We first generated feature data ${\mathbf{y}}\in\mathbb{R}^{D}$ then made the input brain data ${\mathbf{x}}\in\mathbb{R}^{D}$ by translating ${\mathbf{y}}$ with the teacher weight and added observation noise.
-To simulate a situation where the dataset has cluster structures and to control diversity effectively, the training feature vector $\mathbf{y}\in\mathbb{R}^{D}$ was generated from a $D$-dimensional Gaussian mixture (Fig 9A).
+To simulate a situation where the dataset has cluster structures and to control diversity effectively, the training feature vector $\mathbf{y}\in\mathbb{R}^{D}$ was generated from a $D$-dimensional Gaussian mixture (Fig.
+9A).
 We trained a ridge regression model on large training data samples and obtained the student weight.
 To simulate the situation where the trained model encounters clusters that are not available at the training phase, we used two types of test samples, in-distribution and out-of-distribution test samples: in-distribution test samples were generated from one of the clusters used in the training set, whereas out-of-distribution (OOD) samples were generated from the novel cluster that is not included in the training set.
-For these two types of predicted features, we calculated the cluster identification accuracy (Fig 6C) by using $C+1$ cluster centers: $C$ centers from the training set and one cluster center of the OOD test set.
-Fig 9: Simulation analysis of predicting features with a cluster structure (A) Illustration of target latent features setting.
-The latent features were generated from Gaussian mixture distributions.
-$\sigma_{\mathrm{intra}}$ is a parameter that controls the scale of the samples within a cluster.
-$\sigma_{\mathrm{inter}}$ controls the scale between the cluster centers.
-In-distribution test samples are generated from one of the clusters in the training set.
-Out-of-distribution (OOD) test samples are generated from a novel cluster not included in the training set.
-(B) Cluster identification accuracy.
-The $x$-axis represents the number of clusters used in the training, and the $y$-axis shows the cluster identification performance.
-The green line represents the results of in-distribution test samples, and the orange line shows the results for OOD samples.
+For these two types of predicted features, we calculated the cluster identification accuracy (Fig.
+6C) by using $C+1$ cluster centers: $C$ centers from the training set and one cluster center of the OOD test set.
+Fig.
+9: Simulation analysis of predicting cluster-structured features.
+(A) Illustration of target latent features.
+The latent features were generated from Gaussian mixture distributions, with $\sigma_{\mathrm{intra}}$ controlling within-cluster spread and $\sigma_{\mathrm{inter}}$ controlling inter-cluster scaling.
+In-distribution test samples are generated from training clusters while Out-of-distribution (OOD) test samples come from novel clusters.
+(B) Cluster identification accuracy for various numbers of training clusters.
+The $x$-axis represents the number of training clusters, and the $y$-axis shows the cluster identification performance.
+The green and orange lines indicate results for in-distribution and OOD samples, respectively.
 The dashed curve indicates the chance level.
@@
 The lower one shows that the training data covers only the effective axes of the latent feature space, leading to a linear order relative to the feature dimension.
-(D) Relation between latent feature dimension and the sufficient number of clusters for generalization.
+(D) Sufficient number of clusters for generalization as a function of latent feature dimension.
 The $x$-axis represents the dimension of the target latent features.
-The $y$-axis shows the number of clusters that achieve above $0.5$ cluster identification accuracy.
-The three lines indicate the results from different ratios of $\sigma_{\mathrm{intra}}^{2}/\sigma_{\mathrm{inter}}^{2}$.
+The $y$-axis shows the number of clusters achieving above $0.5$ cluster identification accuracy, with curves for different $\sigma_{\mathrm{intra}}^{2}/\sigma_{\mathrm{inter}}^{2}$ ratios.
 (E) Example of successful prediction beyond the training distribution in a 2D output feature space.
-Dark green points represent the in-distribution target features, with their corresponding light green points showing the model’s predictions.
-Dark orange points denote the OOD target features, with light orange points indicating their predictions.
-Despite the OOD target features (dark orange) not being included in the training clusters, they are accurately predicted (light orange), demonstrating the model’s generalization ability.
-This successful generalization highlights the power of compositional prediction, where the model effectively combines individual dimensions to predict novel feature combinations outside the training distribution.
-We first examined zero-shot prediction performance for different numbers of training clusters (Fig 9B).
+In-distribution and OOD target features are depicted in dark green and dark orange, and the model’s predictions of these features are depicted in light green and light orange, respectively.
+Despite the OOD target features (dark orange) not being included in the training clusters, they are accurately predicted (light orange), demonstrating the model’s generalization ability by combining learned feature dimensions.
+We first examined zero-shot prediction performance for different numbers of training clusters (Fig.
+9B).
 We fixed the feature dimension $D$ and the cluster variance ratio constant.
@@
 When the number of training clusters was small, the cluster identification accuracy was 00%.
-This result indicates that the behavior of the translator became more similar to a classifier, making it difficult to generalize beyond the training set, as observed in NSD cases (see also Fig 6D).
+This result indicates that the behavior of the translator became more similar to that of a classifier, making it difficult to generalize beyond the training set, as observed in the NSD cases (see Fig.
+6D).
 On the other hand, as the number of clusters in the training data increased, it became possible to identify the novel clusters, achieving the same performance as in-distribution test samples.
@@
 All of these results were obtained with a sufficiently large amount of training data, and the number of training clusters was varied while keeping the amount of training data fixed.
-Also, we observed qualitatively similar results in increasing the data diversity by controlling the cluster variance ratio and keeping the number of dimension and training clusters constant (S4 Fig.A).
+Also, we observed qualitatively similar results in increasing the data diversity by controlling the cluster variance ratio while keeping the number of dimensions and training clusters constant (Fig.
+A10A).
 Next, we investigate how diverse the training data needs to be to ensure sufficient generalization.
-There are two possible scenarios for diversifying training samples: either by densely sampling the entire target feature space so that there are no remaining gaps (Fig 9C; top) or by uniformly sampling to the extent that it covers the entire dimension of the target feature space (Fig 9C; bottom).
+There are two possible scenarios for diversifying training samples: either by densely sampling the entire target feature space so that there are no remaining gaps (Fig.
+9C; top) or by uniformly sampling to the extent that it covers the entire dimension of the target feature space (Fig.
+9C; bottom).
 The former scenario requires an exponentially larger number of samples/clusters relative to the dimension, whereas the latter only requires up to a linear order.
-We sought to reveal which of these two scenarios was more likely to be true by varying the dimensions of the feature space and identifying the number of clusters required for generalization (S4 Fig.
-B for identification accuracy in each condition).
+We sought to reveal which of these two scenarios was more likely to be true by varying the dimensions of the feature space and identifying the number of clusters required for generalization (Fig.
+A10B for identification accuracy in each condition).
 Here, we defined the number of clusters required for generalization as the point at which the identification accuracy of OOD samples exceeds $50$%.
-The relationship between the dimension of the feature space and the number of clusters necessary for generalization appears to be linear (Fig 9D).
+The relationship between the dimension of the feature space and the number of clusters necessary for generalization appears to be linear (Fig.
+9D).
 This finding suggests that achieving generalization does not necessarily require an exponentially large diversity that fills the entire feature space.
@@
 Although obtaining a large amount of brain data is hard work, it is important for a dataset to contain sufficiently diverse stimuli covering the effective dimensions of the target feature space to achieve zero-shot prediction.
-We also confirmed this phenomenon with a simple and transparent example ($D=2$, Fig 9E).
+We also confirmed this phenomenon with a simple and transparent example ($D=2$, Fig.
+9E).
 The training data covers sufficient axes in the target feature space, enabling the prediction of locations not present in the training set.
-Based on this low-dimensional intuition, we argue that successful zero-shot prediction requires that training data leads to representations that can serve as a basis for spanning the target feature space.
+Based on this low-dimensional intuition, we argue that successful zero-shot prediction requires training data to leads representations that can serve as a basis for spanning the target feature space.
 Leveraging such bases effectively enables the model to predict novel samples by predicting each basis and combining them.
-This compositional representation, spanning the target feature space, is crucial for zero-shot prediction [62] and reconstructing arbitrary visual images from limited brain data.
-Caveat with evaluation by pairwise identification
-Pairwise identification has been a standard metric for evaluating latent feature decoding [19] or reconstruction performance [39, 29, 38, 34, 36, 40, 41, 44, 45, 46, 57, 58].
-However, our analysis revealed that even with difficulties in accurately identifying semantic clusters the test samples belong to, the pairwise identification performance still surpassed its chance level (Fig 6F).
+This compositional representation, spanning the target feature space, is crucial for zero-shot prediction (Schug et al., 2024) and reconstructing arbitrary visual images from limited brain data.
+While our simulations varied the number and spread of clusters to examine the role of training diversity, we assumed that cluster centers were uniformly distributed across the latent feature space.
+However, this assumption may not accurately reflect the characteristics of actual datasets such as NSD, where cluster centers themselves can potentially be biased in the entire visual space (i.e., focusing on natural scenes in MSCOCO).
+Even with a large number of clusters, generalization to unseen images remains challenging if they occupy only a limited region of the visual space.
+These considerations emphasize that true diversity requires careful control over both the number and the spatial distribution of training clusters or images.
+2.2.3 Caveat with evaluation by pairwise identification
+Pairwise identification has been a standard metric for evaluating latent feature decoding (Horikawa and Kamitani, 2017) or reconstruction performance (Beliy et al., 2019; Shen et al., 2019a; Mozafari et al., 2020; Qiao et al., 2020; Ren et al., 2021; Gaziv et al., 2022; Takagi and Nishimoto, 2023a; Ozcelik and VanRullen, 2023; Scotti et al., 2023; Denk et al., 2023; Koide-Majima et al., 2024).
+However, our analysis revealed that even with difficulties in accurately identifying specific semantic clusters the test samples belong to, the pairwise identification performance still surpassed its chance level (Fig.
+6F).
+This result highlights a fundamental limitation of pairwise identification: its susceptibility to overestimation when the dataset or target features contain a strong high-level categorical structure.
 Here we critically examine this metric and demonstrate that significant results can be easily obtained when the target or predicted features exhibit certain structures.
 Pairwise identification is calculated as the accuracy with which the predicted features (either the output of a translator or features extracted from the output of a generator) can correctly identify the corresponding true ones, in pairs consisting of a true sample and one of the remaining samples in the test set.
+We refer to the latter remaining sample as a candidate sample in the following.
 If the candidate sample belongs to the same category as the true sample, the identification is expected to be difficult.
-Conversely, if the candidate sample belongs to a different category from the true one, identification becomes easier.
-Fig 10: Expected pairwise sample identification performance in categorically structured data.
+Conversely, if the candidate sample belongs to a different category from the true sample, identification becomes easier.
+This characteristic makes the metric highly dependent on categorical distinctions rather than the actual quality of feature prediction or reconstruction.
+Fig.
+10: Expected pairwise sample identification performance in categorically structured data.
 The $x$-axis represents the number of categories in the test set.
@@
 Samples from different categories can be distinguished with a pairwise identification accuracy of $q$, where q varies between $0.5$ and $1$.
-In this model, we assume the test set comprises $k$ categories with test samples equally distributed across each category for simplicity.
+Here, we assume the test set comprises $k$ categories with test samples equally distributed across each category for simplicity.
 We model the situation mentioned above by setting the expected identification accuracy at the chance level (i.e., $0.5$) when a candidate sample belongs to the same category as the true one.
-Conversely, when the candidate sample belongs to a different category than the true sample, we set the expected identification accuracy to a parameter $q$, ranging from 0.5 to 1.
+Conversely, when the candidate sample belongs to a different category from the true sample, we set the expected identification accuracy to a parameter $q$, ranging from $0.5$ to $1$.
 This parameter $q$ reflects the ease of identification across categories.
@@
 $\mathrm{Acc}=\frac{1}{k}\cdot 0.5+\left(1-\frac{1}{k}\right)\cdot q$ (4)
-(See Materials and Methods “Expected identification accuracy in imprecise reconstructions” for the derivation).
-Fig 10 illustrates the relationship between pairwise identification accuracy and the number of categories in the test set through line plots of expected values.
+(see Methods “Expected identification accuracy in imprecise reconstructions” for the derivation).
+Fig.
+10 illustrates the relationship between pairwise identification accuracy and the number of categories in the test set through line plots of expected values.
 Inset figures depict the underlying similarity structure of the test set.
-Notably, even when identification within categories fails completely and succeeds only between two categories, the pairwise identification accuracy can still reach a high value of up to $75$%.
-This finding underscores a significant limitation of the pairwise identification metric: evaluating reconstruction performance based solely on whether this metric is above the chance level may lead to misleading conclusions.
-Preserved image information across hierarchical DNN layers
+Notably, even when identification within categories fails completely and succeeds only between two categories, pairwise identification accuracy can still reach a high value of up to $75$%.
+This highlights a major limitation: above-chance performance may simply reflect an ability to differentiate broad categories rather than accurately reconstruct crucial visual details.
+Indeed, the same pattern emerges even under a hold-out split when the samples in the held-out cluster can still be categorized broadly.
+Consequently, performance evaluation only relying on this metric can lead to misleading conclusions about the model’s actual capabilities and generalizability.
+2.2.4 Preserved image information across hierarchical DNN layers
 The reconstruction of arbitrary visual images requires compositional latent features that can be effectively mapped into the image space.
-Hierarchical DNN features are often employed as such representations.
-However, do these features truly map to the image space?
-A common narrative suggests that the hierarchical processing discard pixel-level information through progressively expanding receptive fields.
-This view, however, is not entirely accurate.
-Fig 11: Image recovery check for hierarchical latent features.
+As our case study has suggested, hierarchical DNN features from VGG19 have been found to be suitable for zero-shot prediction or reconstruction tasks due to their compositional representations.
+At the same time, however, the extent to which these features truly map to the image space remains unclear.
+Indeed, a common narrative suggests that the hierarchical processing discards pixel-level information through progressively expanding receptive fields.
+Yet, this view is not entirely accurate.
+Fig.
+11: Image recovery check for hierarchical latent features.
 The images show the ability to recover visual inputs from different layers of the VGG19 network.
 For each of the three sample images, three different optimization methods were applied.
-The left column shows results from pixel optimization, which directly optimizes pixel values to minimize the loss between the generated image features and the target image features [65].
-The middle column displays optimization with a weak image prior, simultaneously optimizing both the weight parameters of the image prior models and their latent features [68].
-The right column presents optimization with a pre-trained image prior, which optimizes the latent features of a parameter-fixed image generator model to minimize the loss between the output image features and the target image features [69].
+The left column shows results from pixel optimization, which directly optimizes pixel values to minimize the loss between the generated image features and the target image features (Mahendran and Vedaldi, 2015).
+The middle column displays optimization with a weak image prior, simultaneously optimizing both the weight parameters of the image prior models and their latent features (Ulyanov et al., 2018).
+The right column presents optimization with a pre-trained image prior, which optimizes the latent features of a parameter-fixed image generator model to minimize the loss between the output image features and the target image features (Dosovitskiy and Brox, 2016).
 Rows correspond to different layers of the VGG19 network, progressing from earlier layers at the top to later layers at the bottom.
@@
 Reasonable image recovery is possible even from relatively high-level layers of the VGG19 network, challenging the notion that higher layers in deep neural networks discard image-level visual information.
-For example, the latent features of auto-encoder models [63, 51, 64] can represent images in a low-dimensional space while preserving their reversibility, which is reasonable considering that the model’s output is trained to match the input.
-Furthermore, Mahendran and Vedaldi [65] showed that input images can be reconstructed with reasonable accuracy even from relatively high-level layers of a DNN designed for an object recognition task.
-It has also been argued in the neuroscience field that large receptive field sizes do not necessarily impair neural coding capacity as long as the number and density of units remain constant [66, 67].
+For example, the latent features of auto-encoder models (Hinton and Salakhutdinov, 2006; Kingma and Welling, 2014; van den Oord et al., 2017) can represent images in a low-dimensional space while preserving their reversibility, which is reasonable considering that the model’s output is trained to match the input.
+Furthermore, Mahendran and Vedaldi (2015) showed that input images can be recovered with reasonable accuracy even from relatively high-level layers of a DNN designed for an object recognition task.
+It has also been argued in the neuroscience field that large receptive field sizes do not necessarily impair neural coding capacity as long as the number and density of units remain constant (Zhang and Sejnowski, 1999; Majima et al., 2017).
 These results challenge the notion that higher-level layers in DNNs discard all pixel-level information.
-To further illustrate this point, we performed a recovery check on each intermediate layer of the VGG19 network used in the iCNN methods (Fig 11; see also Fig 7).
-Given the DNN features of a target image, we optimized input pixel values to make the image’s latent features similar to the targets (See Materials and Methods “Recovery check of a single layer by iCNN”).
+To further illustrate this point, we performed a recovery check on each intermediate layer of the VGG19 network used in the iCNN methods (Fig.
+11; see also Fig.
+7).
+Given the DNN features of a target image, we optimized input pixel values to make the image’s latent features similar to the targets (see Methods “Recovery check of a single layer by iCNN”).
 We observed that input images can be recovered with reasonable accuracy from relatively high-level layers (around the 11th layer out of the total 19th layers).
 Furthermore, by introducing image generator networks to add constraints on image statistics, reasonable recovery can be achieved from even higher layers.
-By utilizing a weak image prior [68], which contains only information about the structure of images without any prior information on natural images, input images can be recovered from the 13th layer.
-When using an image generator that has learned natural image information [29, 69], input images can be recovered even from the 15th layer.
+By utilizing a weak image prior (Ulyanov et al., 2018), which contains only information about the structure of images without any prior information on natural images, input images can be recovered from the 13th layer.
+When using an image generator that has learned natural image information (Shen et al., 2019b; Dosovitskiy and Brox, 2016), input images can be recovered even from the 15th layer.
 These observations suggest that, even when feature representations shift from lower to higher levels through hierarchical processing, pixel-level information is not largely discarded; rather, much of the input information is preserved across almost the entire level.
@@
 Conversely, utilizing high-level image features, such as the output of DNNs, or features from other modalities, such as text annotations, is not a rational choice for visual image reconstruction tasks.
-These latent features make it challenging to recover the corresponding input image (Figure 7B) and are insufficient as surrogate representations of perceived visual images.
+These latent features make it challenging to recover the corresponding input image (Fig.
+7B) and are insufficient as surrogate representations of perceived visual images.
 Although recent text-to-image models and predicted text features can easily generate images, the outputs should not be interpreted as reconstruction results.
 Instead, it is more appropriate to view them as visualizations of decoded semantic information.
-Such visualizations are valuable for illustrating purposes.
-However, it is crucial to recognize the significant distinction between semantic visualization and reconstruction.
-How are we fooled by hallucinations of generative AIs?
-Generative AIs have recently made remarkable progress, with models now capable of producing high-resolution and realistic images from text input [23] or generating text of a quality indistinguishable from human-written content [24].
+While such visualizations are valuable for illustrating purposes, it is crucial to recognize the significant distinction between semantic visualization and reconstruction.
+2.2.5 How are we fooled by hallucinations of generative AIs?
+Generative AIs have recently made remarkable progress, with models now capable of producing high-resolution and realistic images from text input (Ramesh et al., 2021) or generating text of a quality indistinguishable from human-written content (Brown et al., 2020).
 However, due to the complex internal structure of these models and the vast amounts of data they are trained on, we are often fooled by the outputs of generative AIs.
@@
 This is likely a result of generative AIs being trained on a large amount of data and producing highly coherent and contextual responses.
-We may also believe these models are unbiased and can represent all possible data points, even though they inherently contain biases from their training data and developers [70].
-As we have observed, the generative AI-based reconstruction methods exhibit realistic appearances but poor generalizability (Fig 2).
+We may also believe these models are unbiased and can represent all possible data points, even though they inherently contain biases from their training data and developers (Messeri and Crockett, 2024).
+As we have observed, the generative AI-based reconstruction methods exhibit realistic appearances but poor generalizability (Fig.
+2).
 Could similar issues occur in visual image reconstruction studies as well?
@@
 Traditionally, we have held two beliefs: (1) generating realistic images from brain activity is challenging, and (2) if the reconstruction pipeline effectively captures the brain’s representation under natural image perception, the model’s output should also appear realistic.
-Based on these beliefs, realistic reconstruction is often considered as an indication of accurately reflecting the actual visual experience.
-Fig 12: The illustration of how realistic appearances can be misleading in visual reconstruction.
+Based on these beliefs, realistic reconstruction is often considered an indication of accurately reflecting the actual visual experience.
+Fig.
+12: The illustration of how realistic appearances can be misleading in visual reconstruction.
 On the left, Venn diagrams depict the relationship between two events: $R$, where the output has realistic appearances, and $T$, where the output truthfully reflects the visual experience.
@@
 Consequently, it becomes invalid to infer that the visual images are accurately reflected in the generator outputs solely because they appear realistic.
-Rather, as shown in Fig 12, the probability $\Pr[T|R]$ may become smaller as the generative AIs produce more convincing outputs (see also Fig 7).
+Rather, as shown in Fig.
+12, the probability $\Pr[T|R]$ may become smaller as the generative AIs produce more convincing outputs (see also Fig.
+7).
 This perspective emphasizes the need for careful evaluation of reconstruction performance, considering the possibility of hallucinations by generators.
 While pursuing realistic reconstructions to improve reconstruction fidelity is undoubtedly important, it would be counterproductive to obsess over naturalistic appearance to the point of neglecting the original goal of reconstructing perceived visual images.
-Discussion
+3 Discussion
 In this study, we critically examined generative AI-based visual image reconstruction methods to assess their true capabilities and limitations.
@@
 Our study emphasizes the need for more rigorous evaluation and careful interpretation of results in visual image reconstruction research, particularly when using generative AI-based methods.
-While our study critically examined the limitations of recent text-guided, diffusion-based methods for visual image reconstruction, it is important to acknowledge that these approaches provide new and potentially promising directions for brain decoding research.
-Although we emphasized the importance of achieving zero-shot prediction, it is crucial to recognize that most brain decoding studies focus on classification tasks, which, while not zero-shot, have provided insights into neural representations [1, 2].
-Moreover, the visualization of semantic contents (e.g., the supplementary movies in Horikawa et al.
-[4]), can have significant utility in visual communication, even if it does not constitute zero-shot prediction.
-It is also worth noting that the individual components of these text-guided, diffusion-based methods are already being utilized in various brain decoding applications.
-For instance, text latent features derived from deep neural networks and large language models (LLMs) have shown promise in analyzing semantic information from brain activity [71, 72, 73].
-Furthermore, diffusion models are not limited to text-to-image generation; they can be trained to generate images only from visual latent features, as Cheng et al.
-[32] demonstrated in successful reconstruction of subjective experiences.
-By leveraging these components and exploring their potential synergies, researchers can continue to push the boundaries of brain decoding and visual image reconstruction, while keeping in mind the challenges and limitations highlighted in our study.
+While our study highlights the limitations of text-guided diffusion models for visual image reconstruction, it is important to acknowledge that these methods offer promising directions for brain decoding research.
+For instance, they can generate “ROI optimal stimuli,” which create images that activate a certain ROI activities maximally while not activating other ROIs activities, through the learned mapping between brain activities and latent features (Ozcelik et al., 2022; Ozcelik and VanRullen, 2023).
+Although current studies focus on well-known functional ROIs (e.g., face, word, place, and body regions), this approach can extend to less understood brain regions.
+By generating optimal stimuli and identifying robust visual patterns, we can formulate new hypotheses about functional representations in previously uncharacterized brain areas.
+This data-driven approach complements traditional hypothesis-driven methods, potentially uncovering novel functional regions overlooked by conventional analyses.
+Moreover, while we emphasized the importance of zero-shot prediction, it is crucial to recognize that most brain decoding studies focus on classification tasks.
+Although these tasks are not zero-shot, they have nonetheless yielded valuable insights into neural representations (Haxby et al., 2001; Kamitani and Tong, 2005).
+In that sense, the text-guided or diffusion-based methods can also be utilized as tools for the visualization of decoded semantic contents (e.g., the supplementary movies in Horikawa et al., 2013).
+Such visualization can be highly useful for visually conveying decoded information, even if it does not constitute zero-shot prediction.
+Additionally, it is worth noting that the individual components of text-guided reconstruction methods have already been utilized in various brain decoding applications.
+For instance, text latent features from deep neural networks and large language models (LLMs) have shown promise in analyzing semantic information from brain activity (Tang et al., 2023; Caucheteux et al., 2023; Zhou et al., 2024).
+Furthermore, diffusion models can extend beyond text-to-image generation as MindEye2 generates images from visual latent spaces.
+Notably, Cheng et al.
+(2023) successfully reconstructed subjective experiences using a diffusion model within a carefully designed experiment.
+Leveraging these components and exploring their potential synergies, researchers can advance brain decoding and visual image reconstruction while addressing the challenges highlighted in our study.
 The recent trend of collecting and sharing large-scale visual neural datasets, such as those by Hebart et al.
-[14] and Xu et al.
-[15], is a welcome development in the field of neuroscience.
+(2023) and Xu et al.
+(2024), is a welcome development in the field of neuroscience.
 These datasets provide valuable resources for researchers to investigate brain function and advance our understanding of visual processing.
-The NSD is a particularly notable example, as it was created with the goal of extensively sampling brain responses to a wide range of natural visual stimuli [12, 13].
-The NSD has been widely utilized in various studies [74, 75, 76], demonstrating its value to the research community.
+The NSD is a particularly notable example, as it was created with the goal of extensively sampling brain responses to a wide range of natural visual stimuli (Allen et al., 2022; Naselaris et al., 2021).
+The NSD has been widely utilized in various studies (Prince et al., 2022; Gifford et al., 2023; Conwell et al., 2024), demonstrating its value to the research community.
 While our results suggest that the semantic and visual diversity of the NSD stimuli may not be as high as initially thought, and there is substantial overlap between the training and test sets provided by the NSD authors, this does not diminish the overall importance and usefulness of the dataset.
 However, to fully leverage the NSD and other publicly available large-scale datasets for developing generalizable and zero-shot prediction models, it is crucial to consider the data split between training and test sets carefully.
-When aiming for generalizable predictions, such as in visual image reconstruction, researchers should verify whether significantly similar stimuli are included in both the training and test sets.
-If necessary, redesigning the training and test split can ensure that the models are tested on their ability to predict novel stimuli.
-Moreover, recent advancements in functional alignment and inter-site neural code conversion methods [77, 78, 79] hold promise for combining datasets from different sources, enabling truly larger-scale data analysis in neuroscience.
+While many large-scale datasets provide designated training and test splits, these splits are often not optimally designed to evaluate zero-shot prediction performance.
+When aiming for generalizable predictions beyond training examples as in visual image reconstruction, researchers should carefully verify whether significantly similar stimuli are included in both the training and test sets (Fig.
+A5).
+If significant overlap exists, redesigning the training-test split becomes necessary to ensure the test set contains stimuli substantially dissimilar from those in the training set, thereby enabling genuine evaluation of generalization capabilities (Fig.
+A7).
+Moreover, recent advancements in functional alignment and inter-site neural code conversion methods (Haxby et al., 2011; Yamada et al., 2015; Wang et al., 2024) hold promise for combining datasets from different sources, enabling truly larger-scale data analysis in neuroscience.
 These techniques allow researchers to align brain activity patterns across individuals and measurement sites even when stimuli are not shared across datasets.
 By leveraging these methods, researchers can pool data from various sources, increasing the sample size and diversity of the combined dataset, mitigating the limitations of individual datasets, and enhancing the development of generalizable and zero-shot prediction models.
-Investigating neural responses to natural stimuli is a highly valuable approach to understanding brain function and representation [80, 81].
+Investigating neural responses to natural stimuli is a highly valuable approach to understanding brain function and representation (Nastase et al., 2020; Hasson et al., 2020).
 As our brains have developed while being exposed to natural scenes, it is crucial to use natural stimuli, especially in model training.
@@
 With the increasing scale of neural data and the growing complexity of analysis pipelines, there is a risk that the learned mappings may produce unexpected shortcuts, just as we have demonstrated that the text-guided reconstruction methods exploited the semantic and visual overlap between training and test sets.
-In the field of comparative and developmental psychology, researchers often prioritize using not natural but simple stimuli for better experimental controls and more precise inferences about infants’ cognitive abilities [82, 83].
+In the field of comparative and developmental psychology, researchers often prioritize using not natural but simple stimuli for better experimental controls and more precise inferences about infants’ cognitive abilities (Kominsky et al., 2022; Frank, 2023).
 Drawing inspiration from this approach, we argue that the evaluation of visual image reconstruction should not be limited to complex natural stimuli alone.
@@
 Kavasidis et al.
-[84] collected a dataset of EEG signals recorded during natural image perception.
-Their visual stimuli consisted of $2,000$ images selected from $40$ object categories in ImageNet ($50$ images per category) and the test set contained the same categories that are included in the training set (See also [85, 15] for other issues with the data set).
+(2017) collected a dataset of EEG signals recorded during natural image perception.
+Their visual stimuli consisted of $2,000$ images selected from $40$ object categories in ImageNet ($50$ images per category) and the test set contained the same categories that are included in the training set (see also Li et al.
+(2018) and Xu et al.
+(2024) for other issues with the data set).
 Denk et al.
-[57] attempted to develop music reconstruction methods from fMRI activity patterns.
-Their music stimuli consisted of $540$ music pieces selected from $10$ music genres, and the test set contained the same genres as the training set [86].
+(2023) attempted to develop music reconstruction methods from fMRI activity patterns.
+Their music stimuli consisted of $540$ music pieces selected from $10$ music genres, and the test set contained the same genres as the training set (Nakai et al., 2021).
 Orima et al.
-[87] attempted reconstructing perceived texture images from EEG signals.
+(2024) attempted to reconstruct perceived texture images from EEG signals.
 Their texture stimuli consisted of $191$ image patches extracted from $21$ natural textures, and they performed a reconstruction analysis in a leave-one-out manner.
-It should be carefully examined whether these studies may suffer from output dimension collapse, merely decoding the information from trained categories from brain activity.
-Experimental design for train-test stimuli setups requires careful consideration.
+It should be carefully examined whether these studies may suffer from output dimension collapse, merely decoding the broad category-level information observed in the training set.
+Experimental design for training–test stimulus setups requires careful consideration.
 Dado et al.
-[88] conducted an image reconstruction analysis from the multi-unit activity of a macaque using images generated from latent features of generative models.
-Importantly, all of the test stimuli were generated from the averaged latent features of the categories used in their training phase.
+(2024) conducted an image reconstruction analysis from the multi-unit activity of a macaque using images generated from latent features of generative models.
+Importantly, all of the test stimuli were generated from the averaged latent features of the categories used in their training phase, suggesting the test stimuli are highly biased to the training set (see also Fig.
+A5).
 While the authors addressed potential biases by redesigning the training and test split, researchers should exercise prudence when utilizing the dataset.
 Our inspection of the movie stimuli from Nishimoto et al.
-[89] revealed that many frames in the test movie stimuli were nearly identical to those in the training set (S5 Fig).
+(2011) revealed that many frames in the test movie stimuli were nearly identical to those in the training set (Fig.
+A11).
 This similarity likely results from temporally adjacent video frames being split between the training and test stimuli.
-While our preliminary analysis of their latent features (motion energy features) did not show unusual clustering (note also that the study presents a way of retrieving movie instances via a brain encoding model, rather than reconstruction in the current sense), caution is required when using the stimulus set to extract other types of features [90, 91].
+While our preliminary analysis of their latent features (motion energy features) did not show unusual clustering (note also that the study presents a way of retrieving movie instances via a brain encoding model, rather than reconstruction in the current sense), caution is required when using the stimulus set to extract other types of features (Huth et al., 2012, 2016).
 Claims of predicting arbitrary instances or achieving zero-shot prediction warrant thorough scrutiny.
 There is a widespread practice of using test data for fine-tuning, which can also be a questionable procedure.
-Some studies have proposed methods that involve fine-tuning models using the entire test brain data, but not the test stimuli, following initial training with the training stimuli and brain data [39, 43].
+Some studies have proposed methods that involve fine-tuning models using the entire test brain data, but not the test stimuli, following initial training with the training stimuli and brain data (Beliy et al., 2019; Chen et al., 2023).
 While the absence of test labels (stimuli) during fine-tuning may help avoid obvious overfitting, these approaches still treat test brain data as previously observed information, rather than as a proxy for novel data potentially encountered in real-world situations.
@@
 Although such procedures may effectively improve performance on existing benchmarks or competitions, it is crucial to recognize that incorporating test data information during any stage of training can undermine the validity of neuroscientific claims and limit the real-world applicability of the methods.
-The issue of double dipping, which refers to the use of the same dataset for both data/variable selection and selective analysis (inference and prediction), has been widely recognized in neuroscience [92, 93].
+The issue of double dipping, which refers to the use of the same dataset for both data/variable selection and selective analysis (inference and prediction), has been widely recognized in neuroscience (Kriegeskorte et al., 2009; Button, 2019).
 In classification tasks, while selecting input variables using test data is problematic, using the same output labels (target variables) for training and test sets is not inherently flawed, given that the nature of classification assumes consistent categories across datasets.
@@
 Additionally, the independence of test stimuli from the model training processes requires careful examination, particularly when using pre-trained deep neural network (DNN) models and foundation models like CLIP or diffusion models.
-These models are typically trained on vast amounts of data available on the internet [22, 24, 50], which likely includes public datasets such as MS-COCO [94] that are used in the NSD.
+These models are typically trained on vast amounts of data available on the internet (Radford et al., 2021; Brown et al., 2020; Rombach et al., 2022), which likely includes public datasets such as MS-COCO (Lin et al., 2014) that are used in the NSD.
 This overlap raises potential concerns about the true independence of test stimuli, as we cannot rule out the possibility that these pre-trained models have acquired representations specifically tailored to the stimuli used for the model training.
 To address these issues and ensure a more rigorous evaluation of model generalization capabilities, researchers may consider using test stimuli that are not publicly available on the internet.
-This could include self-created stimuli or carefully curated datasets that have not been used in the training of widely used AI models [29, 32].
+This could include self-created stimuli or carefully curated datasets that have not been used in the training of widely used AI models (Shen et al., 2019b; Cheng et al., 2023).
 Such an approach would provide a more stringent test of a model’s ability to generalize to truly novel inputs.
+Although we highlighted how stimulus overlap and limited diversity can lead to an overestimation of visual image reconstruction performance, such considerations do not apply to all reconstruction tasks.
+The dimensionality of the target space varies depending on the domain we seek to reconstruct; for instance, movement reconstruction often involves low-dimensional outputs (e.g., movement direction or velocity), where a small number of brain activity samples may suffice.
+In such cases, robust reconstruction is achievable by memorizing the brain-target pairs that cover the output space and classifying the test brain data into trained targets or interpolating between them.
+However, in visual image reconstruction with high-dimensional output space, evaluating zero-shot prediction is essential.
+Models must demonstrate true generalization to unseen stimuli rather than reflecting training data biases for establishing both reliability and practical utility.
 One of the remaining challenges in visual image reconstruction is the development of metrics for evaluating the quality and accuracy of the reconstructed images.
@@
 Following this, quantitative metrics should be employed for a more objective, high-throughput evaluation.
-However, as our analysis has suggested, it can be misleading to evaluate reconstruction by heavily relying on identification performance based on the relative similarity among alternatives [58].
+However, as our analysis has suggested, it can be misleading to evaluate reconstruction by heavily relying on identification performance based on the relative similarity among alternatives (Koide-Majima et al., 2024).
 Even in cases where the reconstructed images only capture superficial information, such as categories or overall brightness, identification metrics can still be high.
@@
 It is crucial to develop more appropriate similarity metrics that can accurately measure the perceptual similarity between the reconstructed and original images.
-One promising approach is to leverage image quality assessment (IQA) techniques from the computer vision field [55, 95].
+One promising approach is to leverage image quality assessment (IQA) techniques from the computer vision field (Fu et al., 2024; Ding et al., 2020).
 These techniques are designed to quantify the perceptual quality of images and can be adapted to the specific requirements of visual image reconstruction.
-Our findings have significant implications for research integrity and responsible dissemination of scientific results.
-Visual image reconstruction methods have gained attention not only from neuroscientists but also from the general public and policymakers, sparking discussions about their potential applications and risks [8].
+Our findings have profound implications for research integrity and responsible dissemination of scientific results.
+Visual image reconstruction methods have gained attention not only from neuroscientists but also from the general public and policymakers, sparking discussions about their potential applications and risks (UNESCO, 2023).
 These stakeholders often contemplate the possibilities of seamless information communication through the brain, such as in brain-machine interfaces (BMIs), or the dangers of unauthorized access to private information from brain activity.
@@
 Most reconstruction methods analyze previously acquired brain data offline.
-The brain data used for reconstructing images are often averaged over multiple presentations of the test image, with only a few studies demonstrating single-trial reconstruction results [28, 32].
-Further, it has been argued that subject cooperation is essential for reliably training and testing decoding models [71].
+The brain data used for reconstructing images are often averaged over multiple presentations of the test image, with only a few studies demonstrating single-trial reconstruction results (Miyawaki et al., 2008; Cheng et al., 2023).
+Further, it has been argued that subject cooperation is essential for reliably training and testing decoding models (Tang et al., 2023).
 Given these realities, public expectations often exceed current capabilities, and meeting these high demands in the short term is challenging.
@@
 In light of these challenges, while the field of visual image reconstruction from brain activity holds great promise, it is our responsibility as researchers to ensure that its current capabilities and limitations are accurately communicated to all stakeholders.
-Materials and methods
+3.1 Recommendations
+Finally, we present several guidelines for critically testing visual image reconstruction methods.
+These suggestions build upon the limitations and challenges identified in our study and provide a pathway for future improvements in reconstruction research.
+3.1.1 Stimulus design and data splits
+Expand and control diversity.
+Collect or curate training datasets that span sufficient axes in the feature space so that new (unseen) visual images can be predicted.
+When possible, include artificial or carefully-designed stimuli as well as natural stimuli in the test set to provide clearer interpretability and control.
+Natural image reconstruction is not necessarily the ultimate goal, given that humans perceive both natural and artificial images.
+Avoid overlaps in training and test stimuli.
+To evaluate true zero-shot capacity, ensure that test images do not overlap semantically or visually with training images.
+Identifying and removing near-duplicates or highly similar images from the test set helps prevent hidden “shortcut” solutions in which the model simply memorizes or classifies into known stimuli.
+Use multiple and independent test sets.
+Consider separate test sets with varying complexity—e.g., natural images, artificial shapes, and out-of-distribution samples—to comprehensively assess generalizability.
+Disentangling performance across these varied sets can reveal whether a method is genuinely reconstructing novel content or only handling a narrow range of stimuli.
+3.1.2 Model specification and latent feature choice
+Confirm the generator’s recovery capability.
+Perform “recovery checks” by feeding true latent features (extracted directly from the original images) into the image generator.
+If the generator fails to reproduce the original images faithfully, it cannot serve as a valid reconstruction module.
+This step clarifies whether errors in the test phase stem from the latent translator or from a generator prone to hallucinations.
+Use compositional, image-preserving features.
+Favor latent features (such as mid-level DNN layers) that retain sufficient image-level detail and compositional representations.
+This ensures that, with perfect translation from the brain, the original image can be reconstructed accurately.
+In contrast, purely semantic or text-based features often discard important visual details, limiting reconstruction fidelity.
+Mitigate output dimension collapse.
+Choose or design translators (e.g., modular or sparse voxel selection approaches) to reduce collapse onto limited training-set clusters.
+Avoid overfitting to narrow categories by ensuring that the model maintains high-dimensional predictive capacity across all important feature dimensions.
+3.1.3 Evaluation metrics and result transparency
+Prioritize perceptual resemblance across diverse targets.
+Before focusing on quantitative metrics, visually confirm that reconstructions capture the perceptual features of all target stimuli.
+Testing fidelity on diverse or out-of-distribution samples is crucial for confirming genuine reconstruction rather than mere classification or retrieval.
+Avoid overemphasizing photorealism, as it can mask inaccuracies.
+Include extensive examples so readers can visually assess quality and variation.
+Avoid cherry-picking and post-hoc selection.
+Generative models can produce multiple plausible outputs from a single latent feature.
+Selecting only the best-looking or most accurate images artificially inflates performance estimates.
+Present results transparently (e.g., showing random draws or evaluating robustness across different seeds) to offer a fair depiction of each model’s true reliability.
+Use robust metrics beyond pairwise identification.
+Pairwise identification is easy to implement, but can overestimate performance especially in categorically structured data.
+Carefully design the selection of candidates when conducting identification analysis.
+In addition, supplement it with more stringent evaluation, such as fine-grained semantic checks, and advanced image-quality metrics (e.g., SSIM, DreamSim, other learned perceptual measures).
+Distinguish clearly between high-level semantic alignment and perceptual similarity.
+3.1.4 Collaboration and ethical communication
+Interdisciplinary collaborations.
+Close collaboration among neuroscientists, machine learning researchers, and cognitive scientists is vital for designing robust experiments, interpreting results correctly, and addressing complex technical pitfalls (e.g., data leakage, improper splits, or hallucinations by diffusion models).
+Transparent reporting and data-sharing.
+Provide open-source code, clearly document training–test splits, and release relevant stimuli annotations to enable reproducibility.
+Transparency fosters collective progress, allowing others to replicate or extend your findings under more controlled or diverse conditions.
+Realistic public and policy discourse.
+Communicate clearly that current reconstruction methods do not equate to unconstrained “mind reading” and often depend on carefully curated data.
+Highlight the role of subject cooperation, offline averaging, and limited generalizability so that stakeholders—such as policymakers, journalists, and the public—avoid overestimating immediate real-world capabilities.
+In sum, authentic visual image reconstruction from brain activity requires careful management of dataset diversity and overlap, prudent model specification (especially in latent feature selection), and rigorous evaluation metrics beyond simple identification.
+By adhering to these recommendations, researchers can reduce the risk of reporting spurious reconstructions, bringing us closer to methods that genuinely reflect an individual’s perceptual experience.
+4 Conclusions
+Our critical analysis of recent generative AI-based visual image reconstruction methods revealed several limitations and challenges.
+We demonstrated that the apparent success of text-guided reconstruction methods primarily stems from a combination of classification into trained categories and hallucination through text-to-image diffusion, rather than genuine zero-shot reconstruction of novel images, which was the original goal of reconstruction studies.
+Our formal analysis revealed that predicting features with limited diversity can lead to output dimension collapse, where predictions become confined to patterns similar to the training set.
+Our simulation analysis demonstrated that successful zero-shot prediction requires training data with sufficient diversity to span the effective dimensions of the target feature space.
+We also pointed out that standard identification metrics can be misleading, especially when the target set has an underlying similarity structure.
+Additionally, we provided evidence that much of the input information is preserved at almost all hierarchical levels of deep neural networks.
+Finally, we pointed out that recent realistic reconstructions produced by generative AI models may appear convincing but do not necessarily reflect accurate representations of perceived visual experiences.
+These findings emphasize the need for more rigorous evaluation methods, diverse datasets, and careful interpretation of results in visual image reconstruction research.
+Future work should focus on prioritizing accurate reconstruction rather than naturalistic appearance.
+This objective would be achieved by utilizing compositional representations that can effectively span the feature space while maintaining the ability to recover original stimuli from latent features.
+As the field continues to attract attention from both researchers and the public, our results have important implications for research integrity and responsible development of neurotechnology, highlighting the need to balance scientific advancement with realistic expectations about current technological capabilities.
+Methods
 Datasets
-We utilized two datasets: the Natural Scene Dataset (NSD) [12] and the Deeprecon dataset [29].
+We utilized two datasets: the Natural Scene Dataset (NSD; Allen et al., 2022) and the Deeprecon dataset (Shen et al., 2019b).
 Both datasets comprise visual stimuli and corresponding fMRI activity collected when subjects perceived the stimuli.
-In the NSD dataset, eight subjects were presented with MSCOCO images [94], yielding $30,000$ brain activity samples per subject, which is three times the amount provided by the Deeprecon dataset.
-The Deeprecon dataset includes fMRI activity data from subjects presented with both ImageNet images [9] and artificial images.
+In the NSD dataset, eight subjects were presented with MSCOCO images (Lin et al., 2014), yielding $30,000$ brain activity samples per subject, which is three times the amount provided by the Deeprecon dataset.
+The Deeprecon dataset includes fMRI activity data from subjects presented with both ImageNet images (Deng et al., 2009) and artificial images.
 It contains roughly $8,000$ brain samples per subject.
@@
 The artificial images were only used as test data to check the generalizability performance of the proposed reconstruction methods.
-In both datasets, we adopted the train-test split used in previous studies and utilized data from the first subject (S1 in the NSD and Subject 1 in the Deeprecon).
+In both datasets, we adopted the training–test split used in previous studies and utilized data from the first subject (S1 in the NSD and Subject 1 in the Deeprecon).
 Text-guided reconstruction methods require text annotations of images.
@@
 Reconstruction methods
-We utilized three image reconstruction methods: StableDiffusionReconstruction [44], Brain-Diffuser [45], and iCNN [29].
-Each method employs two common steps: first, translating brain activity patterns into latent features of the stimuli, and second, generating images from these latent features using an image generator (see also Fig 1).
-In the StableDiffusionReconstruction method, the latent features are the VAE [51] features calculated from stimulus images and the CLIP text features [22] from the image annotations.
-The generator is StableDiffusion [50].
+We utilized three image reconstruction methods: StableDiffusionReconstruction (Takagi and Nishimoto, 2023a), Brain-Diffuser (Ozcelik and VanRullen, 2023), MindEye2 (Scotti et al., 2024), and iCNN (Shen et al., 2019b).
+Each method employs two common steps: first, translating brain activity patterns into latent features of the stimuli, and second, generating images from these latent features using an image generator (Fig.
+1).
+In the StableDiffusionReconstruction method, the latent features are the VAE (Kingma and Welling, 2014) features calculated from stimulus images and the CLIP text features (Radford et al., 2021) from the image annotations.
+The generator is StableDiffusion (Rombach et al., 2022).
 They first generate low-resolution images from the translated VAE features, and those images are further fed into the StableDiffusion model with translated text features to generate images.
 The generated images are regarded as reconstructed images from brain activity.
-In the Brain-Diffuser method, the latent features are the VDVAE [52], CLIP vision features from stimulus images and CLIP text features from the text annotations of the image.
-The generator is Versatile diffusion [49].
+In the Brain-Diffuser method, the latent features are the VDVAE (Child, 2021), CLIP vision features from stimulus images and CLIP text features from the text annotations of the image.
+The generator is Versatile diffusion (Xu et al., 2022).
 Similar to StableDiffusionReconstruction, low-resolution images are first generated from the translated VDVAE features, and these images are further used for the input of the versatile diffusion model with the translated vision and text features.
 The generated images are regarded as reconstructed images from brain activity.
-In the iCNN method, the latent features are the intermediate output of the VGG19 layer [56].
-As a generator, they used the pre-trained image generator [69], and they solved the optimization problem to minimize the discrepancy between the VGG19 features calculated from the generated images and the translated VGG19 features.
+In the MindEye2 methods, the latent features are the variant of CLIP vision model features (OpenCLIP ViT/bigG-14).
+The generator is multiple Stable diffusion XL (SDXL) models (Podell et al., 2023).
+They first generate images from translated OpenCLIP ViT/bigG-14 features by unCLIP technique (Ramesh et al., 2022; Scotti et al., 2024).
+They then generate the final reconstruction by SDXL, integrating the generated images and text caption predicted from the translated OpenCLIP ViT/bigG-14 features by GiT Image2Text modules (Wang et al., 2022).
+In the iCNN method, the latent features are the intermediate output of the VGG19 layer (Simonyan and Zisserman, 2015).
+As a generator, they used the pre-trained image generator (Dosovitskiy and Brox, 2016), and they solved the optimization problem to minimize the discrepancy between the VGG19 features calculated from the generated images and the translated VGG19 features.
 Well-optimized images are regarded as reconstructed images.
 UMAP visualization
-To investigate dataset diversity, we employed uniform manifold approximation and projection (UMAP), a non-linear dimensionality reduction technique [54], to learn a projection from a latent features space to a lower dimension (UMAP embedding space).
+To investigate dataset diversity, we employed uniform manifold approximation and projection (UMAP), a nonlinear dimensionality reduction technique (McInnes et al., 2018), to learn a projection from a latent features space to a lower dimension (UMAP embedding space).
 We used both the training and test CLIP text features to learn the UMAP projection.
 These features were combined and standardized beforehand.
-The UMAP hyperparameters followed the official UMAP guide for clustering usage (https://umap-learn.readthedocs.io/en/latest/clustering.html#umap-enhanced-clustering) with cosine distance as a distance metric (the visualization result with the default UMAP parameters is shown in S1 Fig.).
-The learned UMAP was also used to project the features predicted from brain activity (Fig 6B).
+The hyperparameters followed the official guide for clustering usage (https://umap-learn.readthedocs.io/en/latest/clustering.html) with cosine distance as a distance metric.
+The learned UMAP was also used to project the features predicted from brain activity (Fig.
+6B).
 After standardizing the predicted features using the same mean and standard deviation parameters used in UMAP projection learning, we projected the predicted features onto the UMAP embedding space.
 Simulation with clustered data
-We conducted a simulation analysis to illustrate output dimension collapse (Fig 8) and to examine the generalization performance beyond the training data (Fig 9).
+We conducted a simulation analysis to illustrate output dimension collapse (Fig.
+8) and to examine the generalization performance beyond the training data (Fig.
+9).
 These analyses involve a teacher-student learning task.
 The teacher model generated the pairs of target features and brain activity as training samples, and the student model learned a mapping from the training samples.
-To imitate the feature translation situation from brain activity, observation noise was added to the brain data.
 To imitate the feature translation situation from brain activity, observation noise was added to the brain data.
@@
 $\sigma^{2}_{\mathrm{noise}}$ is the scalar value representing the variance of observation noise.
-Using $N$ training samples, $X_{\mathrm{tr}}=[\mathbf{x}^{(1)},\ldots,\mathbf{x}^{(N)}]^{\top}$ and $Y_{\mathrm{tr}}=[\mathbf{y}^{(1)},\ldots,\mathbf{y}^{(N)}]^{\top}$, we trained the student model.
+For $N$ training samples, $X_{\mathrm{tr}}=[\mathbf{x}^{(1)},\ldots,\mathbf{x}^{(N)}]^{\top}$ and $Y_{\mathrm{tr}}=[\mathbf{y}^{(1)},\ldots,\mathbf{y}^{(N)}]^{\top}$, we trained the student model.
 by the ridge regression algorithm.
@@
 where $\lambda$ is the regularization parameter.
-After training the student weight $W$, we illustrated the phenomenon of output dimension collapse as in Fig 8 by predicting randomly generated data.
+After training the student weight $W$, we illustrated the phenomenon of output dimension collapse as in Fig.
+8 by predicting randomly generated data.
 First, we generated random test latent features from the Gaussian distribution with a mean of $\mathbf{0}$ and a variance equal to the training set variance scaled by a single scalar value ($9.0$).
@@
 We set $D=512$, $\sigma_{\mathrm{intra}}^{2}=10/512$, $\sigma_{\mathrm{inter}}^{2}=100/512$, and $\sigma_{\mathrm{noise}}^{2}=10$.
-We evaluated the zero-shot performance of the student model as in Fig 9.
+We evaluated the zero-shot performance of the student model as in Fig.
+9.
 We prepared two types of test samples: in-distribution test samples and out-of-distribution test samples.
@@
 For in-distribution test samples, we chose the cluster center randomly without replacement.
-We mainly focused on the typical cluster identification performance for dimension $D$, the number of clusters in the training set $C$, and the ratio of variance about the cluster structure $\sigma_{\mathrm{intra}}^{2}/\sigma_{\mathrm{inter}}^{2}$.
-We used a sufficiently large number of training samples $N=500,000$.
-This size was constant when the above parameters were varied.
+We mainly explored the dependency of the typical cluster identification performance on the following hyperparameters: dimension $D$, the number of clusters in the training set $C$, and the ratio of variance about the cluster structure $\sigma_{\mathrm{intra}}^{2}/\sigma_{\mathrm{inter}}^{2}$.
+We used a sufficiently large number of training samples $N=500,000$, and we kept $N$ constant while changing the above hyperparameters, especially $C$.
 Other parameters were also fixed in this simulation as follows: $\sigma_{\mathrm{intra}}^{2}+\sigma_{\mathrm{inter}}^{2}=110/D$, ${\bar{A}_{ij}}\sim\mathcal{N}(0,D^{-1/2})$, $\sigma_{\mathrm{noise}}^{2}=0.25$, $\lambda=1.0$, $n=100$ and $t=32$.
@@
 If we have a sufficiently large number of test samples, the above identification accuracy can be approximated as
-$\mathrm{Acc}$ $\underset{\phantom{n\to\infty}}{=}\frac{1}{n(n-1)}\left(\sum_{(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in Z}S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})\right)+\frac{1}{n(n-1)}\left(\sum_{(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in\bar{Z}}S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})\right)$ (13)
-$\underset{\phantom{n\to\infty}}{=}\frac{|Z|}{n(n-1)}\left(\frac{1}{|Z|}\sum_{(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in Z}S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})\right)+\frac{|\bar{Z}|}{n(n-1)}\left(\frac{1}{|\bar{Z}|}\sum_{(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in\bar{Z}}S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})\right)$ (14)
-=$\underset{n\to\infty}{=}\frac{1}{k}\cdot\mathbb{E}_{p(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-}\mid(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in Z)}[S]+\left(1-\frac{1}{k}\right)\cdot\mathbb{E}_{p(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-}\mid(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in\bar{Z})}[S]$ (15)
+$\mathrm{Acc}$ $\underset{\phantom{n\to\infty}}{=}\frac{1}{n(n-1)}\left(\sum_{(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in Z}S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})\right)$
+$\hskip 99.58464pt+\frac{1}{n(n-1)}\left(\sum_{(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in\bar{Z}}S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})\right)$ (13)
+$\underset{\phantom{n\to\infty}}{=}\frac{|Z|}{n(n-1)}\left(\frac{1}{|Z|}\sum_{(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in Z}S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})\right)$
+$\hskip 99.58464pt+\frac{|\bar{Z}|}{n(n-1)}\left(\frac{1}{|\bar{Z}|}\sum_{(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in\bar{Z}}S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})\right)$ (14)
+=$\underset{n\to\infty}{=}\frac{1}{k}\cdot\mathbb{E}_{p(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-}\mid(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in Z)}[S]$
+$\hskip 99.58464pt+\left(1-\frac{1}{k}\right)\cdot\mathbb{E}_{p(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-}\mid(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in\bar{Z})}[S]$ (15)
 $\underset{\phantom{n\to\infty}}{=}\frac{1}{k}\cdot 0.5+\left(1-\frac{1}{k}\right)\cdot q,$ (16)
@@
 Recovery check of a single layer by iCNN
-We performed a recovery check analysis using a single layer from the iCNN method in Fig 11.
-The iCNN method generates an image by optimizing pixel values to make the image’s latent features similar to the target latent features [29].
-In the pixel optimization condition (the left columns of each recovery image in Fig 11), we directly optimized the pixel values of images to minimize the mean squared loss between the latent features of the image as well as the total-variance (TV) loss of pixel values [65].
-Additionally, the iCNN method can incorporate image generator networks (middle and right columns of each recovery image in Fig 11) to add constraints on image statistics.
+We performed a recovery check analysis using a single layer from the iCNN method in Fig.
+11.
+The iCNN method generates an image by optimizing pixel values to make the image’s latent features similar to the target latent features (Shen et al., 2019b).
+In the pixel optimization condition (the left columns of each recovery image in Fig.
+11), we directly optimized the pixel values of images to minimize the mean squared loss between the latent features of the image as well as the total-variance (TV) loss of pixel values (Mahendran and Vedaldi, 2015).
+Additionally, the iCNN method can incorporate image generator networks (middle and right columns of each recovery image in Fig.
+11) to add constraints on image statistics.
 Instead of optimizing the pixel values, we optimized the parameters related to the generator networks to minimize the mean squared loss between the latent features obtained through the generator networks and the target latent features.
-As a weak image prior, we used Deep Image Prior (DIP) [68].
+As a weak image prior, we used Deep Image Prior (DIP; Ulyanov et al., 2018).
 DIP utilizes a hierarchical U-Net architecture as an inherent prior for image tasks, capturing the statistical regularities of images without relying on a specific dataset.
@@
 Subsequently, the latent features and weight parameters of DIP were optimized to minimize the difference between the network’s output and target DNN features.
-For the pre-trained image prior [69], we used the same generator model as in Shen et al.
-[29] optimizing the latent features of the pre-trained networks.
+For the pre-trained image prior (Dosovitskiy and Brox, 2016), we used the same generator model as in Shen et al.
+(2019b) optimizing the latent features of the pre-trained networks.
 Acknowledgments
-We thank our laboratory team, especially Eizaburo Doi, Haibao Wang, Kenya Otsuka, Hideki Izumi, and Matthias Mildenberger for their invaluable comments and suggestions on the manuscript.
-This work was supported by the Japan Society for the Promotion of Science (JSPS: KAKENHI grants JP20H05954, JP20H05705 to Y.K, JP21K17821 to Y.N, and 22KJ1801 to K.S.), Japan Science and Technology Agency (JST: CREST grants JPMJCR18A5, and JPMJCR22P3 to Y.K.), and New Energy and Industrial Technology Development Organization (NEDO: commissioned project, JPNP20006 to Y.K.).
-Supporting information
-S1 Fig.
-UMAP visualization of CLIP text features in the Natural Scene Dataset (NSD) using default parameters.
-Similar to Fig 4A in the main text, this scatter plot visualizes the distribution of semantic concepts within the NSD.
+We thank our laboratory team, especially Eizaburo Doi, Haibao Wang, Kenya Otsuka, Hideki Izumi, and Matthias Mildenberger, for their invaluable feedback and insightful suggestions on the manuscript.
+Appendix A Supplementary figures
+Fig.
+A1: Reconstructions from the NSD dataset using a sample size matched to the deeprecon dataset.
+The figure follows the format of Fig.
+2 in the main text.
+The text-guided reconstruction and MindEye2 methods did not show a large performance drop, even when the training sample size of the NSD dataset was reduced to match that of the Deeprecon dataset.
+This result supports that the degraded performance in the Deeprecon is not simply due to its smaller sample size compared to the NSD.
+Fig.
+A2: UMAP visualization of CLIP text features in the Natural Scene Dataset (NSD) using default parameters.
+This scatter plot, analogous to Fig.
+4A in the main text, depicts the distribution of text features within the NSD.
 The gray points represent training samples, and the orange points represent test samples.
-Unlike Fig 4A, which used parameters optimized for clustering visualization, this figure employs default UMAP settings.
-Despite the difference in parameters, the visualization still reveals distinct clusters and substantial overlap between training and test samples.
-S2 Fig.
-UMAP visualization of CLIP text features in the Deeprecon dataset.
-This scatter plot visualizes the distribution of semantic concepts within the Deeprecon data, with the gray points representing training samples and the orange points indicating test samples.
-Unlike the NSD dataset (Fig 4A in the main text), this visualization demonstrates a clearer separation between training and test samples.
+Unlike Fig.
+4A, which used the parameters optimized for clustering visualization, this figure employs the default UMAP settings.
+Despite the absence of parameter tuning, the plot still reveals discernible clusters and considerable overlap between training and test samples, indicating that the observed clustering pattern does not hinge on specialized UMAP configurations.
+Fig.
+A3: UMAP visualization of CLIP text features in the Deeprecon dataset.
+This scatter plot visualizes the distribution of semantic features within the Deeprecon data, with the gray points representing training samples and the orange points indicating test samples.
+Unlike the NSD dataset (Fig.
+4A in the main text), this visualization demonstrates a clearer separation between training and test samples.
 This distinction highlights the Deeprecon dataset’s intentional design to differentiate object categories between training and test sets.
-S3 Fig.
-Confusion matrices from the hold-out analyses.
+Fig.
+A4: UMAP visualization of the latent features of MindEye2 (OpenCLIP ViT/bigG-14) in the NSD and Deeprecon datasets.
+The plots follow the format of Fig.
+4A and Fig.
+A3.
+While these features are extracted from images, not from text captions, the plot still reveals a cluster structure, with considerable overlap between the training and test sets in the NSD dataset, whereas the Deeprecon dataset exhibits noticeably less overlap.
+Fig.
+A5: Distribution of similarity scores between training and test stimuli.
+Violin plots show the distribution of DreamSim-based similarity scores (Fu et al., 2024) between each dataset’s training set and its test stimuli (green) compared to the similarity distributions between the training set and an independent dataset (orange; randomly selected 1000 images in CC3M; Sharma et al., 2018).
+For each dataset, we plotted the distribution of similarity scores from the top $5\%$ most similar training images to either test or independent set images.
+Results remained consistent when varying the percentage or number of selected training images.
+A large deviation between the test- and independent-set distributions indicates that the test set is biased toward the training set, while smaller deviations suggest greater test set independence.
+This analysis was performed on seven datasets: Deeprecon ($1,200$ training and $90$ test images) and NSD (8,859 training and $982$ test images), and was further extended to THINGS-fMRI1 ($8,640$ training and $100$ test images; Hebart et al., 2023), THINGS-MEG ($22,248$ training and $200$ test images; Hebart et al., 2023), DreamDiffusion ($1,330$ training and $333$ test images; Bai et al., 2024), Brain2GAN (4,000 training and 200 test images; Dado et al., 2024), and TVSD ($22,248$ training and $100$ test images; Papale et al., 2025).
+Most datasets exhibited higher training-test similarity compared to training-independent similarity, suggesting these splits are not optimal for evaluating zero-shot prediction.
+In contrast, the Deeprecon dataset was specifically designed to exclude overlapping training categories, facilitating a more suitable evaluation of generalizability performance.
+The distributions of training-independent similarity were consistent across the datasets, suggesting comparable levels of representativeness among the datasets.
+Fig.
+A6: Confusion matrices from the hold-out and naive splits.
 (A) Clustering results of CLIP text features.
-This panel shows the outcome of applying $k$-means clustering to the UMAP embedding of CLIP text features from the NSD dataset (Fig 4A in the main text).
-These clustering results were used to determine the assignment of training and test samples under the hold-out split condition (Fig 6A in the main text).
+The $k$-means clustering was applied to the UMAP embedding of CLIP text features from the NSD (Fig.
+4A in the main text).
+These clustering results were used to make a training–test split under the hold-out split condition (Fig.
+6A).
 (B) Confusion probability matrices for cluster identification.
-These matrices illustrate the performance of cluster identification under different conditions.
-The top row shows results under the hold-out split conditions, while the bottom row presents results under the naive split conditions.
-For each row, the left matrix represents CLIP text features, and the right matrix represents CLIP vision features.
-In each matrix, the cell at position $(i,j)$ indicates the proportion of samples originally from cluster $i$ that were predicted to belong to cluster $j$.
-As visualized in the confusion matrices, the hold-out split condition, where entire semantic clusters are excluded from training, reveals the challenges in zero-shot prediction of novel concepts.
-In contrast, the naive split condition, where samples from all clusters are included in the training, demonstrates the model’s performance when semantic concepts overlap between training and test sets.
-S4 Fig.
-Extended simulation results for clustered data analysis.
+Left and right matrices represent CLIP text and vision features, respectively, with each cell $(i,j)$ indicating the proportion of samples from cluster $i$ predicted as cluster $j$.
+The top row shows results for the hold-out split, where entire semantic clusters are excluded from training, highlighting challenges in zero-shot prediction of novel categories.
+The bottom row shows the naive split, where semantic clusters appear in both training and test sets, demonstrating the translators’ ability to identify learned semantic categories.
+Fig.
+A7: Reconstruction analysis under hold-out split conditions using the NSD dataset.
+Stimulus images were clustered via UMAP of CLIP text features into “held-out” and “training” groups; models were trained only on training clusters and evaluated on held-out clusters using independent brain data.
+(A) DreamSim similarity distributions for various numbers of held-out clusters.
+The figure shows similarities between the training set and the hold-out test set (green) versus those between the training set and an independent dataset (CC3M; orange) as in Fig.
+A5.
+The x-axis indicates the number of excluded clusters.
+Despite removals, the test set remains more similar to the training set than to CC3M.
+The “similarity-matched hold-out” condition aligns these similarities with Deeprecon (see Fig.
+A5).
+The “naive split” (right side of the dashed line) randomly removes samples from the original training set so that the total number of training samples matches that of the similarity-matched hold-out split.
+Unlike the similarity-matched hold-out split, the naive split allows cluster overlap between training and test sets.
+(B, C) Reconstructions under different split conditions for StableDiffusionReconstruction (B) and Brain-Diffuser (C).
+Each row shows the original test image (top), followed by reconstructions from the naive split, the 1-cluster hold-out, and the similarity-matched hold-out.
+All splits use the same number of training samples for fair comparison.
+For StableDiffusionReconstruction, regression parameters were adjusted to accommodate the smaller training set, yet the outputs remained low-quality.
+Although the naive-split results retain some semantic similarity to the test images, using a 1-cluster hold-out led to visibly and semantically different reconstructions from the original test images.
+Brain-Diffuser reconstructions reveal clearer differences across conditions: naive-split outputs closely resemble the test images, 1-cluster hold-out reconstructions preserve the overall layout with minor semantic changes, and similarity-matched hold-out reconstructions deviate substantially in both visual and semantic content.
+Fig.
+A8: Recovery check for MindEye2.
+The panels follow the format of Fig.
+7 in the main text.
+MindEye2 uses a two-stage generation process.
+Left (SDXL unCLIP): In the initial stage, latent features are fed into a fine-tuned SDXL (Podell et al., 2023) with unCLIP technique (Ramesh et al., 2022; Scotti et al., 2024) to generate images, which closely match the originals.
+Right (MindEye2): These initial images are then refined in a second stage using base SDXL and captions predicted from the latent features (via GiT Image2Text modules; Wang et al., 2022).
+Because this second stage incorporates text information, the final MindEye2 outputs show lower fidelity than the SDXL unCLIP results.
+Fig.
+A9: Impact of the nonlinear translator on MindEye2 reconstructions.
+The first row shows the test images (Deeprecon), and the second row presents the full MindEye2 reconstructions, which use a nonlinear translator followed by a two-stage generator (SDXL unCLIP and then SDXL with predicted captions).
+To isolate the effects of the generator utilizing GiT Image2Text modules, the third and fourth rows compare only the first-stage outputs (SDXL unCLIP) using a nonlinear translator (third row) versus a linear translator (fourth row).
+Nonlinear translators often revert to object categories seen in the training set, whereas linear translators do not exhibit this tendency, suggesting that nonlinear models may be more prone to output dimension collapse.
+Fig.
+A10: Extended simulation results for clustered data analysis.
 (A) Relationship between cluster identification accuracy and cluster variance ratio.
@@
 The $x$-axis represents the cluster variance ratio $\sigma_{\mathrm{intra}}^{2}/\sigma_{\mathrm{inter}}^{2}$, and the $y$-axis represents the cluster identification performance.
-The green lines represent results for in-distribution test samples, and the orange lines show outcomes for out-of-distribution (OOD) samples.
-As the cluster variance ratio increases, clusters tend to overlap more, leading to difficulty with identification in the in-distribution condition.
-However, larger cluster variance also allows for more extensive coverage of the feature space, potentially improving OOD prediction.
-Consequently, OOD curves typically show a peak at intermediate variance ratios, reflecting a balance between cluster distinctiveness and feature space coverage.
+The green and orange lines indicate in-distribution and out-of-distribution (OOD) test samples, respectively.
+Higher variance ratios increase cluster overlap, making in-distribution identification harder.
+However, they also expand feature space, improving OOD prediction.
+OOD performance peaks at intermediate variance ratios, where cluster separability and feature space coverage are balanced.
 (B) Heatmap of cluster identification accuracy.
@@
 The cluster variance ratio ($\sigma_{\mathrm{intra}}^{2}/\sigma_{\mathrm{inter}}^{2}$) is fixed at $0.1$ for all simulations.
-The left matrix shows results for in-distribution samples, while the right matrix displays outcomes for out-of-distribution (OOD) samples.
-In each matrix, cell $(i,j)$ indicates the cluster identification accuracy when the feature dimension is $i$ and the number of clusters used in training is $j$.
-This view illustrates the interplay between feature dimensionality and the number of training clusters in determining the model’s generalization capabilities.
-Notably, in the OOD condition, as the feature dimensionality increases, a greater number of training clusters becomes necessary to achieve a robust zero-shot prediction.
-S5 Fig.
-Frame similarity analysis in the dataset of Nishimoto et al.
-[89].
+The left matrix represents in-distribution samples, and the right represents OOD samples.
+Each cell $(i,j)$ indicates the accuracy for the feature dimension $i$ and $j$ training clusters.
+In the OOD condition, higher feature dimensionality requires more training clusters for robust zero-shot prediction and this relationship is linear, not exponential.
+Fig.
+A11: Frame similarity analysis in the dataset of (Nishimoto et al., 2011).
 Movie scenes were detected from the movie stimuli.